### v1

# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Inhomogeneous Coefficients

Abstract
We consider the adaptive-rank integration of general time-dependent advection-diffusion partial differential equations (PDEs) with spatially variable coefficients. We employ a standard finite-difference method for spatial discretization coupled with diagonally implicit Runge-Kutta schemes for temporal discretization. The fully discretized scheme can be written as a generalized Sylvester equation, which we solve with an adaptive-rank algorithm structured around three key strategies: (i) constructing dimension-wise subspaces based on an extended Krylov strategy, (ii) developing an effective Averaged-Coefficient Sylvester (ACS) preconditioner to invert the reduced system for the coefficient matrix efficiently, and (iii) efficiently computing the residual of the equation without explicitly reverting to the full-rank form. The proposed approach features a computational complexity of $\mathcal{O}(N{r^{2}}+{r^{3}})$, where $r$ represents the rank during the Krylov iteration and $N$ is the resolution in one dimension, commensurate with the constant-coefficient case [10]. We present numerical examples that illustrate the computational efficacy and complexity of our algorithm.
Keywords: Adaptive-rank, extended Krylov subspaces, implicit Runge-Kutta integrators, advection-diffusion equations, Sylvester preconditioning, matrix-equations.
1 Introduction
We propose an adaptive-rank numerical scheme for approximating the solution of the time-dependent variable-coefficient advection-diffusion equation. A challenge in existing numerical schemes is the curse of dimensionality, where the problem dimension $d$ leads to exponential growth in computational complexity.
These formulations often result in large matrix systems of the form $Ax=b$, with $A$ sparse and banded, of size $N^{d}$-by-$N^{d}$ and bandwidth $\mathcal{O}(N^{d-1}$), where $N$ is the resolution in one dimension and $d$ is the dimensionality. Direct solvers for such systems scale in general as $\mathcal{O}(N^{3d-2})$ [3]. The complexity of unpreconditioned Conjugate Gradient (a type of Krylov method) for diffusion problems scales as $\mathcal{O}(N^{3d/2})$ [6, 19]. Multigrid-preconditioned Krylov methods may improve this scaling to $\mathcal{O}(N^{d})$ [16], which is considered optimal. In all these cases, the presence of the dimension $d$ in the exponent makes computation of the inverse matrix challenging for fine-mesh resolutions of high-dimensional problems.
Our method seeks to overcome this dimensional scaling and provide an efficient approach with a complexity of $\mathcal{O}(N)$ to solving time-dependent variable coefficient advection-diffusion partial differential equations (PDEs). In two dimensions, our approach leverages adaptive-rank approximations of matrix solutions on a tensor-product grid. Adaptive-rank techniques have proven effective in reducing the computational complexity of high-dimensional problems by representing the solution matrix as a sum of the outer products of basis vectors. Specifically, we employ a truncated singular value decomposition (SVD), such that $X=USV^{\top}$, where basis vectors in $U$ and $V$ span the solution space and determine the rank of the solution. The key idea is to dynamically update the basis vectors, as well as the rank (denoted as $r$), to maintain accuracy while minimizing computational cost.
Two primary strategies in the literature exploit low-rank structures for time-dependent PDEs: the dynamical low-rank (DLR) approximation [17, 22, 5, 8, 21] and the step-and-truncate (SAT) method, which has both explicit [18, 7, 13, 12, 14] and implicit [29, 24, 10, 23] formulations. Our work aligns with the SAT framework, where we perform temporal discretization with high-order accuracy and fully implicitly, leading to matrix differential equations to be solved at every time step. For constant coefficients, this system becomes a Sylvester equation of the form:
$A\mathbf{X}_{1}+\mathbf{X}_{1}B^{\top}=\mathbf{X}_{0},$ (1)
where $\mathbf{X}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ represents the solution at time $t_{1}$, $\mathbf{X}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$ is the initial condition at $t_{0}$, and $A$ and $B$ are difference operators arising from the PDE discretization.
Krylov subspace methods have proven effective for solving Sylvester-type equations resulting from the discretization of PDEs [20, 33, 29, 25, 2]. In these methods, the solution bases are chosen from a Krylov (or Krylov-like) subspace of $A$ and $B$ in (1), and the original system is projected onto this reduced subspace using the Galerkin projection. To accelerate the convergence of iterative methods, various extensions of Krylov subspaces have been developed. Notably, the extended Krylov subspace [9] involving $A^{-1}$, $B^{-1}$ has been applied to matrix equations in [33] and later adopted for time-dependent PDEs in [10].
While substantial progress has been made in constant-coefficient problems, systems with variable coefficients remain challenging. Discretizing such systems leads to the so-called generalized Sylvester matrix equation, of the form:
$\sum_{i=1}^{l}A_{i}\mathbf{X}_{1}{B}^{\top}_{i}=\mathbf{X}_{0},$ (2)
where $A_{i}$ and $B_{i}$ represent discretization operators for advection and/or diffusion terms. While these types of matrix equations can be solved by reverting them to the Kronecker form, in [26] the authors argue that, for a chosen approximation space, the Galerkin projection applied to the matrix equation consistently achieves faster convergence than the Kronecker formulation, particularly when the operators are poorly conditioned. Both analytical and numerical evidence in [26] support this claim. Furthermore, they show that the optimality of the Galerkin projection technique for Sylvester equations extends naturally to the generalized Sylvester equations. That is, regardless of the approximation space, the residual decreases as the subspaces are iteratively expanded, though the convergence rate depends significantly on the choice of subspaces, as demonstrated in [10]. Additional validation of this claim is provided in [34], where a matrix-based conjugate gradient method is compared with Galerkin projection techniques. Such results encourage the community to adopt matrix-equation-based strategies for multi-term matrix equations, rather than defaulting to Kronecker formulations.
Several methods have been proposed to solve generalized Sylvester equations iteratively directly in matrix form, either with stationary iterative techniques or Krylov methods. Stationary iterative methods [1, 32] leverage low-rank techniques to enhance efficiency, achieving computational complexity as low as $\mathcal{O}(N)$ when implemented effectively [32]. However, their effectiveness strongly depends on whether or not such a dominating operator could be correctly identified to enable a contractive iterative process. The difficulty of explicitly finding such a dominating term in a general setting was discussed in [25], which proposed using the Kronecker formulation instead.
Krylov methods have also been explored as an alternative for the solution of equations of the Sylvester and generalized Sylvester type. In [1], the authors present a series of techniques to tackle the Generalized Lyapunov equation, which is a special case of (2) where the first two terms are of Sylvester type, i.e. $A_{1}=B_{2}:=A$, $B_{1}=I_{N_{2}}$, and $A_{2}=I_{N_{1}}$. Among these techniques, they discuss a low-rank projection-based scheme, in which they propose constructing extended Krylov subspaces using $A$, following the approach of [33], and employing a standard Krylov basis for the remaining operators. In [27], a low-rank projection-based technique is employed for systems where $N_{2}\ll N_{1}$. The authors opt for the rational Krylov as an approximation space for the Galerkin projection of the original equation, and solve the resulting small-sized equation with either matrix-based Conjugate Gradient or a Kronecker formulation, given that $N_{2}$ is relatively small. In [15], a Krylov-based projection technique is developed for matrix equations of the form (2), where the Krylov basis construction relies on a low-rank decomposition of the operators. The resulting small generalized Sylvester matrix system is solved using the Sherman-Morrison-Woodbury formula through a series of Sylvester equation solves, which leverages the operator’s low-rank structure. Finally, of particular relevance to our study, a basis-update-and-Galerkin (BUG) DLR integrator is proposed in [23] as a low-rank GMRES preconditioner for the generalized Sylvester equation arising from advection-diffusion time-dependent problems with variable coefficients. The approach evolves the basis and the coefficient matrix in the preconditioner using BUG to provide guesses for the basis vectors and coefficient updates. The advantage of the approach is that the approximations involved in the BUG integrator are only present in the preconditioning stage, and not in the converged solution.
In contrast to existing efforts, in this work we propose an optimal implicit time integrator based on SAT adaptive-rank integrators for the variable-coefficient advection-diffusion equation. Instead of evolving the bases and coefficients simultaneously, we evolve them separately to produce an effective, inexpensive implicit solver. In particular, we iteratively construct a set of extended Krylov subspaces (for each dimension) that effectively capture the solution’s range evolution in an iterative manner. A Galerkin projection is then applied to project the full system of equations into a reduced system. The reduced system is solved using Krylov methods like GMRES, with preconditioning based on an approximated operator that transforms the generalized Sylvester equation into a regular Sylvester one, for which suitable efficient solvers exist. To assess the goodness of the solution efficiently, we propose a residual-checking mechanism that leverages the low-rank structure of solutions directly. Such procedure avoids the need to form the solution matrix for norm evaluation. The overall scheme scales as $\mathcal{O}(Nr^{2}+r^{3})$, with $r$ the rank. The main innovations of our proposed algorithm are as follows:
  • Linear “1D-like” Computational Complexity: We introduce a Krylov-based adaptive-rank algorithm for solving equations of the form (2), which arise from implicit integrators for advection-diffusion-type equations. Our method achieves linear computational complexity with respect to $N$ and quadratic complexity with respect to $r$, offering a significant advantage over existing state-of-the-art adaptive-rank approaches. To the best of our knowledge, this is the first study that clearly demonstrates the optimal complexity of $\mathcal{O}(Nr^{2})$ for the advection-diffusion-type equation with variable coefficients.
  • Generalized dimension-wise extended Krylov Subspaces: We propose an averaged-coefficient approximation operator $\tilde{\mathscr{L}}$, which is instrumental not only in building dimension-wise extended Krylov subspaces, but also for preconditioning the reduced projected system. The dimension-wise Krylov subspaces are generated from multiple operators present in ${\mathscr{L}}$, $\tilde{\mathscr{L}}$ and their inverses. Such strategy is numerically demonstrated in this study to be computationally effective with rapid convergence in Krylov iterations.
  • Efficient Preconditioning Strategy: We deploy a GMRES solver for the reduced generalized Sylvester equation from Galerkin projection (of size $r^{2}$), with an Averaging-Coefficient Sylvester (ACS) preconditioner to accelerate the Krylov convergence. The proposed preconditioner is based on a Sylvester direct solver for the averaged-coefficient approximation operator $\tilde{\mathscr{L}}$, which results in a computational complexity of $\mathcal{O}(r^{3})$ (vs. the expected $\mathcal{O}(r^{6})$ of a direct inversion of the Kronecker-product form of the reduced generalized Sylvester equation). This preconditioner is inspired by the recent work in [25], which employed a similar approach for the full-matrix form of the generalized Sylvester equation. A significant enhancement of the algorithm’s efficiency is demonstrated in our numerical tests.
  • Optimized solution representation and low rank residual computation: We employ an SVD-truncated QR procedure for basis augmentation, and an SVD-truncated procedure for the intermediate solution during the Krylov iteration as often as possible. This optimizes the computational efficiency, while preserving the accuracy and structure of the solution. In addition, an efficient low-rank computation of the residual is enabled with $\mathcal{O}(r^{3})$ complexity without explicitly forming the full solution on tensor-product grids. This avoids the computational cost associated with full matrix, which could otherwise spoil the linear computational complexity sought.
The structure of the rest of this paper is as follows. In Section 2, we consider the case of rank-one advection and diffusion coefficients and derive the generalized Sylvester equation that results from a finite-difference discretization in space and backward Euler in time. In Section 3, we present our adaptive-rank solution strategy, including the coefficient-averaged approximation operator, construction of extended-Krylov subspaces, preconditioning technique in solving the reduced system, and low-rank computation of the residual norm. The extension of the technique to high-order temporal discretizations with multi-rank coefficients is provided in Appendix C. In Section 4, numerical results are presented to demonstrate the efficacy in high-order temporal convergence, adaptive rank evolution of Krylov subspaces, optimal computational complexity, as well as accelerated Krylov convergence. Finally, we conclude with a brief highlight of our proposed algorithm and outlook for future research in Section 5.
2 Problem Description and Generalized Sylvester Equation
In what follows, we refer to equations in the form of Equation (1) as the Sylvester equation, where the unknown solution matrices are operated on one dimension at a time. Additionally, we use the term generalized Sylvester equation to describe equations in the form of Equation (2), where each unknown solution matrix is simultaneously operated on across both dimensions.
We establish the following notation conventions based on the nature of the symbols. Lowercase Greek letters, such as $\phi(x,y)$ or $\sigma(x,y)$, denote continuous scalar fields. Uppercase bold italic Greek letters, such as $\bm{\mathit{\Phi}}(x,y)$ or $\bm{\mathit{\Sigma}}(x,y)$, are used for vector-valued and tensor-valued functions. Uppercase Greek letters, like $\Phi$ or $\Sigma$, represent discrete matrices arising from the discretization of continuous scalar functions or differential operators on a grid. Boldface uppercase letters, such as $\mathbf{F}$ or $\mathbf{S}$, represent unknown matrix solutions or matrix right-hand sides in matrix equations. The identity matrix of size $n$ is denoted by ${I}_{n}$.
We aim to solve the advection-diffusion equation:
$\frac{\partial f}{\partial t}=\nabla_{v}\cdot(\bm{\mathit{\Phi}}\cdot\nabla_{v}f)-\nabla_{v}\cdot(\bm{\mathit{\Sigma}}f),$ (3)
where $\bm{\mathit{\Phi}}\in\mathbb{R}^{2\times 2}$ is an isotropic diffusion tensor:
$\bm{\mathit{\Phi}}(x,y)=\begin{bmatrix}\phi^{x}(x,y)&0\\ 0&\phi^{y}(x,y)\end{bmatrix},$
and the advection vector $\bm{\mathit{\Sigma}}\in\mathbb{R}^{2}$ is given by:
$\bm{\mathit{\Sigma}}(x,y)=\begin{bmatrix}\sigma^{x}(x,y)\\ \sigma^{y}(x,y)\end{bmatrix}.$
We assume that the diffusion and advection coefficients are separable and they exhibit a low-rank structure. Specifically, the diagonal diffusion coefficients, $\phi^{x}(x,y)$ and $\phi^{y}(x,y)$, are given by:
$\phi^{x}(x,y)=\sum_{i=1}^{\ell_{x}}\phi^{1,x}_{i}(x)\phi^{2,x}_{i}(y),\quad\phi^{y}(x,y)=\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\phi^{2,y}_{i}(y).$
Similarly, for the advection coefficients we have:
$\sigma^{x}(x,y)=\sum_{i=1}^{k_{x}}\sigma^{1,x}_{i}(x)\sigma^{2,x}_{i}(y),\quad\sigma^{y}(x,y)=\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\sigma^{2,y}_{i}(y).$
Substituting these coefficients in their separable forms into Equation (3), we obtain:
$\frac{\partial f}{\partial t}$ $=\sum_{i=1}^{\ell_{x}}\frac{\partial}{\partial x}\left(\phi^{1,x}_{i}(x)\frac{\partial f(x,y)}{\partial x}\right)\phi^{2,x}_{i}(y)+\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\frac{\partial}{\partial y}\left(\phi^{2,y}_{i}(y)\frac{\partial f(x,y)}{\partial y}\right)$ (4)
$-\sum_{i=1}^{k_{x}}\frac{\partial}{\partial x}\left(\sigma^{1,x}_{i}(x)f(x,y)\right)\sigma^{2,x}_{i}(y)-\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\frac{\partial}{\partial y}\left(\sigma^{2,y}_{i}(y)f(x,y)\right),$
where, in each term, the component independent of the differentiation variable has been pulled out of the differentiation operator. Discretizing this equation using finite differences in space and an implicit method in time naturally leads to a multi-term matrix differential equation where the unknown solution is acted upon from both directions in each term and at the same time. Numerous strategies have been proposed in the literature to solve such equations, most notably the recent work by Palitta and Simoncini in [25]. In this work, inspired by their preconditioning technique [25], we further exploit the low-rank structure of the solution and design a projection-based algorithm with an optimal scaling of $\mathcal{O}(Nr^{2})$, i.e. a linear scaling with respect to the resolution $N$ per spatial dimension, and quadratic scaling with respect to the rank $r$.
2.1 Generalized Sylvester Equation from Implicit Discretization of Advection-Diffusion
To simplify the presentation, we assume rank-one advection and diffusion coefficients, i.e.:
$\phi^{x}=\phi^{1,x}(x)\phi^{2,x}(y),\quad\phi^{y}=\phi^{1,y}(x)\phi^{2,y}(y)^{\top},$
$\sigma^{x}=\sigma^{1,x}(x)\sigma^{2,x}(y),\quad\sigma^{y}=\sigma^{1,y}(x)\sigma^{2,y}(y).$
Thus, Equation (4) simplifies to:
$\frac{\partial f}{\partial t}$ $=\underbrace{\frac{\partial}{\partial x}\left(\phi^{1,x}(x)\frac{\partial f(x,y)}{\partial x}\right)\phi^{2,x}(y)+\phi^{1,y}(x)\frac{\partial}{\partial y}\left(\phi^{2,y}(y)\frac{\partial f(x,y)}{\partial y}\right)}_{\text{diffusion}}$ (5)
${-}\underbrace{\frac{\partial}{\partial x}\left(\sigma^{1,x}(x)f(x,y)\right)\sigma^{2,x}(y){-}\sigma^{1,y}(x)\frac{\partial}{\partial y}\left(\sigma^{2,y}(y)f(x,y)\right)}_{\text{advection}}.$
We discretize this equation in a two-dimensional tensor-product spatial grid with ${N}_{1}$ and ${N}_{2}$ grid points in the $x$- and $y$- directions, respectively. The grid points in the $x$- direction are denoted as $x_{i}$ for $i=0,\dots,N_{1}-1$, and similarly for the $y$-direction. We consider the point-wise multiplication of the advection and diffusion coefficients in the discrete by setting:
$\Phi_{1}$ $=\text{diag}(\phi^{1,y}(x_{i})),\Sigma_{1}=\text{diag}(\sigma^{1,y}(x_{i})),\quad i=0,\cdots,N_{1}-1,$
$\Phi_{2}$ $=\text{diag}(\phi^{2,x}(y_{j})),\Sigma_{2}=\text{diag}(\sigma^{2,x}(y_{j})),\quad j=0,\cdots,N_{2}-1.$
We approximate the spatial derivative with a three-point stencil, and obtain the following matrix differential equation for $\mathbf{F}$, for which its $(i,j)$ entry $F_{i,j}$ corresponds to the approximation of $f(x_{i},y_{j})$:
$\frac{\partial\mathbf{F}}{\partial t}=\underbrace{D_{xx}^{\Phi}\mathbf{F}\Phi_{2}^{\top}+\Phi_{1}\mathbf{F}{D_{yy}^{\Phi}}^{\top}}_{\text{\emph{diffusion}}}-\underbrace{D_{x}^{\Sigma}\mathbf{F}\Sigma_{2}^{\top}-\Sigma_{1}\mathbf{F}{D_{y}^{\Sigma}}^{\top}}_{\text{\emph{advection}}}.$ (6)
Here, $D_{xx}^{\Phi},D_{yy}^{\Phi},D_{x}^{\Sigma},D_{y}^{\Sigma}$ represent compositions of difference operators with pointwise multiplication operators. See Appendices A and B for more details. For notational simplicity, let
$T_{1}=D_{xx}^{\Phi},\quad T_{2}=D_{yy}^{\Phi},\quad T_{3}=-D_{x}^{\Sigma},\quad T_{4}=-D_{y}^{\Sigma}.$ (7)
Equation (6) can be written compactly as:
$\frac{\partial\mathbf{F}}{\partial t}=\mathscr{L}(\mathbf{F}),$ (8)
where the right-hand side induces the following mapping:
${\mathscr{L}}:\mathbf{F}\mapsto T_{1}\mathbf{F}\Phi_{2}^{\top}+\Phi_{1}\mathbf{F}T_{2}^{\top}+T_{3}\mathbf{F}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}T_{4}^{\top}.$ (9)
We consider first a backward Euler implicit treatment of the temporal component of the matrix differential equation (8), which gives the following generalized Sylvester equation:
$\mathbf{F}_{1}-\Delta t\mathscr{L}({\bf F}_{1})=\mathbf{F}_{0},$ (10)
with the residual equation defined as:
$\mathbf{R_{\mathscr{L}}}=\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}=\mathbf{0}.$ (11)
An iterative solver naturally updates the solution $\mathbf{F}_{1}$ until the residual is sufficiently small under a suitable norm.
If we vectorize the right-hand side using the Kronecker product identity, $\text{vec}(A\mathbf{F}B^{\top})=(B\otimes A)\text{vec}(\mathbf{F})$, the second term with a factor of $\Delta t$ in (10) can be written as a matrix-vector multiplication of the form:
$\mathbf{F}\mapsto\left[\underbrace{\left(\Phi_{2}\otimes T_{1}\right)+\left(T_{2}\otimes\Phi_{1}\right)+\left(\Sigma_{2}\otimes T_{3}\right)+\left(T_{4}\otimes\Sigma_{1}\right)}_{\in\mathbb{R}^{N_{1}N_{2}\times N_{1}N_{2}}}\right]\text{vec}(\mathbf{F}),$ (12)
The Kronecker form yields a linear system $Ax=b$, which can be solved either directly or iteratively. The Kronecker approach offers valuable theoretical insights into the relationship between the Kronecker structure and the constructed Krylov basis, a connection exploited in the context of Sylvester equations in [20, 23]. We note that, based on the discussion in [26] on the advantages of the dimension-wise Krylov subspaces projection approach vs. the Kronecker formulation, we choose to adopt the former approach for better computational efficiency and iterative convergence.
3 Krylov-Based Adaptive-Rank Solver for Generalized Sylvester Equation
We propose an adaptive-rank algorithm for the generalized Sylvester equation (10). We assume a low-rank factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ at time $t^{(1)}=t^{(0)}+\Delta t$ also in the low-rank format. We represent this as:
[…]
where $U_{0}\in\mathbb{R}^{N_{1}\times r_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, and $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$ and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$ are diagonal matrices with singular values at the corresponding times. Our method involves constructing the bases $U_{1}$ and $V_{1}$ and then projecting the original equation to determine the coefficients of $S_{1}$. A key aspect of the proposed adaptive-rank approach is that we do not form the full solutions explicitly. Instead, we seek bases in a dimension-by-dimension manner and then use them to perform Galerkin projection, resulting in a reduced system for the matrix of coefficients $S_{1}$. An SVD truncation procedure is then applied to update the orthonormal basis sets and realize an efficient low-rank representation of the updated solution. Key questions for solving the generalized Sylvester equation in the form of (10) that we address in this paper include:
  • How do we choose suitable subspaces to build $U_{1}$ and $V_{1}$ from $U_{0}$ and $V_{0}$?
  • How can we solve the reduced equation for $S_{1}$ efficiently?
  • How can we check the residual of the generalized Sylvester equation without forming the matrices explicitly?
Our main innovation involves leveraging an approximate operator denoted by $\tilde{\mathscr{L}}$ (which is of the Sylvester form) to the original operator ${\mathscr{L}}$ in (9). $\tilde{\mathscr{L}}$ approximates the diagonal matrices $\Phi_{i}$ and $\Sigma_{i}$ representing pointwise multiplication of diffusion and advection coefficients, respectively, by scaled identities. This approximation not only allows us to construct dimension-wise Krylov bases $U_{1}$ and $V_{1}$ effectively and efficiently, but also enables effective preconditioning in the spirit of [25]. Then we project Equation (10) onto the Krylov bases, resulting in a reduced system for the coefficient matrix $S_{1}$. In this section, we introduce the main steps of our proposed algorithm as illustrated in Figure 1, highlighting the intuition behind our design and innovations along the way. Section 3.1 covers the construction of the approximated operator $\tilde{\mathscr{L}}$. Section 3.2 discusses the dimension-wise basis construction related to step 1 of the flowchart. In Section 3.3, we present the reduced equation for the matrix of coefficients $S_{1}$ and introduce our ACS preconditioner, corresponding to step 2 (i) in the flowchart. Lastly, Section 3.4 outlines the solution truncation mechanism and the residual computation, as described in steps 2(ii)-(iii) of the flowchart.
Figure 1: Flow-chart of the proposed implicit adaptive-rank algorithm for a generalized Sylvester equation.
3.1 Averaged-coefficient Approximate operator $\tilde{\mathscr{L}}$
Our approach incorporates ideas from [25] and [10] to construct an approximate operator $\tilde{\mathscr{L}}$ and efficiently solve the time-dependent inhomogeneous advection-diffusion equation using a projection-based adaptive-rank integration framework. The approximated operator $\tilde{\mathscr{L}}$ serves two purposes: (i) constructing the solution bases $U_{1}$ and $V_{1}$ as extended Krylov subspaces, and (ii) preconditioning the solution of the coefficient matrix $S_{1}$. We first construct a basis informed by both the original operator $\mathscr{L}$ and the approximate operator $\tilde{\mathscr{L}}$ (defined below). Secondly, we derive a corresponding projected (reduced) system that we solve iteratively using GMRES with a preconditioner derived from $\tilde{\mathscr{L}}$.
We seek an approximation to operator (9) that retains its spectral content and is easier to invert. As proposed in [25], the diagonal operators can be approximated by scaling with an identity matrix, where the scaling factor is the average of the diagonal entries. For instance, let $\Phi_{i}\approx{\alpha}_{i}I$ and $\Sigma_{i}\approx\gamma_{i}I$, where $\alpha_{i}$ and $\gamma_{i}$ are the average of the diagonal entries of $\Phi_{i}$ and $\Sigma_{i}$, respectively. This approximation yields the following mapping:
~${\tilde{\mathscr{L}}}:\mathbf{F}\mapsto\left(\alpha_{2}T_{1}+\gamma_{2}T_{3}\right)\mathbf{F}+\mathbf{F}\left(\alpha_{1}T_{2}+\gamma_{1}T_{4}\right)^{\top}.$ (14)
A backward Euler temporal discretization using (14) yields the residual equation:
$\mathbf{R_{\tilde{\mathscr{L}}}}=\mathbf{F}_{1}-\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}=\mathbf{0},$
which leads to the Sylvester equation:
$(A_{1}+A_{3})\mathbf{F}_{1}+\mathbf{F}_{1}(A_{2}+A_{4})^{\top}=\mathbf{F}_{0}.$ (15)
Here:
$A_{1}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{2}T_{1}\right)\quad A_{3}=\left(\frac{1}{4}I-\Delta t\,\gamma_{2}T_{3}\right),\quad P_{1}=A_{1}+A_{3},$ (16)
$A_{2}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{1}T_{2}\right)\quad A_{4}=\left(\frac{1}{4}I-\Delta t\,\gamma_{1}T_{4}\right),\quad P_{2}=A_{2}+A_{4}.$ (17)
The solution to Equation (15) is an approximation to that of (10), but it can be efficiently computed using low-rank approximations [10, 33, 31, 30]. It can therefore be exploited as a viable preconditioner for a GMRES iterative solution of the generalized Sylvester equation (10). This was done in [25] with $\mathcal{O}(N^{3})$ computational complexity. In this work, we extend this approach by performing a Galerkin projection of both the original operator $\mathscr{L}$ and its approximation $\tilde{\mathscr{L}}$ on the low-rank bases before solving the generalized Sylvester equation with GMRES. This reduces the problem to a lower-dimensional subspace, where preconditioning can be applied more effectively on a smaller problem with $\mathcal{O}(r^{3})$ complexity, resulting in an overall computational complexity of $\mathcal{O}(Nr^{2}+r^{3})$.
3.2 Basis Construction
We begin by defining the generalized Krylov subspaces used for our basis construction at the $m$th iteration:
$\kappa_{m}(M_{1},\cdots,M_{l},U)=[U,\underbrace{M_{1}U,\cdots,M_{l}U}_{\text{1st augmentation}},\ldots,\underbrace{M_{1}^{m+1}U,\cdots,M_{l}^{m+1}U}_{\text{mth augmentation}}].$ (18)
where $M_{1},\cdots,M_{l}$ are operators of interest and their inverses arising in our generalized Sylvester equation, and $l$ is the number of operators used for each basis augmentation step.
In our setting, the basis construction involves several operators from the original generalized Sylvester equation (10) and its approximation (15). We first consider the action of the diagonal operators, which have been averaged to obtain the approximated operators, $\Phi_{1}$, $\Sigma_{1}$, $\Phi_{2}$, and $\Sigma_{2}$. We have found that including these matrices is crucial for accelerating convergence. Using these operators, we construct a standard Krylov subspace. Additionally, we use the inverse operators from the approximated operators $A_{1:4}$ in (16)-(17) to form an inverted Krylov subspace. Finally, we construct dimension-wise extended Krylov subspaces for the approximated operator $\tilde{\mathscr{L}}$ using $P_{1}$ and $P_{2}$. The reasoning behind these choices of operators for subspace construction is as follows. Following [10], we construct an extended Krylov subspace in the $x$-direction for the approximated Sylvester equation (15) using $P_{1}$ and its inverse. To further enrich the basis, we include the inverses of the individual operators $A_{1}$ and $A_{3}$, as they may contain information not captured by the inverse of $P_{1}$. This carefully chosen set of operators results in a compact representation of the solution matrix, enriching the basis while keeping its dimension small.
Combining these, we define the extended Krylov subspaces in the $x$ direction as follows:
$\kappa_{m}\left(P_{1},P_{1}^{-1},A_{1}^{-1},A_{3}^{-1},\Phi_{1},\Sigma_{1},U_{0}\right),$
We then apply an SVD-truncated QR, as outlined in Algorithm 1, to orthogonalize the subspace and obtain a basis $U_{1}^{(m)}\in\mathbb{{\bf R}}^{N_{1}\times r^{(m)}_{x}}$ and remove vectors that are closely linearly dependent. By leveraging the rank-revealing property of SVD, only those basis associated with significant singular values are retained. The same procedures are applied to the $y$-direction to obtain $V^{(m)}_{1}\in\mathbb{{\bf R}}^{N_{2}\times r^{(m)}_{y}}$. For notational simplicity, we will omit the superscript $(m)$ from now on, understanding that $U_{1}\in\mathbb{{\bf R}}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{{\bf R}}^{N_{2}\times r_{y}}$ refer to the bases of the Krylov subspaces at the $m$th iteration. This truncated orthonormalization strategy is particularly advantageous when the underlying processes (diffusion or advection) are weak. In such cases, the approximated coefficient matrices contribute minimally, reducing the impact of the operator application and enabling further basis truncation.
These constructed orthonormal bases will be used to represent the solution in the approximation subspace. The coefficients composing the linear combination of these basis vectors will be obtained by projecting the corresponding generalized Sylvester equation and evolving the matrix of coefficients $S_{1}$.
Input: Krylov Subspace $\kappa_{m}$; Tolerance $\epsilon_{\kappa}$.
Output: Truncated bases ${Q}$.
1
2Compute the reduced QR decomposition: $\{Q,R\}=\texttt{QR}\left(\kappa_{m}\right)$;
3
4Perform reduced SVD decomposition: $T_{1}\tilde{R}T_{2}=\texttt{SVD}(R)$, where $\tilde{R}=\text{diag}(\sigma_{j})$;
5
6Identify the last index $r$ such that $\sigma_{r}>\epsilon_{\kappa}$;
7
8Update the basis vectors: $Q\leftarrow QT_{1}(:,1:r)$;
9
Algorithm 1 SVD-truncated QR, $\mathscr{T}_{\epsilon_{\kappa}}$
3.3 Solving for the Matrix of Coefficients, $S_{1}$: ACS-preconditioned GMRES
With the orthonormal bases $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ at hand, we perform a Galerkin projection to derive a reduced generalized Sylvester equation for the matrix $S_{1}\in\mathbb{R}^{r_{x}\times r_{y}}$. The reduced Sylvester equation for $S_{1}$ arises from projecting the residual via the Galerkin condition $U_{1}^{\top}\mathbf{R}_{\mathscr{L}}V_{1}=0$, which yields:
${\bf S}_{1}-\Delta t\left(\tilde{T}_{1}{\bf S}_{1}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}{\bf S}_{1}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}_{1}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}{\bf S}_{1}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}_{1},\qquad\text{with}\qquad\tilde{{\bf B}}_{1}=(U_{1}^{\top}U_{0})S_{0}(V_{1}^{\top}V_{0})^{\top}.$ (19)
This matrix-based equation is equivalent to the classical Kronecker formulation:
$\underbrace{\left[I_{r_{x}r_{y}\times r_{x}r_{y}}-\Delta t\left(\tilde{\Phi}_{2}\otimes\tilde{T}_{1}+\tilde{T}_{2}\otimes\tilde{\Phi}_{1}+\tilde{\Sigma}_{2}\otimes\tilde{T}_{3}+\tilde{T}_{4}\otimes\tilde{\Sigma}_{1}\right)\right]}_{\mathcal{A}}\text{vec}({\bf S}_{1})=\text{vec}(\tilde{{\bf B}}_{1}).$ (20)
The projected operators are defined as follows:
$\tilde{T}_{1}$ $=U_{1}^{\top}T_{1}U_{1},\quad\tilde{T}_{3}=U_{1}^{\top}T_{3}U_{1},\quad\tilde{\Phi}_{1}=U_{1}^{\top}\Phi_{1}U_{1},\quad\tilde{\Sigma}_{1}=U_{1}^{\top}\Sigma_{1}U_{1},$
$\tilde{T}_{2}$ $=V_{1}^{\top}T_{2}V_{1},\quad\tilde{T}_{4}=V_{1}^{\top}T_{4}V_{1},\quad\tilde{\Phi}_{2}=V_{1}^{\top}\Phi_{2}V_{1},\quad\tilde{\Sigma}_{2}=V_{1}^{\top}\Sigma_{2}V_{1}.$
Solving Equation (19) using the Kronecker formulation $\mathcal{A}\text{vec}({{\bf S}_{1}})=\text{vec}({\tilde{{\bf B}}_{1}})$ with a direct solver requires a complexity of $\mathcal{O}(r^{6})$, where $r=\text{max}(r_{x},r_{y})$, which may be too expensive even for moderate $r$-values. Instead, we solve this equation using preconditioned GMRES, where the operator-times-vector operation is performed using a matrix-based strategy, with complexity $\mathcal{O}(r^{3})$. This is possible due to the Kronecker structure of our operator, yielding the equivalence between (19) and (20). Given a vector $s_{k}=\text{vec}(\mathbf{S}_{k})$, the product of the matrix $\mathcal{A}$ times a vector $s_{k}$ is organized as follows:
  1. 1.
    Given $s_{k}$, reshape to $\mathbf{S}_{k}\in\mathbb{R}^{r_{x}\times r_{y}}$, such that $\text{vec}(\mathbf{S}_{k})=s_{k}$.
  2. 2.
    Compute $\hat{\mathbf{Z}}_{k}\leftarrow\mathbf{S}_{k}-\Delta t\left(\tilde{T}_{1}\mathbf{S}_{k}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}\mathbf{S}_{k}\tilde{T}_{2}^{\top}+\tilde{T}_{3}\mathbf{S}_{k}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}\mathbf{S}_{k}\tilde{T}_{4}^{\top}\right)$, with complexity $\mathcal{O}(r^{3})$.
  3. 3.
    Reshape $\hat{z}_{k}=\text{vec}(\hat{\mathbf{Z}}_{k})$.
The computational complexity of the GMRES solve also depends crucially on the total number of iterations needed to invert the linear system. To bound the number of GMRES iterations with respect to the basis dimension (rank) $r$, we require effective preconditioning, such that the eigenvalues of the preconditioned matrix are clustered and the number of iterations becomes independent of the system size $r$ [4]. Here, we propose a (left) preconditioner $\mathcal{P}$ such that the system:
$\mathcal{P}^{-1}\mathcal{A}\text{vec}({\bf S}_{1})=\mathcal{P}^{-1}\text{vec}(\tilde{{\bf B}}_{1}).$ (21)
is well conditioned. The preconditioning step in GMRES requires applying $\mathcal{P}^{-1}$ to an arbitrary vector provided by the algorithm.
For fast GMRES convergence, $\mathcal{P}$ should approximate the left-hand side of Equation (20) well while being easier to invert than the original system. Here, we consider $\mathcal{P}=I_{r_{y}}\otimes\tilde{P}_{1}+\tilde{P}_{2}\otimes I_{r_{x}}$ arising from the Galerkin projection of Equation (15). In particular, let $U_{1}$ and $V_{1}$ be the orthonormal basis for the generalized Krylov subspaces, and let $\tilde{P}_{1}=U_{1}^{\top}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{\top}P_{2}V_{1}$, then
$\mathcal{P}:\mathbf{S}\mapsto{\tilde{P}_{1}}\mathbf{S}+\mathbf{S}{\tilde{P}_{2}^{\top}}.$ (22)
To compute $\mathcal{P}^{-1}\hat{z}_{k}=z_{k}$, we solve the corresponding Sylvester equation:
$\tilde{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\tilde{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k},$
where $z_{k}=\text{vec}(\mathbf{Z}_{k})\in\mathbb{R}^{r_{x}r_{y}}$. Specific steps are as follows:
  1. 1.
    Given $\hat{z}_{k}$, reshape to $\hat{\mathbf{Z}}_{k}\in\mathbb{R}^{r_{x}\times r_{y}}$, such that $\text{vec}(\hat{\mathbf{Z}}_{k})=\hat{z}_{k}$.
  2. 2.
    Solve $\mathcal{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\mathcal{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k}$ with complexity $\mathcal{O}(r^{3})$, for example with the Bartels Stewart algorithm [11].
  3. 3.
    Reshape $z_{k}=\text{vec}(\mathbf{Z}_{k})$.
The proposed preconditioner effectively bounds the number of GMRES iterations required to solve Equation (19) independently of $r$, as will be shown later in this study, rendering the computational complexity of the $S$ matrix solve as $O(r^{3})$, much better than the expected $O(r^{6})$ scaling of a naïve implementation of GMRES.
3.4 Truncation, Efficient Residual Computation, and Algorithm Flowchart
At the end of each iteration, we need to compute the residual to accept or reject the low-rank approximation from Krylov subspaces. This process should take advantage of the low-rank representation of the solution while avoiding forming the residual on the full grid.
Once the solution ${\bf F}_{1}=U_{1}S_{1}V_{1}^{\top}$ is obtained, we apply truncation to control rank growth and reduce the residual computational load. Since $S_{1}$ from the GMRES solve is not necessarily diagonal, we follow the approach in [14] to convert it into a singular value form (also scaling as $\mathcal{O}(r^{3})$) and discard singular values below a threshold $\epsilon$. The truncation results in a solution ${\bf F}_{1}=U_{1}S_{1}V_{1}^{T}$, with $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$, and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$, with $r_{1}\leq\text{min}(r_{x},r_{y})$. This procedure is summarized in Algorithm 2.
Following a similar procedure to [10], we devise a cost-effective way to compute the residual using the identity:
$\|\mathbf{R_{{\mathscr{L}}}}\|$ $=\|\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}\|$
$=\underbrace{\left\|R_{U}\begin{bmatrix}{S_{1}}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&-\Delta t\,{S_{1}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,{S_{1}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{S_{1}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{S_{1}}\\ \end{bmatrix}R_{V}^{\top}\right\|}_{\in\mathbb{R}^{5r_{1}\times 5r_{1}}}$
$=\underbrace{\left\|R_{U}\begin{bmatrix}{S_{1}}-\tilde{B}_{1}&\mathbf{0}\\ \mathbf{0}&-\Delta t\,(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1})\\ \end{bmatrix}R_{V}^{\top}\right\|}\leq\epsilon_{\texttt{tol}},$
where $\mathcal{R}_{\text{ranks}}={\ell_{x}+k_{x}+\ell_{y}+k_{y}}$ is the total rank number of the advection and diffusion coefficients, and where $R_{U}$ and $R_{V}$ are the upper triangular matrices arising from the following QR decomposition:
{_$\{\_\ ,R_{U}\}$ $=\texttt{QR}\left(\begin{bmatrix}{U}_{1}&{T_{1}}{U}_{1}&\Phi_{1}{U}_{1}&{T_{3}}{U}_{1}&\Sigma_{1}{U}_{1}\end{bmatrix}\right),$ (23)
{_$\{\_\ ,R_{V}\}$ $=\texttt{QR}\left(\begin{bmatrix}{V}_{1}&\Phi_{2}{V}_{1}&{T_{2}}{V}_{1}&\Sigma_{2}{V}_{1}&{T_{4}}{V}_{1}\end{bmatrix}\right),$
where the orthogonal columns from the decomposition are ignored. The above residual result follows because, if ${\bf F}_{0}=U_{0}S_{0}V_{0}^{\top}$ (where $U_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ and $V_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ have orthonormal columns) then, since $U_{1}$ and $V_{1}$ contain $U_{0}$ and $V_{0}$, respectively, we have $U_{1}(U_{1})^{\top}{\bf F}_{0}V_{1}(V_{1})^{\top}={\bf F}_{0}$.
Input: Bases $U$, $V$; matrix of coefficients $S$; Tolerance $\epsilon$.
Output: Truncated bases $\tilde{U}$, $\tilde{V}$; truncated singular values $\tilde{S}$.
1
2Perform reduced SVD decomposition: $T_{1}\tilde{S}T_{2}=\texttt{SVD}(S)$, where $\tilde{S}=\text{diag}(\sigma_{j})$;
3
4Identify the last index $r_{1}$ such that $\sigma_{r_{1}}/\|S\|>\epsilon$;
5
6Update matrix of singular values: $\tilde{S}\leftarrow\tilde{S}(1:r_{1},1:r_{1})$;
7
8Update left singular vectors: $\tilde{U}\leftarrow UT_{1}(:,1:r_{1})$;
9
10Update the right singular vectors: $\tilde{V}\leftarrow VT_{2}(:,1:r_{1})$;
11
Algorithm 2 Truncated SVD, $\mathcal{T}_{\epsilon}$
3.5 Generalization to Multi-rank Coefficients
We generalize next the adaptive-rank approach from rank-one advection and diffusion coefficients to the multi-rank case, using backward Euler. The corresponding pseudo-algorithm is presented in Algorithm 3. High-order temporal discretizations are considered later in this study. Discretization of the multi-rank advection-diffusion equation yields a multi-term generalized Sylvester equation of the form:
$\mathbf{F}_{1}-\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}_{1}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}_{1}T_{2,j}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}_{1}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}_{1}T_{4,l}^{\top}}_{\mathscr{L}(\mathbf{F}_{1})}\right)={\bf F}_{0}.$ (24)
[…]
$\alpha_{x,i}$ $=\text{average}(\Phi_{i}^{2,x}),\quad\alpha_{y,j}$ $=\text{average}(\Phi_{j}^{1,y}),\quad\gamma_{x,k}$ $=\text{average}(\Sigma_{k}^{2,x}),\quad\gamma_{y,l}$ $=\text{average}(\Sigma_{l}^{1,y}),$ (25)
to arrive at the approximated Sylvester equation:
${P_{1}}\mathbf{F}_{1}+\mathbf{F}_{1}{P_{2}^{\top}}=\mathbf{F}_{0},$ (26)
where
$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{3,k}}=\sum_{i=1}^{\ell_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\alpha_{x,i}T_{1,i}\right)}+\sum_{k=1}^{k_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\gamma_{x,k}T_{3,k}\right)},$
$P_{2}$ $=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{4,l}}=\sum_{j=1}^{\ell_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\alpha_{y,j}T_{2,j}\right)}+\sum_{l=1}^{k_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{y,l}T_{4,l}\right)}.$
The bases $U_{1}$ and $V_{1}$ are constructed from orthonormalization of dimension-wise generalized Krylov subspaces from equations (24) and (26) as follows:
$U_{1}$ $=\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right)\right),$ (27)
$V_{1}$ $=\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right)\right),$ (28)
In the multi-rank setting, after a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$, we obtain a generalized Sylvester equation similar to (29) for the matrix of coefficients $S_{1}$:
${\bf S}_{1}-\Delta_{t}\left(\sum_{i=1}^{\ell_{x}}\tilde{T}_{1,i}{\bf S}_{1}\tilde{\Phi}_{i}^{2,x\top}+\sum_{j=1}^{\ell_{y}}\tilde{\Phi}_{j}^{1,y}{\bf S}_{1}\tilde{T}_{2,j}^{\top}+\sum_{k=1}^{k_{x}}\tilde{T}_{3,k}{\bf S}_{1}\tilde{\Sigma}_{k}^{2,x\top}+\sum_{l=1}^{k_{y}}\tilde{\Sigma}_{l}^{1,y}{\bf S}_{1}\tilde{T}_{4,l}^{\top}\right)=\tilde{{\bf B}}_{1}$ (29)
where the projected operators are analogous to those defined in the rank-one case.
[…]
$\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}=\tilde{B}_{1},$ (30)
with:
${\tilde{P}_{1}}$ $=\sum_{i=1}^{\ell_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{x}}-\Delta_{t}\alpha_{x,i}\tilde{T}_{1,i}\right)}_{\tilde{A}_{1,i}}+\sum_{k=1}^{k_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{x}}-\Delta_{t}\gamma_{x,k}\tilde{T}_{3,k}\right)}_{\tilde{A}_{3,k}},$
$\tilde{P}_{2}$ $=\sum_{j=1}^{\ell_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{y}}-\Delta_{t}\alpha_{y,j}\tilde{T}_{2,j}\right)}_{\tilde{A}_{2,j}}+\sum_{l=1}^{k_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{y}}-\Delta_{t}\gamma_{y,l}\tilde{T}_{4,l}\right)}_{\tilde{A}_{4,l}}.$
This yields the preconditioner operator:
$\mathscr{P}:\mathbf{S}\mapsto\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}$ (31)
The whole process is summarized in Algorithm 3.
// This algorithm solves Eq. (24) for $\mathbf{F}_{1}=U_{1}S_{1}V_{1}^{\top}$, where $\mathbf{F}_{0}=U_{0}S_{0}V_{0}^{\top}$
Input: Initial condition matrices $U_{0}$, $V_{0}$, $S_{0}$;
Operators $P_{1}$, $P_{2}$, $\{T_{1,i}\}_{i=1}^{\ell_{x}}$, $\{T_{2,j}\}_{j=1}^{\ell_{y}}$, $\{T_{3,k}\}_{k=1}^{k_{x}}$, $\{T_{4,l}\}_{l=1}^{k_{y}}$;
Tolerances $\epsilon_{\kappa}$, $\epsilon$, $\epsilon_{\texttt{tol}}$;
Maximum iterations max_iter;
Output: Updated bases $U_{1}$, $V_{1}$;
Truncated singular values $S_{1}$
1Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eq. (26);
2
3for $m=1$ to max_iter do
4        // Step K1.
5        Truncate to tolerance $\epsilon_{\kappa}$;
6        $U_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right)\right)$;
7
8       $V_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right)\right)$;
9
10       // Step K2.
11        Solve the reduced equation (29) for $S_{1}$ using GMRES-ACS (31) to tolerance $\epsilon_{\texttt{GMRES}}$;
12        // Step K3.
13        Truncate $\{U_{1},S_{1},V_{1}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}\})$ to tolerance $\epsilon$;
14        // Step K4.
15        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},T_{1,1:\ell_{x}}U_{1},\Phi_{1:\ell_{y}}^{1,y}U_{1},T_{3,1:k_{x}}U_{1},\Sigma_{1:k_{y}}^{1,y}U_{1}]\right)$ and
16        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},\Phi_{1:\ell_{x}}^{2,x}V_{1},T_{2,1:\ell_{y}}V_{1},\Sigma_{1:k_{x}}^{2,x}V_{1},T_{4,1:k_{y}}V_{1}]\right)$ ;
17
18       Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(S_{1}-\tilde{B}_{1},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1}\right)\right)R_{V}^{\top}\right\|$;
19        if $\|\mathbf{R}\|/\|\mathbf{F}_{0}\|\geq\epsilon_{\texttt{tol}}$ then
20               Reject solution and return to augment bases further;
21
22       else
23
24              Break;
25               Exit loop;
26
27
Algorithm 3 Backward Euler Adaptive-Rank Integrator for the Generalized Sylvester Equation. 00footnotetext: The MATLAB implementation of the algorithm is publicly available on GitHub, Adaptive-rank-ACS, https://github.com/helkahza1/Adaptive-rank-ACS
3.6 Computational Complexity
We now analyze the computational complexity of Algorithm 3 for backward Euler. For conciseness, we consider the same spatial resolution per dimension, i.e., $N=N_{1}=N_{2}$. Starting from an initial condition $U_{0}S_{0}V_{0}^{\top}$, where $U_{0}\in\mathbb{R}^{N\times r_{0}}$, $V_{0}\in\mathbb{R}^{N\times r_{0}}$, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$, the $m$-th extended Krylov iteration generates $U_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, $V_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, and $S_{1}\in\mathbb{R}^{r_{\kappa}\times r_{\kappa}}$, where $r_{\kappa}=((2+\mathcal{R}_{\text{ranks}})m+1)r_{0}$. Using SVD-truncated QR, this dimension is then reduced to smaller ranks, $r_{x}$ and $r_{y}$, for the $x$ and $y$ directions, respectively. The matrix $S_{1}$ from the $m$-th Krylov iteration solve, initially of size $r_{x}\times r_{y}$, is truncated to rank $r_{1}\leq\min(r_{x},r_{y})$ using Algorithm 2. To simplify the complexity analysis, similar to Section 3.3, let $r=\max(r_{x},r_{y})$. The computational complexity of each step is then estimated as follows:
Step K1
Constructing the generalized Krylov basis as outlined in lines (3) and (3) of Algorithm 3 requires performing the operations $A_{1,1:r_{x}}^{-1}U_{0}$ and $A_{3,1:k_{x}}^{-1}U_{0}$, and $A_{2,1:r_{y}}^{-1}V_{0}$ and $A_{4,1:k_{y}}^{-1}V_{0}$ which necessitates solving the systems $AX=U$ and $AX=V$, where the subscripts have been omitted for simplicity. This is typically achieved through an LU factorization of the operator. With 1D finite differences, $A$ has a tridiagonal structure, which can be efficiently factorized using techniques such as the Thomas algorithm, resulting in a complexity of $\mathcal{O}(N)$ per column of $X$ [28]. The same complexity holds for applying $P_{1}^{-1}$ and $P_{2}^{-1}$. Additionally, applying $P_{1}U_{0}$, $\Phi_{1,3}U_{0}$, $\Sigma_{1,3}U_{0}$, $P_{2}V_{0}$, $\Phi_{2,4}V_{0}$, and $\Sigma_{2,4}V_{0}$, can be done with $\mathcal{O}(Nr_{0})$ computational complexity for sparse diagonal and tridiagonal matrices, leading to a total complexity of $\mathcal{O}(N\mathcal{R}_{\text{ranks}}r_{0})$ to form the extended Krylov subspaces.
Regarding truncated orthonormalization in Algorithm 1, using a modified Gram-Schmidt method to compute the reduced QR decomposition of $U_{1}$ and $V_{1}$, in line (1), requires a complexity of $\mathcal{O}\left(Nr_{\kappa}^{2}\right)$. The SVD decomposition of the small-sized upper triangular matrix in line (1) requires complexity $\mathcal{O}\left(r_{\kappa}^{3}\right)$. The basis update in line (1) requires a complexity of $\mathcal{O}\left(Nr_{\kappa}r\right)$, where ${{r}}\leq r_{\kappa}$. Hence, the complexities of lines (3), (3) in Algorithm 3 scale as $\mathcal{O}\left(Nr_{\kappa}^{2}\right)$. Generally, we expect $r_{\kappa}\ll N$ when the solution exhibits a low-rank structure.
Step K2
The application of the operator using matrix-matrix multiplication and summation in the Krylov method (GMRES), as outlined in Section, 3.3, requires a complexity of $\mathcal{O}(r^{3})$. Furthermore, applying the preconditioner involves solving a Sylvester equation of size $r$, which can be achieved using the Bartels-Stewart algorithm [11] with a complexity of $\mathcal{O}(r^{3})$. Hence, the $S$-step has $\mathcal{O}(r^{3})$ complexity.
Step K3
The SVD decomposition of ${\widetilde{S}}$ in line (2) of Algorithm 2 is of $\mathcal{O}(r^{3})$, also independent of $N$. Updating the bases in lines (2) and (2) of Algorithm 2 requires $\mathcal{O}(Nr{r_{1}})$, where ${r_{1}}\leq r$ is the post-truncation rank of the new solution as described in Algorithm 2.
Step K4
The QR decomposition to obtain $R_{U}$ and $R_{V}$ in lines (3) of Algorithm 3 necessitates a complexity of $\mathcal{O}\left(N(\mathcal{R}_{\text{ranks}}r_{1})^{2}\right)$ The residual computation in line (3) is of $\mathcal{O}({r_{1}}^{3})$ (i.e., independent of $N$) due to the matrix multiplications, with an additional $\mathcal{O}({r_{1}}^{2})$ for the Frobenius norm.
We note that henceforth $r$ will represent the maximum of $r_{\kappa}$ and $\mathcal{R}_{\text{ranks}}r_{1}$. We conclude that the overall algorithm scales as $\mathcal{{O}}(Nr^{2}+r^{3})$, which motivates keeping the overall rank $r$ as low as possible. These scalings will be verified later in this study.
3.7 Extension to High-order Temporal Discretization
The DIRK method features a stage-by-stage backward Euler-like implicit solve, with explicit evaluation of the terms on the right side of the previous Runge-Kutta stages [10]. In Appendix C.1, the mathematical formulation of the DIRK schemes is elaborated leading to the following residual equation for the $k$-th RK stage,
${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\quad k=1,2,\ldots,s.$ (32)
Expanding the operator $\mathscr{L}$ in the residual equation above results in the following $k$-th stage generalized Sylvester equation:
$\mathbf{F}^{(k)}-a_{k,k}\Delta t\left(T_{1}\mathbf{F}^{(k)}\Phi_{2}{{}^{\top}}+\Phi_{1}\mathbf{F}^{(k)}T_{2}^{\top}+T_{3}\mathbf{F}^{(k)}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}^{(k)}T_{4}^{\top}\right)=\mathbf{B}^{(k)},\quad k=1,2,\ldots,s.$ (33)
For the DIRK scheme, we broadly adopt the strategy outlined in [10] with minor modifications to improve efficiency. In particular, we: (i) construct the basis using the time-step of the first stage and retain it for subsequent stages, (ii) evolve the projected system for the $S_{1}$ matrix across the DIRK stages, and (iii) truncate and evaluate the residual to determine whether to admit the solution or return to step (i) to further augment the basis. Key modifications vs. [10] include truncating the solution before the residual computation, and only checking the residual at the end of the DIRK step. The reduced matrix is solved similarly to the backward Euler method outlined above, using preconditioned GMRES, where the preconditioner is based on an approximate operator, $\tilde{\mathscr{L}}$. Further details on the discretization of the advection-diffusion equation using the DIRK scheme are provided in Appendix C. Specifically, Appendix C.2 describes the adaptive-rank approach for high-order integration, Appendix C.3 extends the integration to multi-rank coefficients with DIRK schemes and presents a pseudo-algorithm.
4 Numerical Experiments
We demonstrate the computational complexity estimates and the overall efficacy of the proposed algorithm for advection-diffusion equations with several numerical examples. We assume homogeneous Dirichlet boundary conditions for all the tests for simplicity. All simulations were performed using MATLAB. For benchmarking purposes, we compare against full-rank algorithm using the Kronecker formulation inverted with MATLAB’s backslash command. Throughout this section, we denote the diffusion Courant number and the advection Courant-Friedrichs-Lewy (CFL) number by $\lambda_{D}$ and $\lambda_{A}$, respectively, defined as:
$\lambda_{D}=\frac{\Delta t\phi_{\text{max}}}{\Delta x^{2}},\qquad\lambda_{A}=\frac{\Delta t\sigma_{\text{max}}}{\Delta x},$ (34)
where $\phi_{\text{max}}$ and $\sigma_{\text{max}}$ are the maximum of the diffusion and advection coefficients, respectively. We set algorithmic tolerances as follows. We first choose the residual tolerance, $\epsilon_{\texttt{tol}}$. One strategy to prevent rapid growth of the extended Krylov subspaces, especially when the matrix equations contain many terms, is to choose the Krylov basis truncation threshold as $\epsilon_{\kappa}\geq\epsilon_{\texttt{tol}}$ to truncate basis vectors that are nearly linearly dependent. However, over-truncation may risk removing important information, slowing down convergence. The GMRES tolerance $\epsilon_{\texttt{GMRES}}$ is chosen as $\epsilon_{\texttt{GMRES}}<10^{-2}\epsilon_{\texttt{tol}}$ to ensure that the error in the inner loop does not pollute the outer loop. The SVD truncation threshold $\epsilon$ are set to $\epsilon<10^{-2}\epsilon_{\texttt{tol}}$. Since truncation occurs before residual evaluations, the error introduced by solution truncation must remain below the residual tolerance. Otherwise, the computed solution may meet the residual threshold, but over-truncation could prevent this acceptance.
Example 4.1: Advection-diffusion with a rank-1 advection coefficient and a rank-3 diffusion coefficient
Consider Equation (3) with a rank-1 variable advection coefficient:
$\sigma^{x}(x,y)$ $=\sigma^{1}_{1}(x)\sigma^{2}_{1}(y),$ $\sigma^{1}_{1}(x)$ $=1-x^{2},$ $\sigma^{2}_{1}(y)$ $=2y,$
$\sigma^{y}(x,y)$ $=\sigma^{1}_{2}(x)\sigma^{2}_{2}(y),$ $\sigma^{1}_{2}(x)$ $=-2x,$ $\sigma^{2}_{2}(y)$ $=1-y^{2},$
and a rank-3 variable diffusion coefficient:
$\phi^{x}(x,y)=\phi^{y}(x,y)=\sum_{i=1}^{3}\phi^{1}_{i}(x)\phi^{2}_{i}(y),$
${\phi}^{1}_{1}(x)$ $=\exp\left(-\left(x-0.3\sin(x)\right)^{2}\right),$ ${\phi}^{2}_{1}(y)$ $=\exp\left(-\left(y-0.3\cos(y)\right)^{2}\right),$
${\phi}^{1}_{2}(x)$ $=\exp\left(-\left(x-0.6\sin(\pi x)\right)^{2}\right),$ ${\phi}^{2}_{2}(y)$ $=\exp\left(-\left(y-0.6\sin(\pi y)\right)^{2}\right),$
${\phi}^{1}_{3}(x)$ $=\exp\left(-\left(x-0.6\sin(2\pi x)\right)^{2}\right),$ ${\phi}^{2}_{3}(y)$ $=\exp\left(-\left(y-0.6\sin(2\pi y)\right)^{2}\right).$
The initial condition is a rank-2 Gaussian:
$f(x,y,0)=0.5\exp\left(-400\left((x-0.3)^{2}+(y-0.35)^{2}\right)\right)+0.8\exp\left(-400\left((x-0.65)^{2}+(y-0.5)^{2}\right)\right).$
We use a spatial grid with $N_{x}=N_{y}=300$, and the time-step $\Delta t$ is set to $\lambda_{D}\Delta x^{2}$, with $\lambda_{D}$ ranging from $280$ to $750$. We test our adaptive-rank algorithm using three different integrators: backward Euler (BE), DIRK2, and DIRK3, as detailed in Tables 3 and 3 in the Appendix. The residual tolerance is $\epsilon_{\texttt{tol}}=10^{-4}$ for BE, DIRK2, and DIRK3. The basis truncation threshold is $\epsilon_{\kappa}=10^{-3}$. The GMRES convergence threshold is $\epsilon_{\texttt{GMRES}}=10^{-8}$, and the SVD truncation threshold is $\epsilon=10^{-8}$.
Figure 2 illustrates the $L_{1}$ error norm vs time stepping size characterized by $\lambda_{D}$ by showcasing the performance of the adaptive-rank integrator for BE, DIRK2, and DIRK3. The expected temporal error order of convergence for our proposed adaptive-rank integrators, depicted by lines with markers, is observed.
Figure 2: Example 4.1. Temporal convergence study of BE, DIRK2, and DIRK3 schemes.
We performed a computational complexity study to demonstrate the proposed algorithm has a linear complexity with respect to $N$, the number of grid points per dimension. The mesh size $N=N_{x}=N_{y}$ varied between $100$ and $31000$ and the time step was fixed at $\Delta t=10^{-3}$. This range of mesh sizes is chosen to prevent the $O(r^{3})$ contribution of the coefficient matrix GMRES solver from polluting the mesh-resolution scaling.
Figure 3(a) displays $N$ vs. computational time. We observe that the complexity remains linear with respect to $N$.
Figure 3: Example 4.1.(a) Computational complexity of DIRK2 adaptive-rank integrator. (b) Log-log plot of GMRES-ACS solver. The execution times were recorded using MATLAB’s timeit function.
Figure 3(b) shows the computational time required for solving the projected generalized Sylvester equation via preconditioned GMRES-ACS vs. rank for a single time step using BE with $N=10,000$ and $\epsilon_{\texttt{GMRES}}=10^{-10}$. To examine the computational scaling corresponding to the expansion of the Krylov basis, we allowed the basis to grow without truncation, recording the simulation time required for the GMRES solve at each basis augmentation iteration. This figure verifies the expected cubic scaling, e.g. $\mathcal{O}(r^{3})$. From the above study, we verify that the predominant complexity of our approach is $\mathcal{O}(Nr^{2}+r^{3})$, effectively reducing the complexity of full rank methods, yet maintain their high-order convergence.
We investigate next the impact of the proposed preconditioner in the S-step by comparing the performance of the GMRES solver with and without the preconditioner. Figure 4 showcases the GMRES residual history as a function of the number of iterations for both preconditioned and unpreconditioned cases for mesh sizes ranging from $200$ to $2500$, in a single time step with $\Delta t=0.01$. It is observed that preconditioning significantly accelerates convergence, with the convergence rate independent of the mesh size, which is not the case for unpreconditioned GMRES.
Figure 4: Example 4.1. Semi-log plot of GMRES residuals versus iterations with and without preconditioning. The simulation is conducted for a single time step with $\Delta t=0.01$, on meshes of size $N_{x}=N_{y}=200$ in (a), $N_{x}=N_{y}=1000$ in (b), and $N_{x}=N_{y}=2500$ in (c).
Example 4.2: Adaptive Rank Evolution with Manufactured Solution
We assess next the performance of the proposed method to capture the rank adaptively using a manufactured solution defined as:
$f(x,y,t)$ $=\left(\sin(\pi x)\sin(\pi y)\right)\exp\left[-0.1\pi(t-2)^{2}\right]$
$\quad+\left(\sin(2\pi x)\sin(2\pi y)\right)\exp\left[-0.5\pi(t-10)^{2}\right]$
$\quad+\left(e^{\left[-100(x-0.3)^{2}-100(y-0.3)^{2}\right]}+e^{\left[-100(x+0.3)^{2}-100(y+0.3)^{2}\right]}\right)e^{\left[-0.5\pi(t-15)^{2}\right]}.$
This solution is initially rank-9 and fluctuates in rank over time, with significant changes at $t=1$, $10$, and $15$. Based on this solution, we analytically compute the source term $S(x,y,t)$. The diffusion coefficients are set as previously described in Example 4.1. We use a spatial grid of $N=1000$, the BE integrator and simulate until $T_{f}=18$. The relative residual tolerance is $\epsilon_{\texttt{tol}}=10^{-8}$. The GMRES convergence threshold is set to $\epsilon_{\texttt{GMRES}}=10^{-12}$. The singular-value truncation threshold is set to $\epsilon=10^{-5}$. The Krylov basis truncation threshold is $\epsilon_{\kappa}=10^{-5}$. We use a time step $\Delta t=\lambda_{D}\Delta x^{2}$ with $\lambda_{D}=1300$, or equivalently, $\Delta t=\lambda_{A}\Delta x$ with $\lambda_{A}=17$.
Figure 5(a) tracks the adaptive rank in time with our proposed adaptive rank integrator (in dashed red) vs the full-rank integrator (in solid blue). Agreement is excellent between the two. Figure 5(b) illustrates the evolution of the extended Krylov subspace over time, before and after applying the truncation step in Algorithm 3. Our approach effectively filters out low-relevance basis vectors, maximizing computational efficiency. The extended Krylov basis-augmentation history is shown in the semi-transparent red bars in Figure 5(b) (corresponding to right y-axis). It can be seen that the algorithm performs several ($\sim 3$) iterations of basis augmentation around the $t=1$, $10$, and $15$, to find an appropriate subspace to resolve the rank growth induced by the source term, settling into $\sim 1$ iteration otherwise. This demonstrates the effectiveness of the Krylov subspace formulation proposed here to find suitable low-rank subspaces as allowed by the dynamics of the solution.
Figure 5: Example 4.2. (a) Rank evolution of the adaptive-rank method using the backward Euler temporal update with a rank-3 manufactured solution. (b) Left axis: evolution of the extended Krylov subspace size over time, both before and after truncation. Right axis: number of Krylov basis iterations required for solution convergence.
Example 4.3: Swirling Deformation Flow
We consider solving the following advection-diffusion equation:
$f_{t}-\left[(1-x^{2})\cdot 2y\cdot f\right]_{x}+\left[2x\cdot(1-y^{2})\cdot f\right]_{y}=\nu\Delta f,\quad x,y\in[-1,1]$ (35)
The initial condition is a bi-Gaussian function given by:
$f(x,y,0)=0.5\exp\left(-\frac{1}{2\sigma^{2}}\left[(x-0.3)^{2}+(y-0.35)^{2}\right]\right)+0.8\exp\left(-\frac{1}{2\sigma^{2}}\left[(x-0.65)^{2}+(y-0.5)^{2}\right]\right),$
where $\sigma=0.15$. We choose the diffusion coefficient $\nu=10^{-3}$, small enough to weaken the diffusion process. To satisfy the cell-Reynolds number (to preserve monotonicity while keeping the spatial discretization linear; see Appendix B), we set $\Delta x=9.52\times 10^{-4}$, resulting in a mesh with $N_{x}=N_{y}=2100$. This test showcases the capability of our proposed low-rank integrator to handle fine meshes, ensuring that the numerical solution adheres to the required monotonicity criteria and avoids unphysical behavior. The residual tolerance is set to $\epsilon_{\texttt{tol}}=10^{-5}$, the basis truncation threshold is set to $\epsilon_{\kappa}=10^{-3}$, the SVD truncation threshold is set to $\epsilon=10^{-5}$, and the GMRES acceptance threshold is set to $\epsilon_{\texttt{GMRES}}=10^{-10}$.
Figure 6: Example 4.3. Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=7$ on a $2100\times 2100$. (a) Evolution of rank over time. (b)–(e) Solution profiles at $t=0$, $t=2.8487$, $t=5.7088$, and $t=9.9962$. The final simulation time is $T_{f}=10$.
Figure 6 demonstrates our algorithm’s ability to capture the solution rank for this challenging case. The solution rotates according to the prescribed advection coefficients and diffuses slowly over time. As the solution structure becomes more complex due to deformation, the rank increases to accurately represent its physical features. Notably, the algorithm dynamically adjusts the solution rank, which does not exceed $r_{1}=20$, two orders of magnitude smaller than the grid dimension $N=2100$.
Additionally, we evaluate the algorithm’s effectiveness in dynamically capturing the solution rank, even with large time steps. To this end, we use the proposed adaptive-rank algorithm with DIRK3 temporal integration and employ subsequent larger time steps, $\Delta t$, corresponding to CFL numbers $\lambda_{A}=1$ and $\lambda_{A}=15$. All other parameters are kept consistent with the simulation depicted in Figure 6, specifically $\epsilon_{\kappa}=10^{-3}$, $\epsilon_{\texttt{tol}}=10^{-5}$, $\epsilon=10^{-5}$, and $\epsilon_{\texttt{GMRES}}=10^{-10}$. Figure 7 shows the evolution of the generalized Krylov subspace over time, both before and after truncation, highlighting significant reductions in basis size and thus algorithmic complexity. The figure also demonstrates that the algorithm effectively captures the solution rank evolution even with fairly large (but dynamically relevant) advective CFLs at comparable performance. This demonstrates the algorithm’s capability to construct an appropriate Krylov basis, accurately evaluate the effects of the advection and diffusion operators, and dynamically truncate the subspace as needed.
Figure 7: Example 4.3. Evolution of the Extended Krylov subspace over time before and after truncation for different $\lambda_{A}$ values using DIRK3: (a) and (b) correspond to $\lambda_{A}=1$, while (c) and (d) correspond to $\lambda_{A}=15$.
Example 4.4: Constructed Steady-State Solution
In this example, we construct diffusion and advection coefficients that balance each other, resulting in a non-trivial steady-state solution. Specifically, we seek an equilibrium solution satisfying the steady-state condition:
$0=\nabla\cdot\left(\mathbf{\Phi}\cdot\nabla f_{\text{eq}}-\mathbf{\Sigma}f_{\text{eq}}\right),$ (36)
a solution of which is:
$\mathbf{\Sigma}=\mathbf{\Phi}\cdot\nabla\log(f_{\text{eq}}).$
This equation defines the steady-state flow from a provided diffusion coefficient and a steady-state solution. We choose the steady-state solution in the domain $0\leq x,y\leq 1$ to be:
$f_{\text{eq}}(x,y)=\left(x(1-x)\right)^{2}\left(y(1-y)\right)^{2},$
which satisfies homogeneous Dirichlet boundary conditions. The diffusion coefficients are defined as:
$\phi^{x}(x,y)=\phi^{y}(x,y)$ $=\phi^{1}(x)\phi^{2}(y),$
$\phi^{1}(x)$ $=x^{2}(1-x)^{2},$
$\phi^{2}(y)$ $=y^{2}(1-y)^{2}.$
The variable advection coefficients are then given by:
$\sigma^{x}(x,y)=\sigma^{1}(x)\sigma^{2}(y),\qquad D_{y}(x,y)=\chi^{1}(x)\chi^{2}(y),$
where:
$\sigma^{1}(x)$ $=2x(1-3x+2x^{2}),\qquad\sigma^{2}(y)=y^{2}(1-y)^{2},$
$\chi^{1}(x)$ $=x^{2}(1-x)^{2},\qquad\chi^{2}(y)=2y(1-3y+2y^{2}).$
We select the following rank-1 initial condition:
$f(x,y,0)=\left|\sin(2\pi x)\sin(2\pi y)\right|,\quad 0\leq x,y\leq 1.$
Since the solution will converge to a steady state up to a constant, we correct the mass of the initial condition to match that of the steady-state solution. To correct the mass of the solution, we adopt the LoMaC strategy outlined in [14, 13, 12], as implemented in [10]. Figure 8 shows the error of the numerical steady-state solution relative to the prescribed analytical steady state as a function of mesh size. The advection CFL number, $\lambda_{A}$, for a fixed time step $\Delta t=1000$ ranges from $19,050$ to $577,500$ as the mesh is refined, indicating the exceptionally large time steps allowed by the algorithm. The results confirm that second-order spatial convergence is achieved and demonstrate that, under these conditions, large time steps can be taken while still accurately reaching the correct steady state with asymptotic convergence.
Figure 8: Example 4.4. Log-log plot of error in the numerical steady-state solution compared to the prescribed analytical steady state as a function of mesh size. The advection CFL number, $\lambda_{A}$, ranges from $19,050$ to $577,500$. The plot demonstrates the retrieval of second-order spatial convergence in the steady state.
5 Conclusion
We have proposed a computationally efficient adaptive-rank implicit algorithm for solving time-dependent advection-diffusion partial differential equations with variable coefficients. We discretize the PDE using second-order centered finite differences in space and implicitly in time with up to third-order accuracy using DIRK, resulting in a generalized Sylvester matrix equation. Our algorithm is based on several key strategies: constructing dimension-wise basis from extended Krylov subspaces, projecting the original matrix equation onto a reduced system to solve for a coefficient matrix, devising an effective Averaged-Coefficient Sylvester (ACS) preconditioner for the reduced system, and a computationally efficient procedure for the residual evaluation. This combined approach achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{3})$, with $r$ being the rank of the solution during Krylov iteration process, effectively mitigating the curse of dimensionality and making it comparable in efficiency to constant-coefficient cases [10]. We have verified our algorithm through a series of numerical experiments. Overall, our adaptive-rank algorithm offers a scalable and efficient solution for time-dependent advection-diffusion type PDEs with variable coefficients. It effectively balances computational cost and accuracy by efficiently constructing and truncating the Krylov basis and dynamically adjusting the solution rank. The ability to handle large meshes and time steps without sacrificing accuracy makes it a promising tool for a wide range of applications. Future research directions include extending the algorithm to higher dimensions and nonlinear PDEs.
Acknowledgements
This work was partially supported by the Multifaceted Mathematics Integrated Capability Centers (MMICCs) program of DOE Office of Applied Scientific Computing Research (ASCR). Los Alamos National Laboratory (LANL) is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). J.Q. and H.E. were supported by NSF grant NSF-DMS-2111253. J.Q. was supported by Air Force Office of Scientific Research FA9550-22-1-0390, Department of Energy DE-SC0023164 and Air Force Office of Scientific Research (AFOSR) FA9550-24-1-0254 via the Multidisciplinary University Research Initiatives (MURI) Program. H.E. was also supported by MMICCs at LANL during the summer of 2024.
Appendix A Discretization of the Diffusion Operator
We provide a detailed derivation of the matrix formulation of the discretization of the diffusion operator for the two-dimensional advection-diffusion equation with separable diffusion coefficients. We consider the discretization of the diffusion term in the $x$-direction, with other directions treated similarly. Consider the diffusion term in the $x$-direction:
$\frac{\partial}{\partial x}\left(\phi^{x}(x,y)\frac{\partial u(x,y)}{\partial x}\right),$ (37)
where the diffusion coefficient is separable:
$\phi^{x}(x,y)=\phi^{1}(x)\phi^{2}(y).$ (38)
Our goal is to discretize this term using a 3-point stencil on a uniform grid with spacing $\Delta x$ in the $x$-direction. Let $u_{i,j}\approx u(x_{i},y_{j})$, with $i=0,\cdots,N_{x}-1$ and $j=0,\cdots,N_{y}-1$.
We approximate the spatial derivative using central differences. The discrete approximation at grid point $(i,j)$ is:
$\left[\frac{\partial}{\partial x}\left(\phi^{x}\frac{\partial u}{\partial x}\right)\right]_{i,j}\approx\frac{1}{\Delta x}\left[\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i+\frac{1}{2},j}-\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i-\frac{1}{2},j}\right],$ (39)
where the fluxes at the cell faces are approximated by:
$\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i+\frac{1}{2},j}$ $\approx\phi^{x}_{i+\frac{1}{2},j}\cdot\frac{u_{i+1,j}-u_{i,j}}{\Delta x},$ (40)
$\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i-\frac{1}{2},j}$ $\approx\phi^{x}_{i-\frac{1}{2},j}\cdot\frac{u_{i,j}-u_{i-1,j}}{\Delta x}.$ (41)
Here $\phi^{x}_{i+\frac{1}{2},j}=\phi^{1}_{i+\frac{1}{2}}\phi^{2}_{j}=\phi^{1}(x_{i+\frac{1}{2}})\phi^{2}(y_{j}).$ Substituting back into Eq. (39), we obtain:
$\left[\frac{\partial}{\partial x}\left(\phi^{x}\frac{\partial u}{\partial x}\right)\right]_{i,j}\approx\frac{1}{\Delta x^{2}}\left[\phi^{1}_{i+\frac{1}{2}}(u_{i+1,j}-u_{i,j})-\phi^{1}_{i-\frac{1}{2}}(u_{i,j}-u_{i-1,j})\right]\phi^{2}_{j}.$ (42)
To express the discretization in matrix form, we define:
  • $\mathbf{F}$: matrix of solution values, $[\mathbf{F}]_{i,j}=f_{i,j}$.
  • $\Phi^{1}_{+}$: diagonal matrix with elements $\left[\Phi^{1}_{+}\right]_{i,i}=\phi^{1}_{i+\frac{1}{2}}$.
  • $\Phi^{1}_{-}$: diagonal matrix with elements $\left[\Phi^{1}_{-}\right]_{i,i}=\phi^{1}_{i-\frac{1}{2}}$.
  • $\Phi_{2}$: diagonal matrix with elements $\left[\Phi_{2}\right]_{j,j}=\phi^{2}_{j}$.
We also define the forward and backward difference matrices in the $x$-direction for the interior nodes $i=1,\cdots,N_{x}-2$, where $i=0$ and $i=N_{x}-1$ correspond to zero Dirichlet boundary conditions and rest at the boundary:
$D_{+}$ $=\frac{1}{\Delta x}\begin{pmatrix}-1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\\ 0&\cdots&0&0&-1\end{pmatrix},$ (43)
$D_{-}$ $=\frac{1}{\Delta x}\begin{pmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\end{pmatrix}.$ (44)
Using these definitions, the discrete diffusion term in Eq. (42) can be represented in the matrix form as:
$\left(\Phi^{1}_{+}D_{+}-\Phi^{1}_{-}D_{-}\right)\mathbf{F}{\Phi_{2}^{\top}}.$ (45)
We note that the operator multiplying $\mathbf{F}$ from the left is analogous to $D^{\Phi}_{xx}$ in Equation (6). The discretization in the y-direction follows the same strategy.
Appendix B Discretization of the Advection Operator
We provide a detailed derivation of the matrix formulation of the discretization of the advection operator for the two-dimensional advection-diffusion equation with variable, separable advection coefficients. As before, we focus on the discretization of the advection term in the $x$-direction:
$\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right),$ (46)
where the advection coefficient is separable:
$\sigma^{x}(x,y)=\sigma^{1}(x)\sigma^{2}(y).$ (47)
Our goal is to discretize this term on a uniform grid with spacing $\Delta x$ in the $x$-direction and $\Delta y$ in the $y$-direction. Let $u_{i,j}\approx u(x_{i},y_{j})$. To preserve linearity of our discretization without introducing too much numerical dissipation, we employ a central difference scheme to approximate the spatial derivative. This discretization is not monotonicity-preserving unless the so-called cell Reynolds number condition is satisfied:
$\text{Re}_{c}=\frac{\sigma_{\text{max}}\Delta x}{\phi_{\text{max}}}<2,$ (48)
where $\sigma_{\text{max}}$ is the maximum magnitude of the advection velocity, and $\phi_{\text{max}}$ is the maximum diffusion coefficient. This condition is practical in our formulation due to the very favorable scaling of our low-rank representation with mesh size and rank.
The discrete approximation at grid point $(i,j)$ is:
$\left[\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right)\right]_{i,j}\approx\frac{1}{\Delta x}\left[\left(\sigma^{x}u\right)_{i+\frac{1}{2},j}-\left(\sigma^{x}u\right)_{i-\frac{1}{2},j}\right],$ (49)
where the fluxes at the cell faces are approximated by:
$\left(\sigma^{x}u\right)_{i+\frac{1}{2},j}$ $\approx\sigma^{1}_{i+\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i+1,j}+u_{i,j}}{2},$ (50)
$\left(\sigma^{x}u\right)_{i-\frac{1}{2},j}$ $\approx\sigma^{1}_{i-\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i,j}+u_{i-1,j}}{2}.$ (51)
Here, $\sigma^{1}_{i+\frac{1}{2}}=\sigma^{1}(x_{i+\frac{1}{2}})$ is the advection coefficient evaluated at the midpoint between $x_{i}$ and $x_{i+1}$, and similarly for $\sigma^{1}_{i-\frac{1}{2}}$. Substituting back into the discrete approximation, we have:
$\left[\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right)\right]_{i,j}$ $\approx\frac{1}{\Delta x}\left[\sigma^{1}_{i+\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i+1,j}+u_{i,j}}{2}-\sigma^{1}_{i-\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i,j}+u_{i-1,j}}{2}\right]$ (52)
$=\frac{1}{2\Delta x}\left[\sigma^{1}_{i+\frac{1}{2}}(u_{i+1,j}+u_{i,j})-\sigma^{1}_{i-\frac{1}{2}}(u_{i,j}+u_{i-1,j})\right]\sigma^{2}_{j}.$ (53)
To express the discretization in matrix form, we define:
  • $\mathbf{F}$: matrix of solution values, $[\mathbf{F}]_{i,j}=f_{i,j}$.
  • $\Sigma^{1}_{+}$ and $\Sigma^{1}_{-}$: diagonal matrices with elements $\left[\Sigma^{1}_{+}\right]_{i,i}=\sigma^{1}_{i+\frac{1}{2}}$ and $\left[\Sigma^{1}_{-}\right]_{i,i}=\sigma^{1}_{i-\frac{1}{2}}$.
  • $\Sigma_{2}$: diagonal matrix with elements $\left[\Sigma_{2}\right]_{j,j}=\sigma^{2}_{j}$.
We also define the difference matrices in the $x$-direction:
$S_{+}$ $=\frac{1}{2}\begin{pmatrix}1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\\ 0&\cdots&0&0&1\end{pmatrix},$ (54)
$S_{-}$ $=\frac{1}{2}\begin{pmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\end{pmatrix}.$ (55)
Using these definitions, the discrete advection term in the $x$-direction can be represented as:
$\frac{1}{\Delta x}\left(\Sigma^{1}_{+}S_{+}-\Sigma^{1}_{-}S_{-}\right)\mathbf{F}{\Sigma_{2}^{\top}}.$ (56)
We note that the operator multiplying $\mathbf{F}$ from the left is analogous to $D^{\Sigma}_{x}$ in Equation (6). The discretization in the y-direction follows the same strategy.
Appendix C High-order Temporal Discretization
We extend the backward Euler implicit integrator to higher-order DIRK time discretizations for the advection-diffusion operator, as compactly written in (8).
C.1 DIRK Scheme
DIRK methods are characterized by the Butcher tableau, as shown in Table 1.
Table 1: Butcher tableau for an $s$-stage DIRK scheme. Here, $s$ represents the number of stages in DIRK, $c_{i}$ represents the intermediate stage, $a_{ij}$ is a lower-triangular matrix with coefficients used to approximate the solution at each stage, and $b_{j}$ are the quadrature weights used to update the solution in the final step.
$c_{1}$ $a_{11}$ 0 $\ldots$ 0
$c_{2}$ $a_{21}$ $a_{22}$ldots 0
$\vdots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$
$c_{s}$ $a_{s1}$ $a_{s2}$ldots $a_{ss}$
$b_{1}$ $b_{2}$ldots $b_{s}$
We focus on stiffly accurate DIRK methods, where the coefficients satisfy $b_{i}=a_{si}$ for $i=1,\ldots,s$, ensuring that the final solution update corresponds to the solution at the last DIRK stage. For instance the Butcher tables for DIRK2 and DIRK3 are presented in Tables 3 and 3, respectively.
$\gamma$ $\gamma$ 0
$1$ $1-\gamma$ $\gamma$
$1-\gamma$ $\gamma$
Table 2: DIRK2 Butcher table with $\gamma=1-\frac{\sqrt{2}}{2}$.
$x$ $x$ 0 0
$\frac{1+x}{2}$ $\frac{1-x}{2}$ $x$ 0
$1$ $-\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $-\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
$-\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $-\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
Table 3: DIRK3 Butcher table with $x=0.4358665215$.
Similarly to the backward Euler case, the matrix-based discretization of the advection-diffusion equation using the operator $\mathscr{L}$ defined in (8) results in a generalized Sylvester equation that must be solved at each DIRK stage. The corresponding DIRK scheme for the matrix differential equation (8) from $t^{(0)}$ to $t^{(1)}$ is written as:
$\begin{split}\mathbf{F}^{(k)}&=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k}a_{k\ell}\mathbf{Y}_{\ell},\qquad k=1,2,\ldots,s,\end{split}$ (57a)
$\begin{split}\mathbf{Y}_{k}&=\mathscr{L}(\mathbf{F}^{(k)};t^{(k)}),\qquad t^{(k)}=t^{0}+c_{k}\Delta t,\qquad k=1,2,\ldots,s,\end{split}$ (57b)
$\begin{split}\mathbf{F}_{1}&=\mathbf{F}^{(s)}=\mathbf{F}_{0}+\Delta t\sum\limits_{k=1}^{s}b_{k}\mathbf{Y}_{k}.\end{split}$ (57c)
More compactly, within each stage $k$ of the DIRK scheme, one must minimize the residual equation (32) given by:
${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\quad k=1,2,\ldots,s.$
As shown in Equation (33), the residual equations lead to a generalized Sylvester equation to be solved at each stage.
C.2 Adaptive-Rank Strategy for High-Order Temporal Discretization via DIRK
[…]
$\mathbf{R}^{(k)}_{\tilde{\mathscr{L}}}=\mathbf{F}_{1}-a_{kk}\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell}.$
Expanding this residual using $\tilde{\mathscr{L}}$, defined in (14), leads to the $k$th-stage Sylvester equation:
$\underbrace{\left[\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\alpha_{2}T_{1}\right)}_{A_{1}}+\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\gamma_{2}T_{3}\right)}_{A_{3}}\right]\mathbf{F}^{(k)}}_{P_{1}}+\mathbf{F}^{(k)}\underbrace{\left[\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\alpha_{1}T_{2}\right)^{\top}}_{A^{\top}_{2}}+\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\gamma_{1}T_{4}\right)^{\top}}_{A^{\top}_{4}}\right]}_{P_{2}^{T}}=\mathbf{B}^{(k)}.$ (58)
Note that the operators $A_{1:4}$ are fixed throughout the stages since the DIRK method uses constant diagonal Butcher tableau coefficients.
We follow the strategy discussed in [10] to perform high-order integration. The key idea is to consider the basis fixed across DIRK stages, and only evolve the matrix of coefficients ${S}_{1}$ through the stages. The proposed procedure is as follows:
  1. Step 1.
    Prediction of Krylov basis functions: construct a set of orthonormal bases $U_{1}$ and $V_{1}$ from the Krylov-based low-rank implicit solver at the first DIRK stage, i.e. backward Euler with time stepping size $c_{1}\Delta t$.
  2. Step 2.
    For $k=1:s$ (per DIRK stage)
    1. (a)
      Solve reduced Sylvester equation for $S^{(k)}$ using preconditioned GMRES,
      ${\bf S}^{(k)}-\Delta t\left(\tilde{T}_{1}{\bf S}^{(k)}\tilde{\Phi}^{\top}_{2}+\tilde{\Phi}_{1}{\bf S}^{(k)}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}^{(k)}\tilde{\Sigma}^{\top}_{2}+\tilde{\Sigma}_{1}{\bf S}^{(k)}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}^{(k)},$ (59)
      with
      $\tilde{{\bf B}}^{(k)}=U_{1}^{\top}\mathbf{F}_{0}V_{1}+\Delta t\sum\limits_{\ell=1}^{k-1}{a_{k\ell}\tilde{Y_{\ell}}},$
      Here $\tilde{Y_{\ell}}=U_{1}^{\top}{Y}_{\ell}V_{1}\in\mathbb{R}^{r_{x}\times r_{y}}$. The preconditioner used is given by
      ${\bf S}\mapsto\tilde{P}_{1}{\bf S}+{\bf S}\tilde{P}_{2},$ (60)
      where $\tilde{P}_{1}=U_{1}^{T}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{T}P_{1}V_{1}$, and is valid for all stages since the basis constructed is fixed throught the DIRK stages by assumption. Solve for $S^{(k)}$ and obtain the intermediate RK solutions as $U_{1}S^{(k)}V_{1}^{\top}$. To further improve computational efficiency, from (59), we have,
      $\tilde{Y}_{\ell}$ $=\frac{1}{a_{\ell\ell}\Delta t}\left({S^{(\ell)}}-{\tilde{B}^{(\ell)}}\right),$
      leading to an efficient computation of $\tilde{B}^{(k)}$:
      $\tilde{B}^{(k)}=\tilde{B}_{1}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({S^{(\ell)}}-{\tilde{B}^{(\ell)}}\right),$ (61)
      with $\tilde{B}_{1}$ defined in (19). This avoids the need of evaluating the full size $B^{(k)}$ in (32).
  3. Step 3.
    Perform truncation of evolved solution, i.e. $\{U_{1},S_{1},V_{1}^{\top}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}^{\top}\})$, with $S_{1}:=S_{1}^{(s)}$.
  4. Step 4.
    Efficiently evaluate the residual norm,
    $\left\|{{\bf R}^{(s)}_{\mathscr{L}}}\right\|=\left\|{R}_{U}{\text{diag}\left(S_{1}-\tilde{B}^{(s)}_{1},-a_{ss}\Delta t\left(I_{{\mathcal{R}}_{\text{ranks}}}\otimes S_{1}\right)\right)}{R}_{V}^{\top}\right\|,$
    where $R_{U}$ and $R_{V}$ are upper triangular matrices from a reduced QR decomposition as described in (23), which can be done once for all DIRK stages.
  5. 1.
    Compare the computed residual norm $\|{\bf R}^{(s)}_{\mathscr{L}}\|$ with the given error tolerance $\epsilon_{tol}$. If the tolerance is met, then admit the solution with ${\bf F}_{1}=U_{1}S_{1}V_{1}^{\top}$; otherwise, we go back to Step 1 to further augment Krylov subspaces.
C.3 DIRK Generalization to Multi-Rank Advection-Diffusion Coefficients
We now extend the formulation to the multi-rank case, as described in (4). Using a DIRK method for temporal discretization, the $k$-th stage generalized Sylvester equation becomes:
$\mathbf{F}^{(k)}-a_{kk}\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}^{(k)}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}^{(k)}{T_{2,j}}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}^{(k)}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}^{(k)}{T_{4,l}}^{\top}}_{\mathscr{L}(\mathbf{F}^{(k)})}\right)=\mathbf{B}^{(k)},\quad k=1,2,\ldots,s$ (62)
where $\mathbf{B}^{(k)}$ is as defined in (32). The terms $T_{1:4,i}$ represent compositions of difference operators with diffusion and advection coefficients, similar to (7). Left subscripts 1 and 2 correspond to diffusion, while 3 and 4 correspond to advection in the $x$ and $y$ directions, respectively. The right subscript denotes the rank of the respective diffusion and advection coefficients. This equation generalizes (10), which is recovered for $\ell_{x}=\ell_{y}=k_{x}=k_{y}=1$ and equal diffusion and advection coefficients in both the $x$ and $y$ directions. It also generalizes the backward Euler special case with a single stage and $a_{1,1}=1$, yielding $\mathbf{B}^{(1)}=\mathbf{F}_{0}$, consistently with (10).
Using approximations in (25), we arrive at the $k$th stage approximated Sylvester equation:
[…]
where
$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{3,k}}=\sum_{i=1}^{\ell_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-a_{kk}\Delta t\alpha_{x,i}T_{1,i}\right)}+\sum_{k=1}^{k_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-a_{kk}\Delta t\gamma_{x,k}T_{3,k}\right)},$
$P_{2}$ $=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{4,l}}=\sum_{j=1}^{\ell_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-a_{kk}\Delta t\alpha_{y,j}T_{2,j}\right)}+\sum_{l=1}^{k_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-a_{kk}\Delta t\gamma_{y,l}T_{4,l}\right)}.$
In order to obtain a reduced equation for the $k$th stage matrix of coefficients $S^{(k)}$, we perform a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$ to obtain :
${\bf S}^{(k)}-a_{kk}\Delta_{t}\left(\sum_{i=1}^{\ell_{x}}\tilde{T}_{1,i}{\bf S}^{(k)}\tilde{\Phi}_{i}^{2,x\top}+\sum_{j=1}^{\ell_{y}}\tilde{\Phi}_{j}^{1,y}{\bf S}^{(k)}\tilde{T}_{2,j}^{\top}+\sum_{k=1}^{k_{x}}\tilde{T}_{3,k}{\bf S}^{(k)}\tilde{\Sigma}_{k}^{2,x\top}+\sum_{l=1}^{k_{y}}\tilde{\Sigma}_{l}^{1,y}{\bf S}^{(k)}\tilde{T}_{4,l}^{\top}\right)=\tilde{{\bf B}}^{(k)}$ (64)
where the projected operators are analogous to those defined in the rank-one case.
[…]
$\underbrace{\left[\sum_{i=1}^{\ell_{x}}{\tilde{A}_{1,i}}+\sum_{k=1}^{k_{x}}{\tilde{A}_{3,k}}\right]}_{\tilde{P}_{1}}\mathbf{S}^{(k)}+\mathbf{S}^{(k)}\underbrace{\left[\sum_{j=1}^{\ell_{y}}{\tilde{A}_{2,j}}+\sum_{l=1}^{k_{y}}{\tilde{A}_{4,l}}\right]}_{\tilde{P}_{2}^{\top}}=\tilde{{\bf B}}^{(k)}$ (65)
Using the operators $\tilde{P}_{1}$ and $\tilde{P}_{2}$, we define the preconditioner mapping:
${\bf S}\mapsto\tilde{P}_{1}{\bf S}+{\bf S}\tilde{P}_{2}^{T},$ (66)
Algorithm 4 solves the variable advection diffusion equation using DIRK methods.
Input: Initial condition matrices $U_{0}$, $V_{0}$, $S_{0}$;
Operators $P_{1}$, $P_{2}$, $\{T_{1,i}\}_{i=1}^{\ell_{x}}$, $\{T_{2,j}\}_{j=1}^{\ell_{y}}$, $\{T_{3,k}\}_{k=1}^{k_{x}}$, $\{T_{4,l}\}_{l=1}^{k_{y}}$;
Butcher table $\{a_{ij}\}$;
Time step size $\Delta t$;
Tolerances $\epsilon_{\kappa}$, $\epsilon$, ${\epsilon}_{\text{tol}}$;
Maximum iterations max_iter;
Output: Updated bases $U_{1}$, $V_{1}$;
Truncated singular values $S_{1}$;
1 Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eq. (63);
2
3for $m=1$ to max_iter do
4        // Step 1.
5        Truncate to tolerance $\epsilon_{\kappa}$;
6        $U_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right)\right)$;
7        $V_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right)\right)$;
8        Compute $\tilde{B}^{(1)}=(U_{1}^{\top}U_{0})S_{0}(V_{1}^{\top}V_{0})^{\top}$;
9        // Step 2.
10        for $k=1$ to $s$ do
11               Compute $\tilde{B}^{(k)}=\tilde{B}^{(1)}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left(S^{(\ell)}-\tilde{B}^{(\ell)}\right)$;
12               Solve the reduced equation (65) for $S^{(k)}$ using GMRES with ACS preconditioner (66) to tolerance $\epsilon_{\texttt{GMRES}}$;
13               Compute and store $\frac{1}{a_{kk}}\left(S^{(k)}-\tilde{B}^{(k)}\right)$ and proceed to the next stage;
14
15       Truncate $\{U_{1},S_{1},V_{1}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S^{(s)},V_{1}\})$ to tolerance $\epsilon$;
16        // Step 3.
17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},T_{1,1:\ell_{x}}U_{1},\Phi_{1:\ell_{y}}^{1,y}U_{1},T_{3,1:k_{x}}U_{1},\Sigma_{1:k_{y}}^{1,y}U_{1}]\right)$ and
18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},\Phi_{1:\ell_{x}}^{2,x}V_{1},T_{2,1:\ell_{y}}V_{1},\Sigma_{1:k_{x}}^{2,x}V_{1},T_{4,1:k_{y}}V_{1}]\right)$;
19        // Step 4.
20        Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(S_{1}-\tilde{B}^{(s)},-a_{ss}\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1}\right)\right)R_{V}^{\top}\right\|$;
21
22       if $\|\mathbf{R}\|/\|\mathbf{F}_{0}\|\geq\epsilon_{\texttt{tol}}$ then
23               Return to Step 1 to augment bases further;
24
25       else
26               Break;
27               Exit loop;
28
29
30
Algorithm 4 s-Stages Adaptive-Rank DIRK Integrator for the Generalized Sylvester Equation. 00footnotetext: The MATLAB implementation of the algorithm is publicly available on GitHub, Adaptive-rank-ACS, https://github.com/helkahza1/Adaptive-rank-ACS

### v2

@@
-Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Inhomogeneous Coefficients
+Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Variable Coefficients
 Abstract
-We consider the adaptive-rank integration of general time-dependent advection-diffusion partial differential equations (PDEs) with spatially variable coefficients.
-We employ a standard finite-difference method for spatial discretization coupled with diagonally implicit Runge-Kutta schemes for temporal discretization.
-The fully discretized scheme can be written as a generalized Sylvester equation, which we solve with an adaptive-rank algorithm structured around three key strategies: (i) constructing dimension-wise subspaces based on an extended Krylov strategy, (ii) developing an effective Averaged-Coefficient Sylvester (ACS) preconditioner to invert the reduced system for the coefficient matrix efficiently, and (iii) efficiently computing the residual of the equation without explicitly reverting to the full-rank form.
-The proposed approach features a computational complexity of $\mathcal{O}(N{r^{2}}+{r^{3}})$, where $r$ represents the rank during the Krylov iteration and $N$ is the resolution in one dimension, commensurate with the constant-coefficient case [10].
+We consider the adaptive-rank integration of 2D and 3D time-dependent advection-diffusion partial differential equations (PDEs) with variable coefficients.
+We employ a standard finite-difference method for spatial discretization coupled with diagonally implicit Runge-Kutta temporal schemes.
+The discrete equation is a generalized Sylvester equation (GSE), which we solve with an adaptive-rank algorithm structured around three key strategies: (i) constructing dimension-wise subspaces based on an extended Krylov strategy, (ii) developing an effective preconditioner for the reduced coefficient matrix, and (iii) efficiently computing the residual of the equation without explicitly reverting to the full-rank form.
+The low-rank decomposition is performed in 2D using SVD, and with high-order SVD (HOSVD) in 3D to represent the tensor in a compressed Tucker format.
+The computational complexity of the proposed approach is demonstrated numerically to be comparable to the constant-coefficient case [14], scaling as $\mathcal{O}(N{r^{2}}+{r^{d+1}})$ for $d$-dimensional problems (here, $d=2$ or $3$), with $N$ the resolution in one dimension and $r$ the maximal rank during the Krylov iteration (which we find to be largely independent of $N$).
 We present numerical examples that illustrate the computational efficacy and complexity of our algorithm.
-Keywords: Adaptive-rank, extended Krylov subspaces, implicit Runge-Kutta integrators, advection-diffusion equations, Sylvester preconditioning, matrix-equations.
-1 Introduction
-We propose an adaptive-rank numerical scheme for approximating the solution of the time-dependent variable-coefficient advection-diffusion equation.
-A challenge in existing numerical schemes is the curse of dimensionality, where the problem dimension $d$ leads to exponential growth in computational complexity.
-These formulations often result in large matrix systems of the form $Ax=b$, with $A$ sparse and banded, of size $N^{d}$-by-$N^{d}$ and bandwidth $\mathcal{O}(N^{d-1}$), where $N$ is the resolution in one dimension and $d$ is the dimensionality.
-Direct solvers for such systems scale in general as $\mathcal{O}(N^{3d-2})$ [3].
-The complexity of unpreconditioned Conjugate Gradient (a type of Krylov method) for diffusion problems scales as $\mathcal{O}(N^{3d/2})$ [6, 19].
-Multigrid-preconditioned Krylov methods may improve this scaling to $\mathcal{O}(N^{d})$ [16], which is considered optimal.
-In all these cases, the presence of the dimension $d$ in the exponent makes computation of the inverse matrix challenging for fine-mesh resolutions of high-dimensional problems.
-Our method seeks to overcome this dimensional scaling and provide an efficient approach with a complexity of $\mathcal{O}(N)$ to solving time-dependent variable coefficient advection-diffusion partial differential equations (PDEs).
-In two dimensions, our approach leverages adaptive-rank approximations of matrix solutions on a tensor-product grid.
-Adaptive-rank techniques have proven effective in reducing the computational complexity of high-dimensional problems by representing the solution matrix as a sum of the outer products of basis vectors.
-Specifically, we employ a truncated singular value decomposition (SVD), such that $X=USV^{\top}$, where basis vectors in $U$ and $V$ span the solution space and determine the rank of the solution.
-The key idea is to dynamically update the basis vectors, as well as the rank (denoted as $r$), to maintain accuracy while minimizing computational cost.
-Two primary strategies in the literature exploit low-rank structures for time-dependent PDEs: the dynamical low-rank (DLR) approximation [17, 22, 5, 8, 21] and the step-and-truncate (SAT) method, which has both explicit [18, 7, 13, 12, 14] and implicit [29, 24, 10, 23] formulations.
-Our work aligns with the SAT framework, where we perform temporal discretization with high-order accuracy and fully implicitly, leading to matrix differential equations to be solved at every time step.
-For constant coefficients, this system becomes a Sylvester equation of the form:
+Keywords: Adaptive-rank, extended Krylov subspaces, implicit Runge-Kutta integrators, advection-diffusion equations, Sylvester preconditioning, matrix equations.
+1 Introduction:
+We propose an adaptive-rank numerical scheme for approximating solutions of time-dependent variable-coefficient advection-diffusion equations.
+The goal is to reduce the computational complexity of high-dimensional PDE solvers by exploiting the low-rank structure of the solution, building upon existing literature on low-rank PDE solvers [16, 13] and model order reduction (MOR) for solving high-dimensional linear and nonlinear systems [24, 16, 5, 40].
+Two primary strategies in the literature exploit evolving low-rank structures for time-dependent PDEs: the dynamical low-rank (DLR) approximation [21, 27, 8, 11, 26] and the step-and-truncate (SAT) method, which has both explicit [23, 10, 18, 17, 19] and implicit [35, 30, 14, 28, 29] variants.
+Our study employs SAT, discretizing implicitly in time with high-order accuracy, leading to matrix/tensor differential equations to be solved at every time step.
+For constant coefficients in 2D, the advection-diffusion PDE becomes a Sylvester equation (SE) of the form:
 $A\mathbf{X}_{1}+\mathbf{X}_{1}B^{\top}=\mathbf{X}_{0},$ (1)
 where $\mathbf{X}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ represents the solution at time $t_{1}$, $\mathbf{X}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$ is the initial condition at $t_{0}$, and $A$ and $B$ are difference operators arising from the PDE discretization.
-Krylov subspace methods have proven effective for solving Sylvester-type equations resulting from the discretization of PDEs [20, 33, 29, 25, 2].
-In these methods, the solution bases are chosen from a Krylov (or Krylov-like) subspace of $A$ and $B$ in (1), and the original system is projected onto this reduced subspace using the Galerkin projection.
-To accelerate the convergence of iterative methods, various extensions of Krylov subspaces have been developed.
-Notably, the extended Krylov subspace [9] involving $A^{-1}$, $B^{-1}$ has been applied to matrix equations in [33] and later adopted for time-dependent PDEs in [10].
-While substantial progress has been made in constant-coefficient problems, systems with variable coefficients remain challenging.
-Discretizing such systems leads to the so-called generalized Sylvester matrix equation, of the form:
+Here, $N_{1}$ and $N_{2}$ are the 2D grid dimensions.
+Krylov subspace methods have proven effective for solving SEs resulting from the discretization of PDEs [24, 39, 35, 31, 6].
+In these methods, the solution bases are chosen from a Krylov (or Krylov-like) subspace of $A$ and $B$ in (1), and the original system is Galerkin-projected onto this reduced subspace.
+To accelerate the convergence of iterative methods, extensions of Krylov subspaces have been developed [12] involving $A^{-1}$, $B^{-1}$, which have been successfully applied to matrix equations [39, 14].
+While substantial progress has been made for constant-coefficient problems, PDEs with variable coefficients remain challenging.
+Discretizing such systems leads to the so-called generalized Sylvester equation (GSE), of the form:
 $\sum_{i=1}^{l}A_{i}\mathbf{X}_{1}{B}^{\top}_{i}=\mathbf{X}_{0},$ (2)
 where $A_{i}$ and $B_{i}$ represent discretization operators for advection and/or diffusion terms.
-While these types of matrix equations can be solved by reverting them to the Kronecker form, in [26] the authors argue that, for a chosen approximation space, the Galerkin projection applied to the matrix equation consistently achieves faster convergence than the Kronecker formulation, particularly when the operators are poorly conditioned.
-Both analytical and numerical evidence in [26] support this claim.
-Furthermore, they show that the optimality of the Galerkin projection technique for Sylvester equations extends naturally to the generalized Sylvester equations.
-That is, regardless of the approximation space, the residual decreases as the subspaces are iteratively expanded, though the convergence rate depends significantly on the choice of subspaces, as demonstrated in [10].
-Additional validation of this claim is provided in [34], where a matrix-based conjugate gradient method is compared with Galerkin projection techniques.
-Such results encourage the community to adopt matrix-equation-based strategies for multi-term matrix equations, rather than defaulting to Kronecker formulations.
-Several methods have been proposed to solve generalized Sylvester equations iteratively directly in matrix form, either with stationary iterative techniques or Krylov methods.
-Stationary iterative methods [1, 32] leverage low-rank techniques to enhance efficiency, achieving computational complexity as low as $\mathcal{O}(N)$ when implemented effectively [32].
-However, their effectiveness strongly depends on whether or not such a dominating operator could be correctly identified to enable a contractive iterative process.
-The difficulty of explicitly finding such a dominating term in a general setting was discussed in [25], which proposed using the Kronecker formulation instead.
-Krylov methods have also been explored as an alternative for the solution of equations of the Sylvester and generalized Sylvester type.
-In [1], the authors present a series of techniques to tackle the Generalized Lyapunov equation, which is a special case of (2) where the first two terms are of Sylvester type, i.e.
-$A_{1}=B_{2}:=A$, $B_{1}=I_{N_{2}}$, and $A_{2}=I_{N_{1}}$.
-Among these techniques, they discuss a low-rank projection-based scheme, in which they propose constructing extended Krylov subspaces using $A$, following the approach of [33], and employing a standard Krylov basis for the remaining operators.
-In [27], a low-rank projection-based technique is employed for systems where $N_{2}\ll N_{1}$.
-The authors opt for the rational Krylov as an approximation space for the Galerkin projection of the original equation, and solve the resulting small-sized equation with either matrix-based Conjugate Gradient or a Kronecker formulation, given that $N_{2}$ is relatively small.
-In [15], a Krylov-based projection technique is developed for matrix equations of the form (2), where the Krylov basis construction relies on a low-rank decomposition of the operators.
-The resulting small generalized Sylvester matrix system is solved using the Sherman-Morrison-Woodbury formula through a series of Sylvester equation solves, which leverages the operator’s low-rank structure.
-Finally, of particular relevance to our study, a basis-update-and-Galerkin (BUG) DLR integrator is proposed in [23] as a low-rank GMRES preconditioner for the generalized Sylvester equation arising from advection-diffusion time-dependent problems with variable coefficients.
-The approach evolves the basis and the coefficient matrix in the preconditioner using BUG to provide guesses for the basis vectors and coefficient updates.
-The advantage of the approach is that the approximations involved in the BUG integrator are only present in the preconditioning stage, and not in the converged solution.
-In contrast to existing efforts, in this work we propose an optimal implicit time integrator based on SAT adaptive-rank integrators for the variable-coefficient advection-diffusion equation.
-Instead of evolving the bases and coefficients simultaneously, we evolve them separately to produce an effective, inexpensive implicit solver.
-In particular, we iteratively construct a set of extended Krylov subspaces (for each dimension) that effectively capture the solution’s range evolution in an iterative manner.
-A Galerkin projection is then applied to project the full system of equations into a reduced system.
-The reduced system is solved using Krylov methods like GMRES, with preconditioning based on an approximated operator that transforms the generalized Sylvester equation into a regular Sylvester one, for which suitable efficient solvers exist.
-To assess the goodness of the solution efficiently, we propose a residual-checking mechanism that leverages the low-rank structure of solutions directly.
-Such procedure avoids the need to form the solution matrix for norm evaluation.
-The overall scheme scales as $\mathcal{O}(Nr^{2}+r^{3})$, with $r$ the rank.
-The main innovations of our proposed algorithm are as follows:
-  • Linear “1D-like” Computational Complexity: We introduce a Krylov-based adaptive-rank algorithm for solving equations of the form (2), which arise from implicit integrators for advection-diffusion-type equations.
-Our method achieves linear computational complexity with respect to $N$ and quadratic complexity with respect to $r$, offering a significant advantage over existing state-of-the-art adaptive-rank approaches.
-To the best of our knowledge, this is the first study that clearly demonstrates the optimal complexity of $\mathcal{O}(Nr^{2})$ for the advection-diffusion-type equation with variable coefficients.
-  • Generalized dimension-wise extended Krylov Subspaces: We propose an averaged-coefficient approximation operator $\tilde{\mathscr{L}}$, which is instrumental not only in building dimension-wise extended Krylov subspaces, but also for preconditioning the reduced projected system.
-The dimension-wise Krylov subspaces are generated from multiple operators present in ${\mathscr{L}}$, $\tilde{\mathscr{L}}$ and their inverses.
-Such strategy is numerically demonstrated in this study to be computationally effective with rapid convergence in Krylov iterations.
-  • Efficient Preconditioning Strategy: We deploy a GMRES solver for the reduced generalized Sylvester equation from Galerkin projection (of size $r^{2}$), with an Averaging-Coefficient Sylvester (ACS) preconditioner to accelerate the Krylov convergence.
-The proposed preconditioner is based on a Sylvester direct solver for the averaged-coefficient approximation operator $\tilde{\mathscr{L}}$, which results in a computational complexity of $\mathcal{O}(r^{3})$ (vs.
-the expected $\mathcal{O}(r^{6})$ of a direct inversion of the Kronecker-product form of the reduced generalized Sylvester equation).
-This preconditioner is inspired by the recent work in [25], which employed a similar approach for the full-matrix form of the generalized Sylvester equation.
-A significant enhancement of the algorithm’s efficiency is demonstrated in our numerical tests.
-  • Optimized solution representation and low rank residual computation: We employ an SVD-truncated QR procedure for basis augmentation, and an SVD-truncated procedure for the intermediate solution during the Krylov iteration as often as possible.
-This optimizes the computational efficiency, while preserving the accuracy and structure of the solution.
-In addition, an efficient low-rank computation of the residual is enabled with $\mathcal{O}(r^{3})$ complexity without explicitly forming the full solution on tensor-product grids.
-This avoids the computational cost associated with full matrix, which could otherwise spoil the linear computational complexity sought.
-The structure of the rest of this paper is as follows.
-In Section 2, we consider the case of rank-one advection and diffusion coefficients and derive the generalized Sylvester equation that results from a finite-difference discretization in space and backward Euler in time.
-In Section 3, we present our adaptive-rank solution strategy, including the coefficient-averaged approximation operator, construction of extended-Krylov subspaces, preconditioning technique in solving the reduced system, and low-rank computation of the residual norm.
-The extension of the technique to high-order temporal discretizations with multi-rank coefficients is provided in Appendix C.
-In Section 4, numerical results are presented to demonstrate the efficacy in high-order temporal convergence, adaptive rank evolution of Krylov subspaces, optimal computational complexity, as well as accelerated Krylov convergence.
-Finally, we conclude with a brief highlight of our proposed algorithm and outlook for future research in Section 5.
+Several methods have been proposed to solve GSEs either with stationary iterative techniques  [5, 38] or Krylov methods  [5, 39, 32, 42].
+Stationary iterative methods leverage low-rank techniques to enhance efficiency, achieving computational complexity of $\mathcal{O}(N)$ when implemented effectively [38].
+However, their effectiveness critically depends on whether a dominating operator can be correctly identified to enable a contractive iterative process.
+For the Krylov approach, recent approaches leverage a Galerkin projection of the matrix equation on the dimension-wise Krylov bases when the right-hand side is low-rank [36, 40].
+Compared to a standard Kronecker Krylov iteration, dimension-wise Krylov subspaces are inherently richer and yield superior approximations [32, 42, 24].
+Several studies in the literature have considered using atypical extended Krylov subspaces to solve GSEs (see [5, 33, 20] and references therein).
+In [5], the authors present a series of techniques to tackle the Generalized Lyapunov equation, which is a special case of (2) of the form:
+$A\mathbf{X}+\mathbf{X}A^{T}+\sum_{j=1}^{m}M_{j}\mathbf{X}M_{j}^{T}=\mathbf{F}.$ (3)
+Among them, they discuss a low-rank projection-based scheme, in which extended Krylov subspaces are constructed using $A$ and standard Krylov subspaces are used for the remaining operators.
+In [20], a Krylov-based projection technique is developed for Generalized Lyapunov equations with $m=1$.
+The method builds its approximation subspaces by block augmentations generated from $A$ and $M_{1}$.
+In [33], a low-rank projection-based technique is employed for systems of disparate dimension sizes, $N_{2}\ll N_{1}$.
+The authors employ a nested sequence of block rational Krylov subspaces generated from all the operators in the $N_{1}$ direction, and solve the resulting small-sized equation with either matrix-based Conjugate Gradient or a Kronecker formulation (tractable since $N_{2}$ is small).
+In all of these studies, atypical approximation subspaces are formed by augmenting the basis with multiple applications of the operators in the multi‑term matrix equation, a strategy we adopt in this work.
+Another Krylov-based line of inquiry, particularly relevant to our study, employs a basis-update-and-Galerkin (BUG) DLR integrator as a low-rank GMRES preconditioner for the GSE arising from anisotropic diffusion time-dependent problems with variable coefficients [28].
+The BUG preconditioner evolves the basis and the coefficient matrix to provide guesses for the basis vectors and coefficient updates of the outer GMRES iteration, demonstrating significant potential in controlling GMRES iteration count and maximal Krylov rank.
+The approach is compatible with high-order temporal integrators.
+However, the reference offers no discussion of asymptotic computational complexity.
+In this study, we propose to solve the GSE for time-dependent advection-diffusion equations leveraging adaptive-rank approximations of solutions on tensor-product grids with $N$ grid points per dimension, and using extended Krylov subspaces to update the basis vectors as well as the rank (denoted as $r$).
+Our main goal is to minimize the computational cost of the adaptive rank implementation without sacrificing accuracy.
+Key innovations of our proposed algorithm include:
+  • Generalized Extended Krylov (xKrylov) Adaptive-Rank Integrators with Complexity Reduction for 2D and 3D Problems.
+We consider matrix solutions with factorizations of the form $X=USV^{\top}$, where basis vectors in $U$ and $V$ span the solution space and determine the rank of the solution.
+In the 2D setting, we employ a truncated singular value decomposition (SVD) to factorize the solution.
+In 3D, we use high-order singular value decomposition (HOSVD) to represent the 3D tensor in a compressed Tucker format.
+Our proposed approach constructs a set of xKrylov subspaces (for each dimension) that effectively capture the evolution of solution basis in an iterative manner.
+A Galerkin projection is then applied to project the full system to a reduced GSE.
+We propose a residual-checking mechanism that leverages the low-rank structure of solutions directly, thus avoiding forming the full residual matrix for norm evaluations.
+For the rest of the paper, we refer to the xKrylov iteration as the outer iteration.
+  • Effective Preconditioning Strategy: The effectiveness of these projection-based approaches depends on the availability of an effective solver for the reduced GSE (2) of size up to $r$, the maximum size of dimension-wise xKrylov subspaces.
+There are two main challenges to solve it: (i) an increased computational load due to the multi-term nature of the equation, leading to rapid increase of xKrylov subspace sizes, and (ii) slow convergence of iterative methods (e.g., GMRES) resulting from algebraic ill-conditioning when solved iteratively.
+These challenges highlight the need for an effective preconditioner.
+Inspired by [31], we propose an averaged-coefficient Sylvester (ACS) approximation operator to the GSE as a preconditioner to GMRES.
+The approximate operator is of the Sylvester form and compact-equivalent to the original GSE, ensuring excellent performance as a GMRES accelerator [1, 2].
+The direct inversion of the SE features a favorable computational complexity of $\mathcal{O}(r^{d+1})$, to be contrasted with the $\mathcal{O}(r^{3d})$ for $d$-dimensional problems when directly inverting the Kronecker-product form of the GSE.
+We refer to the preconditioned-GMRES reduced-GSE solve as the inner iteration.
+  • Efficient Computational Complexity: As proposed, our method achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$ for $d$-dimensional problems, offering a significant advantage over existing state-of-the-art adaptive-rank approaches.
+The $Nr^{2}$ term corresponds to the outer xKrylov iteration, and the $r^{d+1}$ term to the inner GMRES iteration, assuming an $r$-independent GMRES convergence rate (which we achieve with the ACS preconditioner).
+In principle, the rank $r$ may scale with $N$ [24], but we do not find this in practice.
+The favorable scaling properties of our approach will be verified in our numerical experiments.
+The rest of this paper is organized as follows.
+In Section 2, we consider the case of rank-one advection and diffusion coefficients and derive the GSE that results from a finite-difference discretization in space and backward Euler in time.
+In Section 3, we introduce an adaptive-rank solution strategy for 2D problems, including the coefficient-averaged approximation operator, construction of xKrylov subspaces, preconditioning technique in solving the reduced system, and low-rank computation of the residual norm.
+The extension of the technique to high-order temporal discretizations with multi-rank coefficients is provided in Appendix C, and for 3D problems in Appendix D.
+In Section 4, numerical results are presented to demonstrate the efficacy of the proposed algorithms.
+Finally, we conclude and provide outlook for future research in Section 5.
 2 Problem Description and Generalized Sylvester Equation
-In what follows, we refer to equations in the form of Equation (1) as the Sylvester equation, where the unknown solution matrices are operated on one dimension at a time.
-Additionally, we use the term generalized Sylvester equation to describe equations in the form of Equation (2), where each unknown solution matrix is simultaneously operated on across both dimensions.
-We establish the following notation conventions based on the nature of the symbols.
+For simplicity, we focus this section on 2D, with the extension to 3D elaborated in Appendix D.
+We establish the following notation conventions.
 Lowercase Greek letters, such as $\phi(x,y)$ or $\sigma(x,y)$, denote continuous scalar fields.
-Uppercase bold italic Greek letters, such as $\bm{\mathit{\Phi}}(x,y)$ or $\bm{\mathit{\Sigma}}(x,y)$, are used for vector-valued and tensor-valued functions.
+Uppercase bold italic Greek letters, such as $\bm{\mathit{\Sigma}}(x,y)$ or $\bm{\mathit{\Phi}}(x,y)$, are used for vector-valued and tensor-valued functions.
 Uppercase Greek letters, like $\Phi$ or $\Sigma$, represent discrete matrices arising from the discretization of continuous scalar functions or differential operators on a grid.
@@
 The identity matrix of size $n$ is denoted by ${I}_{n}$.
-We aim to solve the advection-diffusion equation:
-$\frac{\partial f}{\partial t}=\nabla_{v}\cdot(\bm{\mathit{\Phi}}\cdot\nabla_{v}f)-\nabla_{v}\cdot(\bm{\mathit{\Sigma}}f),$ (3)
-where $\bm{\mathit{\Phi}}\in\mathbb{R}^{2\times 2}$ is an isotropic diffusion tensor:
+We aim to solve the time-dependent advection-diffusion equation:
+$\frac{\partial f}{\partial t}=\nabla_{v}\cdot(\bm{\mathit{\Phi}}\cdot\nabla f)-\nabla\cdot(\bm{\mathit{\Sigma}}f),$ (4)
+where $(x,y)\in\Omega=[a_{1},b_{1}]\times[a_{2},b_{2}]$ and $\bm{\mathit{\Phi}}\in\mathbb{R}^{2\times 2}$ is an isotropic diffusion tensor:
 $\bm{\mathit{\Phi}}(x,y)=\begin{bmatrix}\phi^{x}(x,y)&0\\ 0&\phi^{y}(x,y)\end{bmatrix},$
-and the advection vector $\bm{\mathit{\Sigma}}\in\mathbb{R}^{2}$ is given by:
+and $\bm{\mathit{\Sigma}}\in\mathbb{R}^{2}$ is the advection vector, given by:
 $\bm{\mathit{\Sigma}}(x,y)=\begin{bmatrix}\sigma^{x}(x,y)\\ \sigma^{y}(x,y)\end{bmatrix}.$
-We assume that the diffusion and advection coefficients are separable and they exhibit a low-rank structure.
-Specifically, the diagonal diffusion coefficients, $\phi^{x}(x,y)$ and $\phi^{y}(x,y)$, are given by:
-$\phi^{x}(x,y)=\sum_{i=1}^{\ell_{x}}\phi^{1,x}_{i}(x)\phi^{2,x}_{i}(y),\quad\phi^{y}(x,y)=\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\phi^{2,y}_{i}(y).$
+We assume that the diffusion and advection coefficients are separable with finite rank. Specifically, the diagonal diffusion coefficients, $\phi^{x}(x,y)$ and $\phi^{y}(x,y)$, are given by:
+$\phi^{x}(x,y)=\sum_{i=1}^{\ell_{x}}\phi^{1,x}_{i}(x)\phi^{2,x}_{i}(y),\quad\phi^{y}(x,y)=\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\phi^{2,y}_{i}(y).$ (5)
 Similarly, for the advection coefficients we have:
-$\sigma^{x}(x,y)=\sum_{i=1}^{k_{x}}\sigma^{1,x}_{i}(x)\sigma^{2,x}_{i}(y),\quad\sigma^{y}(x,y)=\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\sigma^{2,y}_{i}(y).$
+$\sigma^{x}(x,y)=\sum_{i=1}^{k_{x}}\sigma^{1,x}_{i}(x)\sigma^{2,x}_{i}(y),\quad\sigma^{y}(x,y)=\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\sigma^{2,y}_{i}(y).$ (6)
 Substituting these coefficients in their separable forms into Equation (4), we obtain:
@@
 $-\sum_{i=1}^{k_{x}}\frac{\partial}{\partial x}\left(\sigma^{1,x}_{i}(x)f(x,y)\right)\sigma^{2,x}_{i}(y)-\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\frac{\partial}{\partial y}\left(\sigma^{2,y}_{i}(y)f(x,y)\right),$
-where, in each term, the component independent of the differentiation variable has been pulled out of the differentiation operator.
-Discretizing this equation using finite differences in space and an implicit method in time naturally leads to a multi-term matrix differential equation where the unknown solution is acted upon from both directions in each term and at the same time.
-Numerous strategies have been proposed in the literature to solve such equations, most notably the recent work by Palitta and Simoncini in [25].
-In this work, inspired by their preconditioning technique [25], we further exploit the low-rank structure of the solution and design a projection-based algorithm with an optimal scaling of $\mathcal{O}(Nr^{2})$, i.e.
-a linear scaling with respect to the resolution $N$ per spatial dimension, and quadratic scaling with respect to the rank $r$.
-2.1 Generalized Sylvester Equation from Implicit Discretization of Advection-Diffusion
-To simplify the presentation, we assume rank-one advection and diffusion coefficients, i.e.:
-$\phi^{x}=\phi^{1,x}(x)\phi^{2,x}(y),\quad\phi^{y}=\phi^{1,y}(x)\phi^{2,y}(y)^{\top},$
+where, in each term, the factor independent of the differentiation variable has been pulled out of the derivative.
+To simplify the presentation, we assume for now rank-one advection and diffusion coefficients:
+$\phi^{x}=\phi^{1,x}(x)\phi^{2,x}(y),\quad\phi^{y}=\phi^{1,y}(x)\phi^{2,y}(y),$
 $\sigma^{x}=\sigma^{1,x}(x)\sigma^{2,x}(y),\quad\sigma^{y}=\sigma^{1,y}(x)\sigma^{2,y}(y).$
-Thus, Equation (4) simplifies to:
-$\frac{\partial f}{\partial t}$ $=\underbrace{\frac{\partial}{\partial x}\left(\phi^{1,x}(x)\frac{\partial f(x,y)}{\partial x}\right)\phi^{2,x}(y)+\phi^{1,y}(x)\frac{\partial}{\partial y}\left(\phi^{2,y}(y)\frac{\partial f(x,y)}{\partial y}\right)}_{\text{diffusion}}$ (5)
-${-}\underbrace{\frac{\partial}{\partial x}\left(\sigma^{1,x}(x)f(x,y)\right)\sigma^{2,x}(y){-}\sigma^{1,y}(x)\frac{\partial}{\partial y}\left(\sigma^{2,y}(y)f(x,y)\right)}_{\text{advection}}.$
-We discretize this equation in a two-dimensional tensor-product spatial grid with ${N}_{1}$ and ${N}_{2}$ grid points in the $x$- and $y$- directions, respectively.
-The grid points in the $x$- direction are denoted as $x_{i}$ for $i=0,\dots,N_{1}-1$, and similarly for the $y$-direction.
+This will be generalized to multiple ranks below. Equation (7) then simplifies to:
+$\frac{\partial f}{\partial t}$ $={\frac{\partial}{\partial x}\left(\phi^{1,x}(x)\frac{\partial f(x,y)}{\partial x}\right)\phi^{2,x}(y)+\phi^{1,y}(x)\frac{\partial}{\partial y}\left(\phi^{2,y}(y)\frac{\partial f(x,y)}{\partial y}\right)}$ (8)
+${-}{\frac{\partial}{\partial x}\left(\sigma^{1,x}(x)f(x,y)\right)\sigma^{2,x}(y){-}\sigma^{1,y}(x)\frac{\partial}{\partial y}\left(\sigma^{2,y}(y)f(x,y)\right)}.$
+We discretize this equation in a two-dimensional tensor-product spatial grid with ${N}_{1}$ and ${N}_{2}$ grid points in the $x$ and $y$ directions, respectively.
+The grid points in the $x$ direction are denoted as $x_{i}$ for $i=0,\dots,N_{1}-1$, and similarly for the $y$ direction.
 We consider the point-wise multiplication of the advection and diffusion coefficients in the discrete by setting:
@@
 $\Phi_{2}$ $=\text{diag}(\phi^{2,x}(y_{j})),\Sigma_{2}=\text{diag}(\sigma^{2,x}(y_{j})),\quad j=0,\cdots,N_{2}-1.$
-We approximate the spatial derivative with a three-point stencil, and obtain the following matrix differential equation for $\mathbf{F}$, for which its $(i,j)$ entry $F_{i,j}$ corresponds to the approximation of $f(x_{i},y_{j})$:
-$\frac{\partial\mathbf{F}}{\partial t}=\underbrace{D_{xx}^{\Phi}\mathbf{F}\Phi_{2}^{\top}+\Phi_{1}\mathbf{F}{D_{yy}^{\Phi}}^{\top}}_{\text{\emph{diffusion}}}-\underbrace{D_{x}^{\Sigma}\mathbf{F}\Sigma_{2}^{\top}-\Sigma_{1}\mathbf{F}{D_{y}^{\Sigma}}^{\top}}_{\text{\emph{advection}}}.$ (6)
+We approximate the spatial derivative with a second-order finite-difference scheme using a three-point stencil, and obtain the following matrix differential equation for $\mathbf{F}$, for which its $(i,j)$ entry $F_{i,j}$ corresponds to the approximation of $f(x_{i},y_{j})$:
+$\frac{\partial\mathbf{F}}{\partial t}={D_{xx}^{\Phi}\mathbf{F}\Phi_{2}^{\top}+\Phi_{1}\mathbf{F}{D_{yy}^{\Phi}}^{\top}}-{D_{x}^{\Sigma}\mathbf{F}\Sigma_{2}^{\top}-\Sigma_{1}\mathbf{F}{D_{y}^{\Sigma}}^{\top}}.$ (9)
 Here, $D_{xx}^{\Phi},D_{yy}^{\Phi},D_{x}^{\Sigma},D_{y}^{\Sigma}$ represent compositions of difference operators with pointwise multiplication operators.
 See Appendices A and B for more details.
-For notational simplicity, let
+For notational simplicity, let:
 $T_{1}=D_{xx}^{\Phi},\quad T_{2}=D_{yy}^{\Phi},\quad T_{3}=-D_{x}^{\Sigma},\quad T_{4}=-D_{y}^{\Sigma}.$ (10)
@@
 ${\mathscr{L}}:\mathbf{F}\mapsto T_{1}\mathbf{F}\Phi_{2}^{\top}+\Phi_{1}\mathbf{F}T_{2}^{\top}+T_{3}\mathbf{F}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}T_{4}^{\top}.$ (12)
-We consider first a backward Euler implicit treatment of the temporal component of the matrix differential equation (8), which gives the following generalized Sylvester equation:
+We consider first a backward Euler implicit treatment of the temporal component of the matrix differential equation (11), which gives the following discrete GSE:
 $\mathbf{F}_{1}-\Delta t\mathscr{L}({\bf F}_{1})=\mathbf{F}_{0},$ (13)
-with the residual equation defined as:
-$\mathbf{R_{\mathscr{L}}}=\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}=\mathbf{0}.$ (11)
-An iterative solver naturally updates the solution $\mathbf{F}_{1}$ until the residual is sufficiently small under a suitable norm.
-If we vectorize the right-hand side using the Kronecker product identity, $\text{vec}(A\mathbf{F}B^{\top})=(B\otimes A)\text{vec}(\mathbf{F})$, the second term with a factor of $\Delta t$ in (10) can be written as a matrix-vector multiplication of the form:
-$\mathbf{F}\mapsto\left[\underbrace{\left(\Phi_{2}\otimes T_{1}\right)+\left(T_{2}\otimes\Phi_{1}\right)+\left(\Sigma_{2}\otimes T_{3}\right)+\left(T_{4}\otimes\Sigma_{1}\right)}_{\in\mathbb{R}^{N_{1}N_{2}\times N_{1}N_{2}}}\right]\text{vec}(\mathbf{F}),$ (12)
-The Kronecker form yields a linear system $Ax=b$, which can be solved either directly or iteratively.
-The Kronecker approach offers valuable theoretical insights into the relationship between the Kronecker structure and the constructed Krylov basis, a connection exploited in the context of Sylvester equations in [20, 23].
-We note that, based on the discussion in [26] on the advantages of the dimension-wise Krylov subspaces projection approach vs.
-the Kronecker formulation, we choose to adopt the former approach for better computational efficiency and iterative convergence.
+and associated residual equation:
+$\mathbf{R_{\mathscr{L}}}=\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}.$ (14)
+Iterative solvers update the solution $\mathbf{F}_{1}$ until the residual is sufficiently small under a suitable norm.
+In this work, we opt for Galerkin-projection-based iterative solvers, where the bases spanning $\mathbf{F}_{1}$ are updated by augmenting Krylov-like subspaces, as shown next.
 3 Krylov-Based Adaptive-Rank Solver for Generalized Sylvester Equation
-We propose an adaptive-rank algorithm for the generalized Sylvester equation (10).
-We assume a low-rank factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ at time $t^{(1)}=t^{(0)}+\Delta t$ also in the low-rank format.
-We represent this as:
+We propose an adaptive-rank algorithm for the GSE (13).
+We assume a low-rank factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ at time $t^{(1)}=t^{(0)}+\Delta t$ also in the low-rank format, i.e.:
 […]
 where $U_{0}\in\mathbb{R}^{N_{1}\times r_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, and $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$ and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$ are diagonal matrices with singular values at the corresponding times.
-Our method involves constructing the bases $U_{1}$ and $V_{1}$ and then projecting the original equation to determine the coefficients of $S_{1}$.
-A key aspect of the proposed adaptive-rank approach is that we do not form the full solutions explicitly.
-Instead, we seek bases in a dimension-by-dimension manner and then use them to perform Galerkin projection, resulting in a reduced system for the matrix of coefficients $S_{1}$.
-An SVD truncation procedure is then applied to update the orthonormal basis sets and realize an efficient low-rank representation of the updated solution.
-Key questions for solving the generalized Sylvester equation in the form of (10) that we address in this paper include:
-  • How do we choose suitable subspaces to build $U_{1}$ and $V_{1}$ from $U_{0}$ and $V_{0}$?
+Projection methods involve constructing bases $U_{1}$ and $V_{1}$ and then projecting the original equation onto them to determine the coefficients of $S_{1}$.
+A key aspect of the proposed adaptive-rank approach is that we never form the full solutions explicitly.
+Instead, we seek bases in a dimension-by-dimension manner and then use them to perform a Galerkin projection.
+Key questions for solving the GSE that we address in this study include:
+  • How do we choose suitable subspaces effectively to build $U_{1}$ and $V_{1}$ from $U_{0}$ and $V_{0}$?
   • How can we solve the reduced equation for $S_{1}$ efficiently?
-  • How can we check the residual of the generalized Sylvester equation without forming the matrices explicitly?
-Our main innovation involves leveraging an approximate operator denoted by $\tilde{\mathscr{L}}$ (which is of the Sylvester form) to the original operator ${\mathscr{L}}$ in (9).
-$\tilde{\mathscr{L}}$ approximates the diagonal matrices $\Phi_{i}$ and $\Sigma_{i}$ representing pointwise multiplication of diffusion and advection coefficients, respectively, by scaled identities.
-This approximation not only allows us to construct dimension-wise Krylov bases $U_{1}$ and $V_{1}$ effectively and efficiently, but also enables effective preconditioning in the spirit of [25].
-Then we project Equation (10) onto the Krylov bases, resulting in a reduced system for the coefficient matrix $S_{1}$.
-In this section, we introduce the main steps of our proposed algorithm as illustrated in Figure 1, highlighting the intuition behind our design and innovations along the way.
-Section 3.1 covers the construction of the approximated operator $\tilde{\mathscr{L}}$.
-Section 3.2 discusses the dimension-wise basis construction related to step 1 of the flowchart.
-In Section 3.3, we present the reduced equation for the matrix of coefficients $S_{1}$ and introduce our ACS preconditioner, corresponding to step 2 (i) in the flowchart.
-Lastly, Section 3.4 outlines the solution truncation mechanism and the residual computation, as described in steps 2(ii)-(iii) of the flowchart.
-Figure 1: Flow-chart of the proposed implicit adaptive-rank algorithm for a generalized Sylvester equation.
+  • How can we check the residual of the GSE without forming the matrices explicitly?
+Our main innovation leverages an approximate operator $\tilde{\mathscr{L}}$ (which is of the Sylvester form) to the original operator ${\mathscr{L}}$ in (12).
+This approximation not only allows us to construct dimension-wise Krylov bases $U_{1}$ and $V_{1}$ effectively and efficiently, but also enables effective preconditioning for solving the projected reduced system for $S_{1}$, in the spirit of [31].
+In the following, we cover the construction of the approximated operator $\tilde{\mathscr{L}}$ in Section 3.1; we discuss the dimension-wise basis construction in Section 3.2; we present the reduced equation for the matrix of coefficients $S_{1}$ and introduce our ACS preconditioner in Section 3.3, and outline the residual computation in Section 3.4.
+Finally, Section 3.5 extends the algorithm to the case of multi-rank coefficients.
 3.1 Averaged-Coefficient Approximate Operator $\tilde{\mathscr{L}}$
 Our approach incorporates ideas from [31] and [14] to construct an approximate operator $\tilde{\mathscr{L}}$ and efficiently solve the time-dependent inhomogeneous advection-diffusion equation using a projection-based adaptive-rank integration framework.
-The approximated operator $\tilde{\mathscr{L}}$ serves two purposes: (i) constructing the solution bases $U_{1}$ and $V_{1}$ as extended Krylov subspaces, and (ii) preconditioning the solution of the coefficient matrix $S_{1}$.
-We first construct a basis informed by both the original operator $\mathscr{L}$ and the approximate operator $\tilde{\mathscr{L}}$ (defined below).
-Secondly, we derive a corresponding projected (reduced) system that we solve iteratively using GMRES with a preconditioner derived from $\tilde{\mathscr{L}}$.
-We seek an approximation to operator (9) that retains its spectral content and is easier to invert.
-As proposed in [25], the diagonal operators can be approximated by scaling with an identity matrix, where the scaling factor is the average of the diagonal entries.
-For instance, let $\Phi_{i}\approx{\alpha}_{i}I$ and $\Sigma_{i}\approx\gamma_{i}I$, where $\alpha_{i}$ and $\gamma_{i}$ are the average of the diagonal entries of $\Phi_{i}$ and $\Sigma_{i}$, respectively.
+The approximated operator $\tilde{\mathscr{L}}$ serves two purposes: (i) constructing the solution bases $U_{1}$ and $V_{1}$ as xKrylov subspaces, and (ii) preconditioning the solution of the coefficient matrix $S_{1}$.
+We seek an approximation to ${\mathscr{L}}$ that captures its spectral content and is easier to invert.
+As proposed in [31], the advection diagonal operators can be approximated by scaling an identity matrix, where the scaling factor is the average of the diagonal entries.
+In this work, we approximate not only the advection coefficients by a suitable constant multiple of the identity, but extend this strategy to the scalar diffusion coefficients by averaging them across the computational domain.
+Let $\Phi_{i}\approx{\alpha}_{i}I$ and $\Sigma_{i}\approx\gamma_{i}I$, where $\alpha_{i}$ and $\gamma_{i}$ are the averages of the diagonal entries of $\Phi_{i}$ and $\Sigma_{i}$, respectively.
 This approximation yields the following mapping:
 ~${\tilde{\mathscr{L}}}:\mathbf{F}\mapsto\left(\alpha_{2}T_{1}+\gamma_{2}T_{3}\right)\mathbf{F}+\mathbf{F}\left(\alpha_{1}T_{2}+\gamma_{1}T_{4}\right)^{\top}.$ (16)
-A backward Euler temporal discretization using (14) yields the residual equation:
-$\mathbf{R_{\tilde{\mathscr{L}}}}=\mathbf{F}_{1}-\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}=\mathbf{0},$
-which leads to the Sylvester equation:
-$(A_{1}+A_{3})\mathbf{F}_{1}+\mathbf{F}_{1}(A_{2}+A_{4})^{\top}=\mathbf{F}_{0}.$ (15)
-Here:
+A backward Euler temporal discretization using (16) leads to the SE:
+$(A_{1}+A_{3})\mathbf{F}_{1}+\mathbf{F}_{1}(A_{2}+A_{4})^{\top}=\mathbf{F}_{0},$ (17)
+with:
 $A_{1}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{2}T_{1}\right)\quad A_{3}=\left(\frac{1}{4}I-\Delta t\,\gamma_{2}T_{3}\right),\quad P_{1}=A_{1}+A_{3},$ (18)
 $A_{2}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{1}T_{2}\right)\quad A_{4}=\left(\frac{1}{4}I-\Delta t\,\gamma_{1}T_{4}\right),\quad P_{2}=A_{2}+A_{4}.$ (19)
-The solution to Equation (15) is an approximation to that of (10), but it can be efficiently computed using low-rank approximations [10, 33, 31, 30].
-It can therefore be exploited as a viable preconditioner for a GMRES iterative solution of the generalized Sylvester equation (10).
+The solution to (17) is an approximation to that of (13), but it can be efficiently computed using low-rank approximations [14, 39, 37, 36].
+It can therefore be exploited as a viable preconditioner for a GMRES iterative solution of the GSE (13).
 This was done in [31] with $\mathcal{O}(N^{3})$ computational complexity.
-In this work, we extend this approach by performing a Galerkin projection of both the original operator $\mathscr{L}$ and its approximation $\tilde{\mathscr{L}}$ on the low-rank bases before solving the generalized Sylvester equation with GMRES.
-This reduces the problem to a lower-dimensional subspace, where preconditioning can be applied more effectively on a smaller problem with $\mathcal{O}(r^{3})$ complexity, resulting in an overall computational complexity of $\mathcal{O}(Nr^{2}+r^{3})$.
+In this work, we extend this approach by performing a Galerkin projection of both the original operator $\mathscr{L}$ and its approximation $\tilde{\mathscr{L}}$ on the low-rank bases before solving the GSE with GMRES.
+This reduces the problem to a lower-dimensional subspace, where preconditioning can be applied more effectively on a smaller $d$-dimensional problem with $\mathcal{O}(r^{d+1})$ complexity, resulting in an overall computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$.
 3.2 Basis Construction
-We begin by defining the generalized Krylov subspaces used for our basis construction at the $m$th iteration:
-$\kappa_{m}(M_{1},\cdots,M_{l},U)=[U,\underbrace{M_{1}U,\cdots,M_{l}U}_{\text{1st augmentation}},\ldots,\underbrace{M_{1}^{m+1}U,\cdots,M_{l}^{m+1}U}_{\text{mth augmentation}}].$ (18)
-where $M_{1},\cdots,M_{l}$ are operators of interest and their inverses arising in our generalized Sylvester equation, and $l$ is the number of operators used for each basis augmentation step.
-In our setting, the basis construction involves several operators from the original generalized Sylvester equation (10) and its approximation (15).
-We first consider the action of the diagonal operators, which have been averaged to obtain the approximated operators, $\Phi_{1}$, $\Sigma_{1}$, $\Phi_{2}$, and $\Sigma_{2}$.
-We have found that including these matrices is crucial for accelerating convergence.
-Using these operators, we construct a standard Krylov subspace.
-Additionally, we use the inverse operators from the approximated operators $A_{1:4}$ in (16)-(17) to form an inverted Krylov subspace.
-Finally, we construct dimension-wise extended Krylov subspaces for the approximated operator $\tilde{\mathscr{L}}$ using $P_{1}$ and $P_{2}$.
-The reasoning behind these choices of operators for subspace construction is as follows.
-Following [10], we construct an extended Krylov subspace in the $x$-direction for the approximated Sylvester equation (15) using $P_{1}$ and its inverse.
-To further enrich the basis, we include the inverses of the individual operators $A_{1}$ and $A_{3}$, as they may contain information not captured by the inverse of $P_{1}$.
-This carefully chosen set of operators results in a compact representation of the solution matrix, enriching the basis while keeping its dimension small.
-Combining these, we define the extended Krylov subspaces in the $x$ direction as follows:
-$\kappa_{m}\left(P_{1},P_{1}^{-1},A_{1}^{-1},A_{3}^{-1},\Phi_{1},\Sigma_{1},U_{0}\right),$
-We then apply an SVD-truncated QR, as outlined in Algorithm 1, to orthogonalize the subspace and obtain a basis $U_{1}^{(m)}\in\mathbb{{\bf R}}^{N_{1}\times r^{(m)}_{x}}$ and remove vectors that are closely linearly dependent.
-By leveraging the rank-revealing property of SVD, only those basis associated with significant singular values are retained.
-The same procedures are applied to the $y$-direction to obtain $V^{(m)}_{1}\in\mathbb{{\bf R}}^{N_{2}\times r^{(m)}_{y}}$.
-For notational simplicity, we will omit the superscript $(m)$ from now on, understanding that $U_{1}\in\mathbb{{\bf R}}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{{\bf R}}^{N_{2}\times r_{y}}$ refer to the bases of the Krylov subspaces at the $m$th iteration.
-This truncated orthonormalization strategy is particularly advantageous when the underlying processes (diffusion or advection) are weak.
-In such cases, the approximated coefficient matrices contribute minimally, reducing the impact of the operator application and enabling further basis truncation.
-These constructed orthonormal bases will be used to represent the solution in the approximation subspace.
-The coefficients composing the linear combination of these basis vectors will be obtained by projecting the corresponding generalized Sylvester equation and evolving the matrix of coefficients $S_{1}$.
+We begin by defining the xKrylov subspaces used for our basis construction at the $m$-th outer iteration as:
+$\kappa_{m}(M_{1},\cdots,M_{l},U)=\text{{\color[rgb]{0,0,0}Range}}\Big{\{}U,\underbrace{M_{1}U,\cdots,M_{l}U}_{\text{1st augmentation}},\ldots,\underbrace{M_{1}^{m+1}U,\cdots,M_{l}^{m+1}U}_{\text{mth augmentation}}\Big{\}}.$ (20)
+where $M_{1},\cdots,M_{l}$ are operators of interest and their inverses arising in our GSE, and $l$ is the number of operators used for each basis augmentation step.
+The xKrylov basis construction involves several operators from the original GSE (13) and its approximation (17).
+Following [14], we construct an xKrylov subspace in the $x$-direction for the approximated SE (17) using $P_{1}$ and its inverse.
+We also include the inverses of the individual operators $A_{1}$ and $A_{3}$, as they may contain information not captured by the inverse of $P_{1}$.
+Additionally, we consider the action of the diagonal operators $\Phi_{1}$, $\Sigma_{1}$, $\Phi_{2}$, and $\Sigma_{2}$, which are averaged to obtain $\tilde{\mathscr{L}}$.
+We have found that including all these matrices is effective in capturing the range of the solution.
+Using these operators, we construct a generalized Krylov subspace in the $x$-direction as follows:
+$\kappa_{m}\left(M_{1},\cdots,M_{6},U_{0}\right),\quad\text{where}$ $M_{1}=P_{1},$ $M_{2}=P_{1}^{-1},$ $M_{3}=A_{1}^{-1},$ $M_{4}=A_{3}^{-1},$ $M_{5}=\Phi_{1},$ $M_{6}=\Sigma_{1}.$
+Depending on the structure of the GSE (2), $l$ may be large, and the generated subspace vectors at each augmentation may exhibit linear dependencies.
+To limit subspace inflation, we employ an SVD-based truncation strategy (Algorithm 1), such that only basis vectors corresponding to significant singular values are retained.
+The construction of the dimension-wise xKrylov subspaces (20), following [39], is as follows.
+At each iteration $m$, Algorithm 2 augments the previously constructed basis $\mathcal{V}_{(m-1)}$ by applying powers of operators $M_{1}$ through $M_{l}$ to their corresponding basis blocks $V^{(1)},\dots,V^{(l)}$.
+Each block matrix $V^{(i)}$ may have a rank of at most $r_{0}$.
+We apply Algorithm 2 to orthogonalize the subspace and obtain a basis $\mathcal{U}_{(m)}\in\mathbb{R}^{N_{1}\times r^{(m)}_{x}}$ and $\mathcal{V}_{(m)}\in\mathbb{R}^{N_{2}\times r^{(m)}_{y}}$ in the $x$ and $y$ directions, respectively.
+For notational simplicity, we omit the calligraphic notation and subscript $(m)$ from now on, understanding that $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ refer to the bases of the Krylov subspaces at the $m$th iteration.
+The truncation step in line 2 of Algorithm 2 removes any linear dependencies within each block during the power iterations; line 2 applies a global truncation across all newly generated blocks, eliminating redundancies among them.
+As a result, the basis of the subspace is kept optimal in size.
 Input: Krylov Subspace $\kappa_{m}$; Tolerance $\epsilon_{\kappa}$.
 Output: Truncated bases ${Q}$.
+1 Compute the reduced QR decomposition: $\{Q,R\}=\texttt{QR}\left(\kappa_{m}\right)$;
+2 Perform reduced SVD decomposition: $Z_{1}\tilde{R}Z_{2}=\texttt{SVD}(R)$, where $\tilde{R}=\text{diag}(\sigma_{j})$ Identify the last index $r$ such that $\sigma_{r}>\epsilon_{\kappa}$;
+3 Update the basis vectors: $Q\leftarrow QZ_{1}(:,1:r)$;
+Algorithm 1 SVD-truncated QR, $\mathscr{T}_{\epsilon_{\kappa}}$
+Input: Operators $M_{1},\dots,M_{l}$; augmentation basis blocks $\{V^{(1)},\dots,V^{(l)}\}$; previous Krylov basis ${\mathcal{V}}_{(m-1)}$; tolerance $\epsilon_{\kappa}$.
+Output: Updated Krylov basis $\mathcal{V}_{(m)}$; basis blocks $\{V^{(1)},\dots,V^{(l)}\}$ for the next iteration.
 1
-2Compute the reduced QR decomposition: $\{Q,R\}=\texttt{QR}\left(\kappa_{m}\right)$;
-3
-4Perform reduced SVD decomposition: $T_{1}\tilde{R}T_{2}=\texttt{SVD}(R)$, where $\tilde{R}=\text{diag}(\sigma_{j})$;
+2for $i=1,\dots,l$ do
+3        $V^{(i)}\leftarrow M_{i}V^{(i)}$;
+4        $\widetilde{V}^{(i)}\leftarrow\text{orthogonalize }V^{(i)}\text{ w.r.t. }V_{(m-1)}$ (e.g., modified Gram-Schmidt);
+        $V^{(i)}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}(\widetilde{V}^{(i)})$ ;
+         // truncate to $\epsilon_{\kappa}$ for next iteration using Algorithm 1
 5
-6Identify the last index $r$ such that $\sigma_{r}>\epsilon_{\kappa}$;
-7
-8Update the basis vectors: $Q\leftarrow QT_{1}(:,1:r)$;
-9
-Algorithm 1 SVD-truncated QR, $\mathscr{T}_{\epsilon_{\kappa}}$
+[…]
+$V^{\prime}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}(V^{\prime})$ ;
+  // final orthogonalization and truncation using Algorithm 1
+7 $\mathcal{V}_{(m)}\leftarrow[\mathcal{V}_{(m-1)},V^{\prime}]$;
+Algorithm 2 Construction of xKrylov Subspace $\kappa_{m}^{\texttt{trunc}}$
+We comment next on the performance of the xKrylov iteration.
+In principle, the number of xKrylov iterations may depend on the mesh size, $N$.
+For instance, [24, 4] theoretically show for the 2D and 3D Poisson equations that the dimension of the extended Krylov subspace (denoted by $r^{(m)}$ in the basis construction above) scales as $\mathcal{O}(\kappa^{1/4})\sim\mathcal{O}(N^{1/2})$, with $\kappa\sim N^{2}$ the condition number.
+While this provides an upper bound of the convergence rate (without considering solution smoothness), our numerical results suggest that it may be pessimistic in practice.
+In particular, for moderately tight tolerances, we find that the required Krylov subspace dimensions remain independent of grid resolution for a wide range of grid sizes $N$.
 3.3 Solving for the Matrix of Coefficients, $S_{1}$: ACS-Preconditioned GMRES
-With the orthonormal bases $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ at hand, we perform a Galerkin projection to derive a reduced generalized Sylvester equation for the matrix $S_{1}\in\mathbb{R}^{r_{x}\times r_{y}}$.
-The reduced Sylvester equation for $S_{1}$ arises from projecting the residual via the Galerkin condition $U_{1}^{\top}\mathbf{R}_{\mathscr{L}}V_{1}=0$, which yields:
+With the orthonormal bases $U_{1}\in\mathbb{R}^{N_{1}\times r_{x}}$ and $V_{1}\in\mathbb{R}^{N_{2}\times r_{y}}$ at hand, we perform a Galerkin projection to derive a reduced GSE for the matrix $S_{1}\in\mathbb{R}^{r_{x}\times r_{y}}$.
+The reduced GSE for ${\bf S}_{1}$ arises from projecting the residual via the Galerkin condition $U_{1}^{\top}\mathbf{R}_{\mathscr{L}}V_{1}=0$, yielding:
 ${\bf S}_{1}-\Delta t\left(\tilde{T}_{1}{\bf S}_{1}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}{\bf S}_{1}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}_{1}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}{\bf S}_{1}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}_{1},\qquad\text{with}\qquad\tilde{{\bf B}}_{1}=(U_{1}^{\top}U_{0})S_{0}(V_{1}^{\top}V_{0})^{\top}.$ (21)
@@
 $\tilde{T}_{2}$ $=V_{1}^{\top}T_{2}V_{1},\quad\tilde{T}_{4}=V_{1}^{\top}T_{4}V_{1},\quad\tilde{\Phi}_{2}=V_{1}^{\top}\Phi_{2}V_{1},\quad\tilde{\Sigma}_{2}=V_{1}^{\top}\Sigma_{2}V_{1}.$
-Solving Equation (19) using the Kronecker formulation $\mathcal{A}\text{vec}({{\bf S}_{1}})=\text{vec}({\tilde{{\bf B}}_{1}})$ with a direct solver requires a complexity of $\mathcal{O}(r^{6})$, where $r=\text{max}(r_{x},r_{y})$, which may be too expensive even for moderate $r$-values.
-Instead, we solve this equation using preconditioned GMRES, where the operator-times-vector operation is performed using a matrix-based strategy, with complexity $\mathcal{O}(r^{3})$.
-This is possible due to the Kronecker structure of our operator, yielding the equivalence between (19) and (20).
-Given a vector $s_{k}=\text{vec}(\mathbf{S}_{k})$, the product of the matrix $\mathcal{A}$ times a vector $s_{k}$ is organized as follows:
+Solving the Kronecker formulation $\mathcal{A}\text{vec}({{\bf S}_{1}})=\text{vec}({\tilde{{\bf B}}_{1}})$ in (22) with a direct solver scales as $\mathcal{O}(r^{6})$ in 2D and $\mathcal{O}(r^{9})$ in 3D, where $r=\text{max}(r_{x},r_{y})$, which may be too expensive even for moderate $r$-values.
+Instead, we solve this equation iteratively using preconditioned GMRES.
+Instead of adapting GMRES to the matrix equation, we wrap GMRES around the Kronecker form, but exploiting the matrix-equation form (21) for both the $\mathcal{A}$-times-vector operation and the preconditioning step as indicated below, with overall complexity $\mathcal{O}(r^{3})$ in 2D (and $\mathcal{O}(r^{4})$ in 3D).
+For the GMRES operator-vector product, given a vector $s_{k}=\text{vec}(\mathbf{S}_{k})$, the product of the matrix $\mathcal{A}$ times a vector $s_{k}$ is performed as follows:
   1. 1.
@@
     Reshape $\hat{z}_{k}=\text{vec}(\hat{\mathbf{Z}}_{k})$.
-The computational complexity of the GMRES solve also depends crucially on the total number of iterations needed to invert the linear system.
-To bound the number of GMRES iterations with respect to the basis dimension (rank) $r$, we require effective preconditioning, such that the eigenvalues of the preconditioned matrix are clustered and the number of iterations becomes independent of the system size $r$ [4].
-Here, we propose a (left) preconditioner $\mathcal{P}$ such that the system:
+For efficiency, GMRES requires preconditioning so that the eigenvalues of the preconditioned matrix are sufficiently clustered to deliver $r$-independent convergence rates  [7].
+Here, we employ the ACS preconditioner, with $\mathcal{P}=I_{r_{y}}\otimes\tilde{P}_{1}+\tilde{P}_{2}\otimes I_{r_{x}}$ arising from the Galerkin projection of (17).
+The ACS preconditioner is expected to deliver optimal rank-independent, superlinear convergence (both observed in our numerical experiments), owing to its connection to the theory of compact-equivalent operators for diffusion-like equations, [^1] The averaged-coefficient advection-diffusion equation is compact-equivalent to the variable-coefficient one.
+which are known to be effective preconditioners for Krylov methods [1, 2].
+We consider left preconditioning as:
 $\mathcal{P}^{-1}\mathcal{A}\text{vec}({\bf S}_{1})=\mathcal{P}^{-1}\text{vec}(\tilde{{\bf B}}_{1}).$ (23)
-is well conditioned. The preconditioning step in GMRES requires applying $\mathcal{P}^{-1}$ to an arbitrary vector provided by the algorithm.
-For fast GMRES convergence, $\mathcal{P}$ should approximate the left-hand side of Equation (20) well while being easier to invert than the original system.
-Here, we consider $\mathcal{P}=I_{r_{y}}\otimes\tilde{P}_{1}+\tilde{P}_{2}\otimes I_{r_{x}}$ arising from the Galerkin projection of Equation (15).
-In particular, let $U_{1}$ and $V_{1}$ be the orthonormal basis for the generalized Krylov subspaces, and let $\tilde{P}_{1}=U_{1}^{\top}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{\top}P_{2}V_{1}$, then
-$\mathcal{P}:\mathbf{S}\mapsto{\tilde{P}_{1}}\mathbf{S}+\mathbf{S}{\tilde{P}_{2}^{\top}}.$ (22)
-To compute $\mathcal{P}^{-1}\hat{z}_{k}=z_{k}$, we solve the corresponding Sylvester equation:
+To compute $\mathcal{P}^{-1}\hat{z}_{k}=z_{k}$ for an arbitrary vector $\hat{z}_{k}$ during GMRES, we solve the corresponding SE:
 $\tilde{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\tilde{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k},$
-where $z_{k}=\text{vec}(\mathbf{Z}_{k})\in\mathbb{R}^{r_{x}r_{y}}$. Specific steps are as follows:
+where $z_{k}=\text{vec}(\mathbf{Z}_{k})\in\mathbb{R}^{r_{x}r_{y}}$. Specific steps are:
   1. 1.
@@
   2. 2.
-    Solve $\mathcal{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\mathcal{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k}$ with complexity $\mathcal{O}(r^{3})$, for example with the Bartels Stewart algorithm [11].
+    Solve $\mathcal{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\mathcal{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k}$ with complexity $\mathcal{O}(r^{d+1})$, for example with the Bartels-Stewart algorithm [15].
   3. 3.
     Reshape $z_{k}=\text{vec}(\mathbf{Z}_{k})$.
-The proposed preconditioner effectively bounds the number of GMRES iterations required to solve Equation (19) independently of $r$, as will be shown later in this study, rendering the computational complexity of the $S$ matrix solve as $O(r^{3})$, much better than the expected $O(r^{6})$ scaling of a naïve implementation of GMRES.
-3.4 Truncation, Efficient Residual Computation, and Algorithm Flowchart
-At the end of each iteration, we need to compute the residual to accept or reject the low-rank approximation from Krylov subspaces.
-This process should take advantage of the low-rank representation of the solution while avoiding forming the residual on the full grid.
+As a result, the $S$-matrix solve scales as $O(r^{d+1})$ overall, a great reduction in complexity compared to the $O(r^{3d})$ scaling of a naïve implementation of GMRES with the Kronecker form.
+3.4 Truncation, Efficient Residual Computation
 Once the solution ${\bf F}_{1}=U_{1}S_{1}V_{1}^{\top}$ is obtained, we apply truncation to control rank growth and reduce the residual computational load.
-Since $S_{1}$ from the GMRES solve is not necessarily diagonal, we follow the approach in [14] to convert it into a singular value form (also scaling as $\mathcal{O}(r^{3})$) and discard singular values below a threshold $\epsilon$.
-The truncation results in a solution ${\bf F}_{1}=U_{1}S_{1}V_{1}^{T}$, with $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$, and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$, with $r_{1}\leq\text{min}(r_{x},r_{y})$.
-This procedure is summarized in Algorithm 2.
-Following a similar procedure to [10], we devise a cost-effective way to compute the residual using the identity:
+Since $S_{1}$ resulting from the GMRES solve is not necessarily diagonal, we diagonalize it with SVD [19] [also scaling as $\mathcal{O}(r^{d+1})$] and discard singular values below a threshold $\epsilon$.
+The SVD truncation gives ${\bf F}_{1}=U_{1,\epsilon}S_{1,\epsilon}V_{1,\epsilon}^{\top}$, with $U_{1,\epsilon}\in\mathbb{R}^{N_{1}\times r_{1}}$, $V_{1,\epsilon}\in\mathbb{R}^{N_{2}\times r_{1}}$, and $S_{1,\epsilon}\in\mathbb{R}^{r_{1}\times r_{1}}$, with $r_{1}\leq\min(r_{x},r_{y})$ (Algorithm 3).
+Following [14], we implement a cost-effective way to compute the residual using the identity:
 $\|\mathbf{R_{{\mathscr{L}}}}\|$ $=\|\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}\|$
-$=\underbrace{\left\|R_{U}\begin{bmatrix}{S_{1}}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&-\Delta t\,{S_{1}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,{S_{1}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{S_{1}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{S_{1}}\\ \end{bmatrix}R_{V}^{\top}\right\|}_{\in\mathbb{R}^{5r_{1}\times 5r_{1}}}$
-$=\underbrace{\left\|R_{U}\begin{bmatrix}{S_{1}}-\tilde{B}_{1}&\mathbf{0}\\ \mathbf{0}&-\Delta t\,(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1})\\ \end{bmatrix}R_{V}^{\top}\right\|}\leq\epsilon_{\texttt{tol}},$
-where $\mathcal{R}_{\text{ranks}}={\ell_{x}+k_{x}+\ell_{y}+k_{y}}$ is the total rank number of the advection and diffusion coefficients, and where $R_{U}$ and $R_{V}$ are the upper triangular matrices arising from the following QR decomposition:
-{_$\{\_\ ,R_{U}\}$ $=\texttt{QR}\left(\begin{bmatrix}{U}_{1}&{T_{1}}{U}_{1}&\Phi_{1}{U}_{1}&{T_{3}}{U}_{1}&\Sigma_{1}{U}_{1}\end{bmatrix}\right),$ (23)
-{_$\{\_\ ,R_{V}\}$ $=\texttt{QR}\left(\begin{bmatrix}{V}_{1}&\Phi_{2}{V}_{1}&{T_{2}}{V}_{1}&\Sigma_{2}{V}_{1}&{T_{4}}{V}_{1}\end{bmatrix}\right),$
-where the orthogonal columns from the decomposition are ignored.
-The above residual result follows because, if ${\bf F}_{0}=U_{0}S_{0}V_{0}^{\top}$ (where $U_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ and $V_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ have orthonormal columns) then, since $U_{1}$ and $V_{1}$ contain $U_{0}$ and $V_{0}$, respectively, we have $U_{1}(U_{1})^{\top}{\bf F}_{0}V_{1}(V_{1})^{\top}={\bf F}_{0}$.
-Input: Bases $U$, $V$; matrix of coefficients $S$; Tolerance $\epsilon$.
+$=\underbrace{\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}\\ \end{bmatrix}R_{V}^{\top}\right\|}_{\in\mathbb{R}^{(r_{x}+5r_{1})\times(r_{y}+5r_{1})}}$
+$={\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{S_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,(I_{\mathcal{R}_{\text{ranks}}}\otimes{S_{1,\epsilon}})\end{bmatrix}R_{V}^{\top}\right\|}\leq\epsilon_{\texttt{tol}},$
+where $\mathcal{R}_{\text{ranks}}={\ell_{x}+k_{x}+\ell_{y}+k_{y}}$ is the total rank number of the advection and diffusion coefficients, and $R_{U}$ and $R_{V}$ are the upper triangular matrices arising from the following Q-less QR decomposition:
+{_$\{\_\ ,R_{U}\}$ $=\texttt{QR}\left(\begin{bmatrix}{U}_{1}&{U}_{1,\epsilon}&{T_{1}}{U}_{1,\epsilon}&\Phi_{1}{U}_{1,\epsilon}&{T_{3}}{U}_{1,\epsilon}&\Sigma_{1}{U}_{1,\epsilon}\end{bmatrix}\right),$ (24)
+{_$\{\_\ ,R_{V}\}$ $=\texttt{QR}\left(\begin{bmatrix}{V}_{1}&{V}_{1,\epsilon}&\Phi_{2}{V}_{1,\epsilon}&{T_{2}}{V}_{1,\epsilon}&\Sigma_{2}{V}_{1,\epsilon}&{T_{4}}{V}_{1,\epsilon}\end{bmatrix}\right).$
+The above residual result follows from the property that $U_{1}(U_{1})^{\top}{\bf F}_{0}V_{1}(V_{1})^{\top}={\bf F}_{0}$, with ${\bf F}_{0}=U_{0}S_{0}V_{0}^{\top}$ (where $U_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ and $V_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ have orthonormal columns), since $U_{1}$ and $V_{1}$ contain $U_{0}$ and $V_{0}$, respectively.
+Input: Bases $U$, $V$; matrix of coefficients ${\bf S}$; Tolerance $\epsilon$.
 Output: Truncated bases $\tilde{U}$, $\tilde{V}$; truncated singular values $\tilde{S}$.
@@
 3.5 Generalization to Multi-rank Coefficients
-We generalize next the adaptive-rank approach from rank-one advection and diffusion coefficients to the multi-rank case, using backward Euler.
-The corresponding pseudo-algorithm is presented in Algorithm 3.
-High-order temporal discretizations are considered later in this study.
-Discretization of the multi-rank advection-diffusion equation yields a multi-term generalized Sylvester equation of the form:
+We generalize next the adaptive-rank approach to the multi-rank coefficient case in (5) and (6), using backward Euler for now.
+Discretization of the advection-diffusion equation in this case yields a multi-term GSE of the form:
 $\mathbf{F}_{1}-\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}_{1}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}_{1}T_{2,j}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}_{1}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}_{1}T_{4,l}^{\top}}_{\mathscr{L}(\mathbf{F}_{1})}\right)={\bf F}_{0}.$ (25)
 […]
-$\alpha_{x,i}$ $=\text{average}(\Phi_{i}^{2,x}),\quad\alpha_{y,j}$ $=\text{average}(\Phi_{j}^{1,y}),\quad\gamma_{x,k}$ $=\text{average}(\Sigma_{k}^{2,x}),\quad\gamma_{y,l}$ $=\text{average}(\Sigma_{l}^{1,y}),$ (25)
-to arrive at the approximated Sylvester equation:
 ${P_{1}}\mathbf{F}_{1}+\mathbf{F}_{1}{P_{2}^{\top}}=\mathbf{F}_{0},$ (26)
 where
-$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{3,k}}=\sum_{i=1}^{\ell_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\alpha_{x,i}T_{1,i}\right)}+\sum_{k=1}^{k_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\gamma_{x,k}T_{3,k}\right)},$
-$P_{2}$ $=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{4,l}}=\sum_{j=1}^{\ell_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\alpha_{y,j}T_{2,j}\right)}+\sum_{l=1}^{k_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{y,l}T_{4,l}\right)}.$
-The bases $U_{1}$ and $V_{1}$ are constructed from orthonormalization of dimension-wise generalized Krylov subspaces from equations (24) and (26) as follows:
-$U_{1}$ $=\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right)\right),$ (27)
-$V_{1}$ $=\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right)\right),$ (28)
-In the multi-rank setting, after a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$, we obtain a generalized Sylvester equation similar to (29) for the matrix of coefficients $S_{1}$:
-${\bf S}_{1}-\Delta_{t}\left(\sum_{i=1}^{\ell_{x}}\tilde{T}_{1,i}{\bf S}_{1}\tilde{\Phi}_{i}^{2,x\top}+\sum_{j=1}^{\ell_{y}}\tilde{\Phi}_{j}^{1,y}{\bf S}_{1}\tilde{T}_{2,j}^{\top}+\sum_{k=1}^{k_{x}}\tilde{T}_{3,k}{\bf S}_{1}\tilde{\Sigma}_{k}^{2,x\top}+\sum_{l=1}^{k_{y}}\tilde{\Sigma}_{l}^{1,y}{\bf S}_{1}\tilde{T}_{4,l}^{\top}\right)=\tilde{{\bf B}}_{1}$ (29)
+$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{3,k}}=\sum_{i=1}^{\ell_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\alpha_{x,i}T_{1,i}\right)}+\sum_{k=1}^{k_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\gamma_{x,k}T_{3,k}\right)},$ (27)
+$P_{2}$ $=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{4,l}}=\sum_{j=1}^{\ell_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\alpha_{y,j}T_{2,j}\right)}+\sum_{l=1}^{k_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{y,l}T_{4,l}\right)}.$ (28)
+Also as before, the bases $U_{1}$ and $V_{1}$ are constructed from orthonormalization of dimension-wise xKrylov from equations (25) and (26) as:
+$U_{1}$ $=\kappa^{\texttt{trunc}}_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right),$ (29)
+$V_{1}$ $=\kappa^{\texttt{trunc}}_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right).$ (30)
+In the multi-rank setting, after a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$, we obtain a GSE similar to (21) for the matrix of coefficients ${\bf S}_{1}$:
+${\bf S}_{1}-\Delta_{t}\left(\sum_{i=1}^{\ell_{x}}\tilde{T}_{1,i}{\bf S}_{1}\tilde{\Phi}_{i}^{2,x\top}+\sum_{j=1}^{\ell_{y}}\tilde{\Phi}_{j}^{1,y}{\bf S}_{1}\tilde{T}_{2,j}^{\top}+\sum_{k=1}^{k_{x}}\tilde{T}_{3,k}{\bf S}_{1}\tilde{\Sigma}_{k}^{2,x\top}+\sum_{l=1}^{k_{y}}\tilde{\Sigma}_{l}^{1,y}{\bf S}_{1}\tilde{T}_{4,l}^{\top}\right)=\tilde{{\bf B}}_{1},$ (31)
 where the projected operators are analogous to those defined in the rank-one case.
-[…]
-$\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}=\tilde{B}_{1},$ (30)
+We solve (31) for the ${\bf S}_{1}$ matrix using preconditioned GMRES.
+The preconditioner is constructed by another Galerkin projection to find the projected SE:
+$\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}=\tilde{{\bf B}}_{1},$ (32)
 with:
@@
 $\tilde{P}_{2}$ $=\sum_{j=1}^{\ell_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{y}}-\Delta_{t}\alpha_{y,j}\tilde{T}_{2,j}\right)}_{\tilde{A}_{2,j}}+\sum_{l=1}^{k_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{r_{y}}-\Delta_{t}\gamma_{y,l}\tilde{T}_{4,l}\right)}_{\tilde{A}_{4,l}}.$
-This yields the preconditioner operator:
-$\mathscr{P}:\mathbf{S}\mapsto\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}$ (31)
-The whole process is summarized in Algorithm 3.
+This yields the multi-rank-coefficients ACS preconditioner operator:
+$\mathscr{P}:\mathbf{S}\mapsto\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}.$ (33)
+We summarize the pseudo-algorithm for multi-rank coefficients in Algorithm 4.
+The tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, $\epsilon_{\texttt{tol}}$ in the algorithm are set as follows.
+First, we choose the residual tolerance for xKrylov outer iterations, $\epsilon_{\texttt{tol}}$.
+To prevent rapid growth of the xKrylov subspaces, we choose the Krylov basis truncation threshold as $\epsilon_{\kappa}\leq\epsilon_{\texttt{tol}}$, typically two to three orders of magnitude smaller.
+The GMRES-ACS tolerance for the inner reduced system solve, $\epsilon_{\texttt{GMRES}}$, is initially set to be less than $\epsilon_{\texttt{tol}}$ for the first outer-basis iteration.
+It is subsequently adjusted to be two to three orders of magnitude smaller than the relative residual error from the previous xKrylov outer iteration.
+This adjustment ensures that errors from the inner solve do not pollute the accuracy of the outer solve.
+The SVD truncation threshold $\epsilon$ is chosen to be one to three orders of magnitude smaller than the tolerance $\epsilon_{\text{tol}}$.
+Since solution truncation occurs before residual evaluations, the truncation error in the solution should remain below the residual tolerance to prevent residual pollution by over-truncation.
 // This algorithm solves Eq. (25) for $\mathbf{F}_{1}=U_{1}S_{1}V_{1}^{\top}$, where $\mathbf{F}_{0}=U_{0}S_{0}V_{0}^{\top}$
@@
 Operators $P_{1}$, $P_{2}$, $\{T_{1,i}\}_{i=1}^{\ell_{x}}$, $\{T_{2,j}\}_{j=1}^{\ell_{y}}$, $\{T_{3,k}\}_{k=1}^{k_{x}}$, $\{T_{4,l}\}_{l=1}^{k_{y}}$;
-Tolerances $\epsilon_{\kappa}$, $\epsilon$, $\epsilon_{\texttt{tol}}$;
-Maximum iterations max_iter;
+Tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, $\epsilon_{\texttt{tol}}$;
 Output: Updated bases $U_{1}$, $V_{1}$;
 Truncated singular values $S_{1}$
-1Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eq. (26);
-2
-3for $m=1$ to max_iter do
+1Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eqs. (27), (28);
+2 Compute $l=2+\ell_{x}+\ell_{y}+k_{x}+k_{y}$;
+3 Set $U^{(i)}=U_{0}$ and $V^{(i)}=V_{0}$ for $i=1,\dots,l$;
+4 Set $U_{1}=U_{0}\quad V_{1}=V_{0}$;
+5 while not converged do
 6        // Step K1.
-5        Truncate to tolerance $\epsilon_{\kappa}$;
-6        $U_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right)\right)$;
-7
-8       $V_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right)\right)$;
-9
+7        Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
+8        $U_{1},\{U^{(1)}\cdots U^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}}\};\{U^{(1)}\cdots U^{(l)}\};U_{1};\epsilon_{\kappa}\right)$;
+9        $V_{1},\{V^{(1)}\cdots V^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}}\};\{V^{(1)}\cdots V^{(l)}\};V_{1};\epsilon_{\kappa}\right)$;
+10
 11       // Step K2.
-11        Solve the reduced equation (29) for $S_{1}$ using GMRES-ACS (31) to tolerance $\epsilon_{\texttt{GMRES}}$;
+12        Solve the reduced GSE (31) for ${\bf S}_{1}$ using GMRES-ACS (33) to tolerance $\epsilon_{\texttt{GMRES}}$;
 13        // Step K3.
-13        Truncate $\{U_{1},S_{1},V_{1}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}\})$ to tolerance $\epsilon$;
+14        Truncate $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}\})$ to tolerance $\epsilon$ (Alg. 3);
+15
 16       // Step K4.
-15        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},T_{1,1:\ell_{x}}U_{1},\Phi_{1:\ell_{y}}^{1,y}U_{1},T_{3,1:k_{x}}U_{1},\Sigma_{1:k_{y}}^{1,y}U_{1}]\right)$ and
-16        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},\Phi_{1:\ell_{x}}^{2,x}V_{1},T_{2,1:\ell_{y}}V_{1},\Sigma_{1:k_{x}}^{2,x}V_{1},T_{4,1:k_{y}}V_{1}]\right)$ ;
-17
-18       Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(S_{1}-\tilde{B}_{1},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1}\right)\right)R_{V}^{\top}\right\|$;
+17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
+18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
+19
+20       Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(-\tilde{B}_{1},S_{1,\epsilon},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)R_{V}^{\top}\right\|$;
+21
 22       if $\|\mathbf{R}\|/\|\mathbf{F}_{0}\|\geq\epsilon_{\texttt{tol}}$ then
-20               Reject solution and return to augment bases further;
-21
+23               Reject solution and cycle to augment bases further;
+24
 25       else
-23
+26               Set $U_{1}\leftarrow U_{1,\epsilon}$; $U_{1}\leftarrow V_{1,\epsilon}$; $S_{1}=S_{1,\epsilon}$;
 27               Break;
 28               Exit loop;
-26
-27
-Algorithm 3 Backward Euler Adaptive-Rank Integrator for the Generalized Sylvester Equation.
-00footnotetext: The MATLAB implementation of the algorithm is publicly available on GitHub, Adaptive-rank-ACS, https://github.com/helkahza1/Adaptive-rank-ACS
+29
+30
+Algorithm 4 Backward Euler Adaptive-Rank Integrator for the GSE.
 3.6 Computational Complexity
-We now analyze the computational complexity of Algorithm 3 for backward Euler.
+We analyze next the computational complexity of Algorithm 4 for backward Euler.
 For conciseness, we consider the same spatial resolution per dimension, i.e., $N=N_{1}=N_{2}$.
-Starting from an initial condition $U_{0}S_{0}V_{0}^{\top}$, where $U_{0}\in\mathbb{R}^{N\times r_{0}}$, $V_{0}\in\mathbb{R}^{N\times r_{0}}$, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$, the $m$-th extended Krylov iteration generates $U_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, $V_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, and $S_{1}\in\mathbb{R}^{r_{\kappa}\times r_{\kappa}}$, where $r_{\kappa}=((2+\mathcal{R}_{\text{ranks}})m+1)r_{0}$.
-Using SVD-truncated QR, this dimension is then reduced to smaller ranks, $r_{x}$ and $r_{y}$, for the $x$ and $y$ directions, respectively.
-The matrix $S_{1}$ from the $m$-th Krylov iteration solve, initially of size $r_{x}\times r_{y}$, is truncated to rank $r_{1}\leq\min(r_{x},r_{y})$ using Algorithm 2.
-To simplify the complexity analysis, similar to Section 3.3, let $r=\max(r_{x},r_{y})$.
+Starting from an initial condition $U_{0}S_{0}V_{0}^{\top}$, where $U_{0}\in\mathbb{R}^{N\times r_{0}}$, $V_{0}\in\mathbb{R}^{N\times r_{0}}$, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$, the $m$-th xKrylov iteration generates $U_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, $V_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, and $S_{1}\in\mathbb{R}^{r_{\kappa}\times r_{\kappa}}$, where $r_{\kappa}=(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0}+r^{(m-1)}$.
+Here, $r^{(m-1)}$ denotes the basis size from the previous xKrylov iteration (initialized with $r^{(0)}=r_{0}$), and $r^{\ast}_{0}\leq r_{0}$ is the possibly truncated initial rank as produced by line (2) of Algorithm 2.
+Using SVD-truncated QR, this dimension is then reduced to smaller ranks, $r_{x}$ and $r_{y}$.
+For simplicity, let $\tilde{r}=\max(r_{x},r_{y})$.
+The matrix $S_{1}$ from the $m$-th inner iteration solve, initially of size $r_{x}\times r_{y}$, is truncated to rank $r_{1}\leq\min(r_{x},r_{y})$ using Algorithm 3.
+To simplify the complexity analysis, let $r=\max(r_{0},r^{(m-1)},\tilde{r},(\mathcal{R}_{\text{ranks}}+1)r_{1})$, which we consider independent of the mesh-refinement $N$ for practical purposes.
 The computational complexity of each step is then estimated as follows:
 Step K1
-Constructing the generalized Krylov basis as outlined in lines (3) and (3) of Algorithm 3 requires performing the operations $A_{1,1:r_{x}}^{-1}U_{0}$ and $A_{3,1:k_{x}}^{-1}U_{0}$, and $A_{2,1:r_{y}}^{-1}V_{0}$ and $A_{4,1:k_{y}}^{-1}V_{0}$ which necessitates solving the systems $AX=U$ and $AX=V$, where the subscripts have been omitted for simplicity.
-This is typically achieved through an LU factorization of the operator.
-With 1D finite differences, $A$ has a tridiagonal structure, which can be efficiently factorized using techniques such as the Thomas algorithm, resulting in a complexity of $\mathcal{O}(N)$ per column of $X$ [28].
+Constructing the generalized Krylov basis as outlined in lines (4) and (4) of Algorithm 4 requires performing the operations $A_{1,1:r_{x}}^{-1}U_{0}$ and $A_{3,1:k_{x}}^{-1}U_{0}$, and $A_{2,1:r_{y}}^{-1}V_{0}$ and $A_{4,1:k_{y}}^{-1}V_{0}$, which require solving the systems $AX=U$ and $AX=V$ (subscripts have been omitted for simplicity).
+With 1D finite-differences, $A$ has a tridiagonal structure, which can be efficiently factorized using direct-solver techniques such as the Thomas algorithm, with $\mathcal{O}(N)$ complexity per column of $X$ [34].
 The same complexity holds for applying $P_{1}^{-1}$ and $P_{2}^{-1}$.
-Additionally, applying $P_{1}U_{0}$, $\Phi_{1,3}U_{0}$, $\Sigma_{1,3}U_{0}$, $P_{2}V_{0}$, $\Phi_{2,4}V_{0}$, and $\Sigma_{2,4}V_{0}$, can be done with $\mathcal{O}(Nr_{0})$ computational complexity for sparse diagonal and tridiagonal matrices, leading to a total complexity of $\mathcal{O}(N\mathcal{R}_{\text{ranks}}r_{0})$ to form the extended Krylov subspaces.
-Regarding truncated orthonormalization in Algorithm 1, using a modified Gram-Schmidt method to compute the reduced QR decomposition of $U_{1}$ and $V_{1}$, in line (1), requires a complexity of $\mathcal{O}\left(Nr_{\kappa}^{2}\right)$.
-The SVD decomposition of the small-sized upper triangular matrix in line (1) requires complexity $\mathcal{O}\left(r_{\kappa}^{3}\right)$.
-The basis update in line (1) requires a complexity of $\mathcal{O}\left(Nr_{\kappa}r\right)$, where ${{r}}\leq r_{\kappa}$.
-Hence, the complexities of lines (3), (3) in Algorithm 3 scale as $\mathcal{O}\left(Nr_{\kappa}^{2}\right)$.
-Generally, we expect $r_{\kappa}\ll N$ when the solution exhibits a low-rank structure.
+Additionally, applying $P_{1}U_{0}$, $\Phi_{1,3}U_{0}$, $\Sigma_{1,3}U_{0}$, $P_{2}V_{0}$, $\Phi_{2,4}V_{0}$, and $\Sigma_{2,4}V_{0}$, can be done with $\mathcal{O}(Nr^{\ast}_{0})$ computational complexity for sparse diagonal and tridiagonal matrices, leading to a total complexity of $\mathcal{O}(N(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0})$ to augment the xKrylov subspaces.
+The use of Algorithm 2 to compute the orthonormal truncated bases $U_{1}$ and $V_{1}$ incurs a complexity of $\mathcal{O}\left(N\,\max\left[\left(2+\mathcal{R}_{\text{ranks}}\right)\,r_{0}^{\ast}\,r^{(m-1)},\;\left(\left(2+\mathcal{R}_{\text{ranks}}\right)\,r_{0}^{\ast}\right)^{2}\right]\right)\sim\mathcal{O}(Nr^{2})$.
+Generally, we expect $r$ to be much smaller and independent of $N$ when the solution is sufficiently smooth and exhibits a low-rank structure.
 Step K2
-The application of the operator using matrix-matrix multiplication and summation in the Krylov method (GMRES), as outlined in Section, 3.3, requires a complexity of $\mathcal{O}(r^{3})$.
-Furthermore, applying the preconditioner involves solving a Sylvester equation of size $r$, which can be achieved using the Bartels-Stewart algorithm [11] with a complexity of $\mathcal{O}(r^{3})$.
-Hence, the $S$-step has $\mathcal{O}(r^{3})$ complexity.
+The application of the reduced GSE operator using matrix-matrix multiplication and summation in the Krylov method (GMRES) features a complexity of $\mathcal{O}\left(\tilde{r}^{d+1}\right)$.
+Applying the ACS preconditioner requires solving an SE of size $r_{x}\times r_{y}$, which can be achieved using the Bartels-Stewart algorithm [15] with a complexity of $\mathcal{O}(\tilde{r}^{d+1})$.
+Hence, the $S$-step has $\mathcal{O}(r^{d+1})$ complexity.
 Step K3
-The SVD decomposition of ${\widetilde{S}}$ in line (2) of Algorithm 2 is of $\mathcal{O}(r^{3})$, also independent of $N$.
-Updating the bases in lines (2) and (2) of Algorithm 2 requires $\mathcal{O}(Nr{r_{1}})$, where ${r_{1}}\leq r$ is the post-truncation rank of the new solution as described in Algorithm 2.
+The SVD decomposition of ${\widetilde{S}}$ in line (3) of Algorithm 3 is of $\mathcal{O}(\tilde{r}^{d+1})$ complexity.
+Updating the bases in lines (3) and (3) of Algorithm 3 requires $\mathcal{O}(N\tilde{r}{r_{1}})$ operations, where ${r_{1}}\leq\min(r_{x},r_{y})$ is the post-truncation rank of the new solution from Algorithm 3.
 Step K4
-The QR decomposition to obtain $R_{U}$ and $R_{V}$ in lines (3) of Algorithm 3 necessitates a complexity of $\mathcal{O}\left(N(\mathcal{R}_{\text{ranks}}r_{1})^{2}\right)$ The residual computation in line (3) is of $\mathcal{O}({r_{1}}^{3})$ (i.e., independent of $N$) due to the matrix multiplications, with an additional $\mathcal{O}({r_{1}}^{2})$ for the Frobenius norm.
-We note that henceforth $r$ will represent the maximum of $r_{\kappa}$ and $\mathcal{R}_{\text{ranks}}r_{1}$.
-We conclude that the overall algorithm scales as $\mathcal{{O}}(Nr^{2}+r^{3})$, which motivates keeping the overall rank $r$ as low as possible.
-These scalings will be verified later in this study.
-3.7 Extension to High-order Temporal Discretization
-The DIRK method features a stage-by-stage backward Euler-like implicit solve, with explicit evaluation of the terms on the right side of the previous Runge-Kutta stages [10].
-In Appendix C.1, the mathematical formulation of the DIRK schemes is elaborated leading to the following residual equation for the $k$-th RK stage,
-${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\quad k=1,2,\ldots,s.$ (32)
-Expanding the operator $\mathscr{L}$ in the residual equation above results in the following $k$-th stage generalized Sylvester equation:
-$\mathbf{F}^{(k)}-a_{k,k}\Delta t\left(T_{1}\mathbf{F}^{(k)}\Phi_{2}{{}^{\top}}+\Phi_{1}\mathbf{F}^{(k)}T_{2}^{\top}+T_{3}\mathbf{F}^{(k)}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}^{(k)}T_{4}^{\top}\right)=\mathbf{B}^{(k)},\quad k=1,2,\ldots,s.$ (33)
-For the DIRK scheme, we broadly adopt the strategy outlined in [10] with minor modifications to improve efficiency.
-In particular, we: (i) construct the basis using the time-step of the first stage and retain it for subsequent stages, (ii) evolve the projected system for the $S_{1}$ matrix across the DIRK stages, and (iii) truncate and evaluate the residual to determine whether to admit the solution or return to step (i) to further augment the basis.
-Key modifications vs.
-[10] include truncating the solution before the residual computation, and only checking the residual at the end of the DIRK step.
-The reduced matrix is solved similarly to the backward Euler method outlined above, using preconditioned GMRES, where the preconditioner is based on an approximate operator, $\tilde{\mathscr{L}}$.
-Further details on the discretization of the advection-diffusion equation using the DIRK scheme are provided in Appendix C.
-Specifically, Appendix C.2 describes the adaptive-rank approach for high-order integration, Appendix C.3 extends the integration to multi-rank coefficients with DIRK schemes and presents a pseudo-algorithm.
+The QR decomposition to obtain $R_{U}$ and $R_{V}$ in lines (11)-(4) of Algorithm 4 features a complexity of $\mathcal{O}\left(N\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{2}\right)$.
+The residual computation in line (4) is of $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d+1}\right)$ due to the matrix multiplications, with an additional $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d}\right)$ for the Frobenius norm.
+It follows that, as advertised, the overall complexity of the algorithm scales as $Nr^{2}$ and as $r^{d+1}$, with $d$ the dimensionality, which will be numerically verified in Section 4.
 4 Numerical Experiments
-We demonstrate the computational complexity estimates and the overall efficacy of the proposed algorithm for advection-diffusion equations with several numerical examples.
-We assume homogeneous Dirichlet boundary conditions for all the tests for simplicity.
-All simulations were performed using MATLAB.
-For benchmarking purposes, we compare against full-rank algorithm using the Kronecker formulation inverted with MATLAB’s backslash command.
+We demonstrate the overall efficacy of the proposed algorithm with several 2D and 3D numerical examples.
+We assume homogeneous Dirichlet boundary conditions for simplicity; other boundary conditions (.e.g, periodic, Neumann, etc.) can be easily incorporated in the formulation through the operators $T_{1:4}$ (10).
+All simulations are performed using MATLAB.
+For efficient computational handling in 3D, we employ tools from the Tensor Toolbox [3] and the Htucker package [25].
 Throughout this section, we denote the diffusion Courant number and the advection Courant-Friedrichs-Lewy (CFL) number by $\lambda_{D}$ and $\lambda_{A}$, respectively, defined as:
@@
 where $\phi_{\text{max}}$ and $\sigma_{\text{max}}$ are the maximum of the diffusion and advection coefficients, respectively.
-We set algorithmic tolerances as follows.
-We first choose the residual tolerance, $\epsilon_{\texttt{tol}}$.
-One strategy to prevent rapid growth of the extended Krylov subspaces, especially when the matrix equations contain many terms, is to choose the Krylov basis truncation threshold as $\epsilon_{\kappa}\geq\epsilon_{\texttt{tol}}$ to truncate basis vectors that are nearly linearly dependent.
-However, over-truncation may risk removing important information, slowing down convergence.
-The GMRES tolerance $\epsilon_{\texttt{GMRES}}$ is chosen as $\epsilon_{\texttt{GMRES}}<10^{-2}\epsilon_{\texttt{tol}}$ to ensure that the error in the inner loop does not pollute the outer loop.
-The SVD truncation threshold $\epsilon$ are set to $\epsilon<10^{-2}\epsilon_{\texttt{tol}}$.
-Since truncation occurs before residual evaluations, the error introduced by solution truncation must remain below the residual tolerance.
-Otherwise, the computed solution may meet the residual threshold, but over-truncation could prevent this acceptance.
-Example 4.1: Advection-diffusion with a rank-1 advection coefficient and a rank-3 diffusion coefficient
-Consider Equation (3) with a rank-1 variable advection coefficient:
-$\sigma^{x}(x,y)$ $=\sigma^{1}_{1}(x)\sigma^{2}_{1}(y),$ $\sigma^{1}_{1}(x)$ $=1-x^{2},$ $\sigma^{2}_{1}(y)$ $=2y,$
-$\sigma^{y}(x,y)$ $=\sigma^{1}_{2}(x)\sigma^{2}_{2}(y),$ $\sigma^{1}_{2}(x)$ $=-2x,$ $\sigma^{2}_{2}(y)$ $=1-y^{2},$
+These numbers are measures of the condition number of the associated GSEs.
+For all simulations presented in this section, we did not need to set a maximum number of outer xKrylov iterations because the algorithm consistently converged within a small number of iterations.
+Example 4.1: 3D Advection-Diffusion
+We consider a rank-1 variable advection coefficient in a 3D domain $x,y,z\in[-1,1]$:
+$\sigma^{x}(x,y,z)$ $=\sigma^{1}_{1}(x)\sigma^{2}_{1}(y)\sigma^{3}_{1}(z),$ $\sigma^{1}_{1}(x)$ $=1-x^{2},$ $\sigma^{2}_{1}(y)$ $=2y,$ $\sigma^{3}_{1}(z)$ $=-2z$
+$\sigma^{y}(x,y,z)$ $=\sigma^{1}_{2}(x)\sigma^{2}_{2}(y)\sigma^{3}_{2}(z),$ $\sigma^{1}_{2}(x)$ $=-2x,$ $\sigma^{2}_{2}(y)$ $=1-y^{2},$ $\sigma^{3}_{2}(z)$ $=2z$
+$\sigma^{z}(x,y,z)$ $=\sigma^{1}_{3}(x)\sigma^{2}_{3}(y)\sigma^{3}_{3}(z),$ $\sigma^{1}_{2}(x)$ $=4x,$ $\sigma^{2}_{2}(y)$ $=2y,$ $\sigma^{3}_{3}(z)$ $=1-z^{2}.$
 and a rank-3 variable diffusion coefficient:
-$\phi^{x}(x,y)=\phi^{y}(x,y)=\sum_{i=1}^{3}\phi^{1}_{i}(x)\phi^{2}_{i}(y),$
-${\phi}^{1}_{1}(x)$ $=\exp\left(-\left(x-0.3\sin(x)\right)^{2}\right),$ ${\phi}^{2}_{1}(y)$ $=\exp\left(-\left(y-0.3\cos(y)\right)^{2}\right),$
-${\phi}^{1}_{2}(x)$ $=\exp\left(-\left(x-0.6\sin(\pi x)\right)^{2}\right),$ ${\phi}^{2}_{2}(y)$ $=\exp\left(-\left(y-0.6\sin(\pi y)\right)^{2}\right),$
-${\phi}^{1}_{3}(x)$ $=\exp\left(-\left(x-0.6\sin(2\pi x)\right)^{2}\right),$ ${\phi}^{2}_{3}(y)$ $=\exp\left(-\left(y-0.6\sin(2\pi y)\right)^{2}\right).$
+$\phi^{x}(x,y,z)=\phi^{y}(x,y,z)=\sum_{i=1}^{3}\phi^{1}_{i}(x)\phi^{2}_{i}(y)\phi^{3}_{i}(z),$
+${\phi}^{1}_{1}(x)$ $=\exp\left(-\left(x-0.3\sin(x)\right)^{2}\right),$ ${\phi}^{2}_{1}(y)$ $=\exp\left(-\left(y-0.3\cos(y)\right)^{2}\right),$ ${\phi}^{3}_{1}(z)$ $=\exp\left(-\left(y-0.3\sin(z)\right)^{2}\right),$
+${\phi}^{1}_{2}(x)$ $=\exp\left(-\left(x-0.6\sin(\pi x)\right)^{2}\right),$ ${\phi}^{2}_{2}(y)$ $=\exp\left(-\left(y-0.6\sin(\pi y)\right)^{2}\right),$ ${\phi}^{3}_{2}(z)$ $=\exp\left(-\left(y-0.6\sin(\pi z)\right)^{2}\right),$
+${\phi}^{1}_{3}(x)$ $=\exp\left(-\left(x-0.6\sin(2\pi x)\right)^{2}\right),$ ${\phi}^{2}_{3}(y)$ $=\exp\left(-\left(y-0.6\sin(2\pi y)\right)^{2}\right)$ ${\phi}^{3}_{3}(z)$ $=\exp\left(-\left(y-0.6\sin(2\pi z)\right)^{2}\right).$
 The initial condition is a rank-2 Gaussian:
-$f(x,y,0)=0.5\exp\left(-400\left((x-0.3)^{2}+(y-0.35)^{2}\right)\right)+0.8\exp\left(-400\left((x-0.65)^{2}+(y-0.5)^{2}\right)\right).$
-We use a spatial grid with $N_{x}=N_{y}=300$, and the time-step $\Delta t$ is set to $\lambda_{D}\Delta x^{2}$, with $\lambda_{D}$ ranging from $280$ to $750$.
+$f(x,y,z,t=0)$ $=0.5\exp\left(-400\left((x-0.3)^{2}+(y-0.35)^{2}+(z-0.2)^{2}\right)\right)$
+$+0.8\exp\left(-400\left((x-0.65)^{2}+(y-0.5)^{2}+(z-0.55)^{2}\right)\right).$
+We use a spatial grid with $N=300$ in all dimensions, and the time-step $\Delta t$ is set to $\lambda_{D}\Delta x^{2}$, with $\lambda_{D}$ ranging from $140$ to $1125$.
 We test our adaptive-rank algorithm using three different integrators: backward Euler (BE), DIRK2, and DIRK3, as detailed in Tables 3 and 3 in the Appendix.
-The residual tolerance is $\epsilon_{\texttt{tol}}=10^{-4}$ for BE, DIRK2, and DIRK3.
-The basis truncation threshold is $\epsilon_{\kappa}=10^{-3}$.
-The GMRES convergence threshold is $\epsilon_{\texttt{GMRES}}=10^{-8}$, and the SVD truncation threshold is $\epsilon=10^{-8}$.
-Figure 2 illustrates the $L_{1}$ error norm vs time stepping size characterized by $\lambda_{D}$ by showcasing the performance of the adaptive-rank integrator for BE, DIRK2, and DIRK3.
-The expected temporal error order of convergence for our proposed adaptive-rank integrators, depicted by lines with markers, is observed.
-Figure 2: Example 4.1. Temporal convergence study of BE, DIRK2, and DIRK3 schemes.
-We performed a computational complexity study to demonstrate the proposed algorithm has a linear complexity with respect to $N$, the number of grid points per dimension.
-The mesh size $N=N_{x}=N_{y}$ varied between $100$ and $31000$ and the time step was fixed at $\Delta t=10^{-3}$.
-This range of mesh sizes is chosen to prevent the $O(r^{3})$ contribution of the coefficient matrix GMRES solver from polluting the mesh-resolution scaling.
-Figure 3(a) displays $N$ vs. computational time. We observe that the complexity remains linear with respect to $N$.
-Figure 3: Example 4.1.(a) Computational complexity of DIRK2 adaptive-rank integrator. (b) Log-log plot of GMRES-ACS solver. The execution times were recorded using MATLAB’s timeit function.
-Figure 3(b) shows the computational time required for solving the projected generalized Sylvester equation via preconditioned GMRES-ACS vs.
-rank for a single time step using BE with $N=10,000$ and $\epsilon_{\texttt{GMRES}}=10^{-10}$.
-To examine the computational scaling corresponding to the expansion of the Krylov basis, we allowed the basis to grow without truncation, recording the simulation time required for the GMRES solve at each basis augmentation iteration.
-This figure verifies the expected cubic scaling, e.g.
-$\mathcal{O}(r^{3})$.
-From the above study, we verify that the predominant complexity of our approach is $\mathcal{O}(Nr^{2}+r^{3})$, effectively reducing the complexity of full rank methods, yet maintain their high-order convergence.
-We investigate next the impact of the proposed preconditioner in the S-step by comparing the performance of the GMRES solver with and without the preconditioner.
-Figure 4 showcases the GMRES residual history as a function of the number of iterations for both preconditioned and unpreconditioned cases for mesh sizes ranging from $200$ to $2500$, in a single time step with $\Delta t=0.01$.
-It is observed that preconditioning significantly accelerates convergence, with the convergence rate independent of the mesh size, which is not the case for unpreconditioned GMRES.
+The residual tolerance is $\epsilon_{\texttt{tol}}=10^{-4}$ for BE, DIRK2, and DIRK3, the basis truncation threshold is $\epsilon_{\kappa}=10^{-6}$, the SVD truncation threshold is $\epsilon=10^{-6}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-6}$.
+We investigate numerically the following: (1) high-order temporal convergence, (2) effectiveness of the xKrylov iterations in capturing the range of solution, (3) performance of the proposed ACS preconditioner, and (4) verification of the overall computational complexity.
+High-Order Temporal Convergence: Figure 1(a) illustrates the $L_{1}$ error norm vs time stepping size (characterized by $\lambda_{D}$) by showcasing the performance of the adaptive-rank integrators for BE, DIRK2, and DIRK3.
+The expected asymptotic temporal error order of convergence for our proposed adaptive-rank integrators, depicted by lines with markers, is observed.
+Additionally, the algorithm adeptly captures the rank evolution of the solution in the x, y, and z directions, as illustrated in Figure 1(b).
+Figure 1: Example 4.1. (a) Temporal convergence study of BE, DIRK2, and DIRK3 schemes. (b) Rank evolution of the solution in the $x$, $y$, and $z$ directions over time for $\lambda_{D}=140$.
+Effectiveness of the xKrylov Iterations: We evaluate the xKrylov subspace dimension required to achieve tolerances of $\epsilon_{\texttt{tol}}=10^{-4},10^{-5},10^{-6}$.
+We perform a single time-step integration with a fixed step size $\Delta t=10^{-3}$ with grid refinements in $N$.
+To isolate the effect of $\lambda_{D}$ (a measure of the GSE condition number) on the convergence behavior of xKrylov, we mitigate other potential influencing factors, such as excessive truncation of the Krylov subspace, insufficient convergence of the GMRES-ACS solver, and truncation errors in the constructed solution, by setting related thresholds to $\epsilon=\epsilon_{\texttt{GMRES}}=\epsilon_{\kappa}=10^{-10}$.
+Figure 2 illustrates the xKrylov subspace dimension required for convergence as a function of the diffusion Courant number $\lambda_{D}$ for a fixed $\Delta t$ and varying spatial resolutions $N$.
+Our results suggest that, for a fixed tolerance, the convergence of the Krylov subspaces remains largely insensitive to grid refinement across a broad range of $\lambda_{D}$.
+However, for extremely fine mesh resolutions, beyond typical practical computational needs, the convergence rate of xKrylov scales as $\lambda_{D}^{1/4}$, consistently with the $\kappa^{1/4}$ scaling (with $\kappa$ the condition number) previously reported for the Poisson equation [24, 4].
+Figure 2: Example 4.1. Convergence study using the xKrylov subspace method with DIRK3 adaptive-rank integrator for several outer basis tolerances, $\epsilon_{tol}$.
+Performance of ACS-Preconditioner: We compare the performance of the GMRES solver with and without the ACS preconditioner for the inner solve.
+Figure 3 illustrates the GMRES residual history over iterations for ACS-preconditioned and unpreconditioned cases.
+Results at the first time step ($\Delta t=0.01$) and the third outer-basis augmentation are shown, comparing different mesh sizes.
+The results confirm that ACS-preconditioned GMRES delivers mesh-independent and super-linear convergence rates, which is not the case for unpreconditioned GMRES.
 Figure 3: Example 4.1.
-Semi-log plot of GMRES residuals versus iterations with and without preconditioning.
-The simulation is conducted for a single time step with $\Delta t=0.01$, on meshes of size $N_{x}=N_{y}=200$ in (a), $N_{x}=N_{y}=1000$ in (b), and $N_{x}=N_{y}=2500$ in (c).
-Example 4.2: Adaptive Rank Evolution with Manufactured Solution
-We assess next the performance of the proposed method to capture the rank adaptively using a manufactured solution defined as:
-$f(x,y,t)$ $=\left(\sin(\pi x)\sin(\pi y)\right)\exp\left[-0.1\pi(t-2)^{2}\right]$
-$\quad+\left(\sin(2\pi x)\sin(2\pi y)\right)\exp\left[-0.5\pi(t-10)^{2}\right]$
-$\quad+\left(e^{\left[-100(x-0.3)^{2}-100(y-0.3)^{2}\right]}+e^{\left[-100(x+0.3)^{2}-100(y+0.3)^{2}\right]}\right)e^{\left[-0.5\pi(t-15)^{2}\right]}.$
-This solution is initially rank-9 and fluctuates in rank over time, with significant changes at $t=1$, $10$, and $15$.
-Based on this solution, we analytically compute the source term $S(x,y,t)$.
-The diffusion coefficients are set as previously described in Example 4.1.
-We use a spatial grid of $N=1000$, the BE integrator and simulate until $T_{f}=18$.
-The relative residual tolerance is $\epsilon_{\texttt{tol}}=10^{-8}$.
-The GMRES convergence threshold is set to $\epsilon_{\texttt{GMRES}}=10^{-12}$.
-The singular-value truncation threshold is set to $\epsilon=10^{-5}$.
-The Krylov basis truncation threshold is $\epsilon_{\kappa}=10^{-5}$.
-We use a time step $\Delta t=\lambda_{D}\Delta x^{2}$ with $\lambda_{D}=1300$, or equivalently, $\Delta t=\lambda_{A}\Delta x$ with $\lambda_{A}=17$.
-Figure 5(a) tracks the adaptive rank in time with our proposed adaptive rank integrator (in dashed red) vs the full-rank integrator (in solid blue).
-Agreement is excellent between the two.
-Figure 5(b) illustrates the evolution of the extended Krylov subspace over time, before and after applying the truncation step in Algorithm 3.
-Our approach effectively filters out low-relevance basis vectors, maximizing computational efficiency.
-The extended Krylov basis-augmentation history is shown in the semi-transparent red bars in Figure 5(b) (corresponding to right y-axis).
-It can be seen that the algorithm performs several ($\sim 3$) iterations of basis augmentation around the $t=1$, $10$, and $15$, to find an appropriate subspace to resolve the rank growth induced by the source term, settling into $\sim 1$ iteration otherwise.
-This demonstrates the effectiveness of the Krylov subspace formulation proposed here to find suitable low-rank subspaces as allowed by the dynamics of the solution.
+Log-log plots of GMRES residuals versus iterations, with and without preconditioning.
+Simulations are conducted for a single time step with $\Delta t=0.01$ at the third outer-basis augmentation, using grid resolutions $N_{x}=N_{y}=N_{z}=200$ in (a), $1000$ in (b), and $2500$ in (c).
+Figure 4 (a) illustrates the average condition number of the Galerkin-projected diffusion and advection operators as the dimension of the proposed xKrylov subspaces increases across 12 outer iterations.
+For these results, the grid size is fixed at $N=1000$ in 3D, and a single time step of $\Delta t=0.01$ is used.
+Figures 4 (b) - (d) show the GMRES residual histories for the ACS-preconditioned and unpreconditioned cases, corresponding to the second, fourth, and twelfth outer xKrylov iterations.
+These correspond to xKrylov subspace dimensions of $r=37$, $r=73$, and $r=217$, respectively.
+As the dimension of the xKrylov subspace grows, unpreconditioned GMRES convergence for the projected problem becomes progressively slower, as expected.
+In contrast, the ACS-preconditioner effectively maintains rapid convergence rates independent of subspace size $r$.
 Figure 4: Example 4.1.
-(a) Rank evolution of the adaptive-rank method using the backward Euler temporal update with a rank-3 manufactured solution.
-(b) Left axis: evolution of the extended Krylov subspace size over time, both before and after truncation.
-Right axis: number of Krylov basis iterations required for solution convergence.
-Example 4.3: Swirling Deformation Flow
-We consider solving the following advection-diffusion equation:
-$f_{t}-\left[(1-x^{2})\cdot 2y\cdot f\right]_{x}+\left[2x\cdot(1-y^{2})\cdot f\right]_{y}=\nu\Delta f,\quad x,y\in[-1,1]$ (35)
-The initial condition is a bi-Gaussian function given by:
-$f(x,y,0)=0.5\exp\left(-\frac{1}{2\sigma^{2}}\left[(x-0.3)^{2}+(y-0.35)^{2}\right]\right)+0.8\exp\left(-\frac{1}{2\sigma^{2}}\left[(x-0.65)^{2}+(y-0.5)^{2}\right]\right),$
+(a) Average condition number for Galerkin-projected advection and diffusion operators.
+(b)-(d) Log-log plots of GMRES residuals versus iterations, with and without preconditioning for various subspace sizes $r_{\kappa}=37$, $r_{\kappa}=73$, and $r_{\kappa}=217$, respectively.
+These cases correspond to the highlighted blue data points in (a).
+Simulations are conducted for a single time step with $\Delta t=0.01$ and a grid of $N_{x}=N_{y}=N_{z}=1000$.
+Verification of the Computational Complexity scaling: We verify the $\mathcal{O}(N)$ computational complexity claim next.
+We vary $N$ from $100$ to $30000$ and the time step is fixed at $\Delta t=10^{-3}$.
+Such range of mesh sizes is chosen to prevent the $O(r^{4})$ contributions of the ACS-GMRES solver from polluting the mesh-resolution scaling.
+Figure 5(a) displays $N$ vs.
+computational time for $10$ time steps of integration.
+We observe that the complexity remains linear with respect to $N$.
+The finite y-intercept corresponds to the $\mathcal{O}(r^{4})$ cost of the GMRES-ACS solve in 3D.
+Figure 5(b) displays the computational time required for the preconditioned GMRES-ACS solver vs.
+rank for a single time step, using BE with $N=10,000$ and $\epsilon_{\texttt{GMRES}}=10^{-10}$.
+For this test, we allowed the basis to grow without truncation, recording the simulation time required for the GMRES solve at each basis augmentation iteration.
+The expected $\mathcal{O}(r^{4})$ scaling in 3D is found.
+These studies verify that the complexity scaling of our approach is $\mathcal{O}(Nr^{2}+r^{4})$ in 3D.
+Figure 5: Example 4.1.
+(a) Demonstration of ${\cal O}(N)$ computational complexity of DIRK3 adaptive-rank integrator in 3D.
+(b) Demonstration of ${\cal O}(r^{4})$ computational complexity of the GMRES-ACS solver in 3D.
+The execution times were recorded using MATLAB’s timeit function.
+Example 4.2: 3D Swirling Deformation Flow
+We consider next a 3D swirling deformation flow:
+$f_{t}+\left[(1-x^{2})\cdot 2y\cdot-2z\cdot f\right]_{x}+\left[2x\cdot(1-y^{2})\cdot 2z\cdot f\right]_{y}+\left[-2x\cdot 2y\cdot 2(1-z^{2})\cdot f\right]_{z}=\nu\Delta f,\quad x,y,z\in[-1,1].$ (35)
+We set the diffusion coefficient $\nu=1/500$. The initial condition is a bi-Gaussian function given by:
+$f(x,y,0)$ $=0.5\exp\left(-\frac{1}{2\sigma^{2}}\left[(x-0.3)^{2}+(y-0.35)^{2}+(z-0.3)^{2}\right]\right)$
+$+0.8\exp\left(-\frac{1}{2\sigma^{2}}\left[(x+0.5)^{2}+(y+0.5)^{2}+(z+0.5)^{2}\right]\right),$
 where $\sigma=0.15$.
-We choose the diffusion coefficient $\nu=10^{-3}$, small enough to weaken the diffusion process.
-To satisfy the cell-Reynolds number (to preserve monotonicity while keeping the spatial discretization linear; see Appendix B), we set $\Delta x=9.52\times 10^{-4}$, resulting in a mesh with $N_{x}=N_{y}=2100$.
-This test showcases the capability of our proposed low-rank integrator to handle fine meshes, ensuring that the numerical solution adheres to the required monotonicity criteria and avoids unphysical behavior.
-The residual tolerance is set to $\epsilon_{\texttt{tol}}=10^{-5}$, the basis truncation threshold is set to $\epsilon_{\kappa}=10^{-3}$, the SVD truncation threshold is set to $\epsilon=10^{-5}$, and the GMRES acceptance threshold is set to $\epsilon_{\texttt{GMRES}}=10^{-10}$.
+We choose $N_{x}=N_{y}=N_{z}=1000$.
+The outer basis residual tolerance is $\epsilon_{\texttt{tol}}=10^{-5}$, the basis truncation threshold is $\epsilon_{\kappa}=10^{-5}$, the SVD truncation threshold is $\epsilon=10^{-5}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-10}$.
+This example demonstrates the algorithm’s ability to adapt to dynamic rank evolution.
+The algorithm performs nine xKrylov augmentations at the initial time step, two augmentations at the third and seventh time steps, and a single augmentation otherwise.
+Figure 6 shows rank evolution.
+The solution rank rises sharply in the $x$ and $y$ directions [Figures 6 (a)-(b)].
+The solution rank in the $z$ direction evolves more gently [Figure 6 (c)] and is roughly one-third of that in the $x$ and $y$ directions.
+Figures 6 (d)-(e) showcase how the solution deforms according to the prescribed advection coefficients and diffuses slowly over time.
+Figure 7 shows snapshots of the solution in the $x\text{--}z$ plane at $y=-0.5$.
+As the flow-driven deformation increases the complexity of the solution structure, the rank grows to capture these features; when the diffusion process starts to dominate, the rank decreases as the solution smooths.
+This example shows that the low-rank assumption is not intrinsic to our proposed algorithm.
+Even when the rank becomes moderately large, the algorithm remains capable of capturing the underlying dynamics effectively.
+Naturally, the benefit of a low-rank approximation decreases as the solution approaches a full-rank structure.
 Figure 6: Example 4.2.
-Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=7$ on a $2100\times 2100$.
-(a) Evolution of rank over time.
-(b)–(e) Solution profiles at $t=0$, $t=2.8487$, $t=5.7088$, and $t=9.9962$.
+3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid.
+(a)–(c) Evolution of rank, xKrylov subspace before and after truncation, and number of xKrylov iterations over time for the $x$, $y$, and $z$ directions, respectively.
+(d)–(e) Solution profiles at $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$.
 The final simulation time is $T_{f}=10$.
-Figure 6 demonstrates our algorithm’s ability to capture the solution rank for this challenging case.
-The solution rotates according to the prescribed advection coefficients and diffuses slowly over time.
-As the solution structure becomes more complex due to deformation, the rank increases to accurately represent its physical features.
-Notably, the algorithm dynamically adjusts the solution rank, which does not exceed $r_{1}=20$, two orders of magnitude smaller than the grid dimension $N=2100$.
-Additionally, we evaluate the algorithm’s effectiveness in dynamically capturing the solution rank, even with large time steps.
-To this end, we use the proposed adaptive-rank algorithm with DIRK3 temporal integration and employ subsequent larger time steps, $\Delta t$, corresponding to CFL numbers $\lambda_{A}=1$ and $\lambda_{A}=15$.
-All other parameters are kept consistent with the simulation depicted in Figure 6, specifically $\epsilon_{\kappa}=10^{-3}$, $\epsilon_{\texttt{tol}}=10^{-5}$, $\epsilon=10^{-5}$, and $\epsilon_{\texttt{GMRES}}=10^{-10}$.
-Figure 7 shows the evolution of the generalized Krylov subspace over time, both before and after truncation, highlighting significant reductions in basis size and thus algorithmic complexity.
-The figure also demonstrates that the algorithm effectively captures the solution rank evolution even with fairly large (but dynamically relevant) advective CFLs at comparable performance.
-This demonstrates the algorithm’s capability to construct an appropriate Krylov basis, accurately evaluate the effects of the advection and diffusion operators, and dynamically truncate the subspace as needed.
 Figure 7: Example 4.2.
-Evolution of the Extended Krylov subspace over time before and after truncation for different $\lambda_{A}$ values using DIRK3: (a) and (b) correspond to $\lambda_{A}=1$, while (c) and (d) correspond to $\lambda_{A}=15$.
-Example 4.4: Constructed Steady-State Solution
-In this example, we construct diffusion and advection coefficients that balance each other, resulting in a non-trivial steady-state solution.
-Specifically, we seek an equilibrium solution satisfying the steady-state condition:
-$0=\nabla\cdot\left(\mathbf{\Phi}\cdot\nabla f_{\text{eq}}-\mathbf{\Sigma}f_{\text{eq}}\right),$ (36)
+3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid.
+(a)–(f) snapshots of the solution in the $x\text{--}z$ plane at $y=-0.5$, corresponding to times $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$, respectively.
+Example 4.3: A 2D Steady-State Balanced-Advection-Diffusion Problem
+The last example prescribes diffusion and advection coefficients that exactly balance each other, resulting in a non-trivial steady-state solution.
+We seek an equilibrium solution satisfying the steady-state condition:
+$0=\nabla\cdot\left(\bm{\mathit{\Phi}}\cdot\nabla f_{\text{eq}}-\bm{\mathit{\Sigma}}f_{\text{eq}}\right),$ (36)
 a solution of which is:
-$\mathbf{\Sigma}=\mathbf{\Phi}\cdot\nabla\log(f_{\text{eq}}).$
-This equation defines the steady-state flow from a provided diffusion coefficient and a steady-state solution. We choose the steady-state solution in the domain $0\leq x,y\leq 1$ to be:
+$\bm{\mathit{\Sigma}}=\bm{\mathit{\Phi}}\cdot\nabla\log(|f_{\text{eq}}|).$
+This equation defines the steady-state advective flow $\bm{\mathit{\Sigma}}$ for given diffusion coefficient $\bm{\mathit{\Phi}}$ and steady-state solution $f_{\text{eq}}$.
+We construct a steady-state solution in the domain $0\leq x,y\leq 1$ to be (up to an arbitrary constant):
 $f_{\text{eq}}(x,y)=\left(x(1-x)\right)^{2}\left(y(1-y)\right)^{2},$
 which satisfies homogeneous Dirichlet boundary conditions. The diffusion coefficients are defined as:
-$\phi^{x}(x,y)=\phi^{y}(x,y)$ $=\phi^{1}(x)\phi^{2}(y),$
-$\phi^{1}(x)$ $=x^{2}(1-x)^{2},$
-$\phi^{2}(y)$ $=y^{2}(1-y)^{2}.$
-The variable advection coefficients are then given by:
-$\sigma^{x}(x,y)=\sigma^{1}(x)\sigma^{2}(y),\qquad D_{y}(x,y)=\chi^{1}(x)\chi^{2}(y),$
-where:
-$\sigma^{1}(x)$ $=2x(1-3x+2x^{2}),\qquad\sigma^{2}(y)=y^{2}(1-y)^{2},$
-$\chi^{1}(x)$ $=x^{2}(1-x)^{2},\qquad\chi^{2}(y)=2y(1-3y+2y^{2}).$
-We select the following rank-1 initial condition:
-$f(x,y,0)=\left|\sin(2\pi x)\sin(2\pi y)\right|,\quad 0\leq x,y\leq 1.$
-Since the solution will converge to a steady state up to a constant, we correct the mass of the initial condition to match that of the steady-state solution.
-To correct the mass of the solution, we adopt the LoMaC strategy outlined in [14, 13, 12], as implemented in [10].
-Figure 8 shows the error of the numerical steady-state solution relative to the prescribed analytical steady state as a function of mesh size.
-The advection CFL number, $\lambda_{A}$, for a fixed time step $\Delta t=1000$ ranges from $19,050$ to $577,500$ as the mesh is refined, indicating the exceptionally large time steps allowed by the algorithm.
-The results confirm that second-order spatial convergence is achieved and demonstrate that, under these conditions, large time steps can be taken while still accurately reaching the correct steady state with asymptotic convergence.
+$\phi^{x}(x,y)=\phi^{y}(x,y)$ $=\phi^{1}(x)\phi^{2}(y),\qquad\phi^{1}(x)=x^{2}(1-x)^{2},\qquad\phi^{2}(y)=y^{2}(1-y)^{2}.$
+The variable advection coefficients are $\sigma^{x}(x,y)=\sigma_{1}^{1}(x)\sigma_{1}^{2}(y)$, $\sigma^{y}(x,y)=\sigma_{2}^{1}(x)\sigma_{2}^{2}(y),$ where:
+$\sigma_{1}^{1}(x)$ $=2x(1-3x+2x^{2}),\qquad\sigma_{1}^{2}(y)=y^{2}(1-y)^{2},\qquad\sigma_{2}^{1}(x)=x^{2}(1-x)^{2},\qquad\sigma_{2}^{2}(y)=2y(1-3y+2y^{2}).$
+We select the following rank-1 initial condition: $f(x,y,0)=\left|\sin(2\pi x)\sin(2\pi y)\right|$.
+The outer basis residual tolerance is $\epsilon_{\texttt{tol}}=10^{-3}$, the basis truncation threshold is $\epsilon_{\kappa}=10^{-8}$, the SVD truncation threshold is $\epsilon=10^{-8}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-8}$.
+Figure 8 shows the error of the numerical steady-state solution, reached using ten time steps, relative to the prescribed analytical steady state as a function of mesh size.
+The advection CFL number, $\lambda_{A}$, for a fixed time step $\Delta t=1000$ ranges from $1,190$ to $36,076$ as the mesh is refined.
+The results are obtained by rescaling the numerical solution so that its normalization constant is the same as the analytical one.
+They confirm second-order spatial convergence and demonstrate the ability of our algorithm to take large timesteps to reach the correct steady state with expected asymptotic accuracy.
 Figure 8: Example 4.3.
 Log-log plot of error in the numerical steady-state solution compared to the prescribed analytical steady state as a function of mesh size.
-The advection CFL number, $\lambda_{A}$, ranges from $19,050$ to $577,500$.
-The plot demonstrates the retrieval of second-order spatial convergence in the steady state.
+The advection CFL number, $\lambda_{A}$, ranges from $1,190$ to $36,076$.
+The plot shows second-order spatial convergence in the steady state.
 5 Conclusion
 We have proposed a computationally efficient adaptive-rank implicit algorithm for solving time-dependent advection-diffusion partial differential equations with variable coefficients.
-We discretize the PDE using second-order centered finite differences in space and implicitly in time with up to third-order accuracy using DIRK, resulting in a generalized Sylvester matrix equation.
-Our algorithm is based on several key strategies: constructing dimension-wise basis from extended Krylov subspaces, projecting the original matrix equation onto a reduced system to solve for a coefficient matrix, devising an effective Averaged-Coefficient Sylvester (ACS) preconditioner for the reduced system, and a computationally efficient procedure for the residual evaluation.
-This combined approach achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{3})$, with $r$ being the rank of the solution during Krylov iteration process, effectively mitigating the curse of dimensionality and making it comparable in efficiency to constant-coefficient cases [10].
-We have verified our algorithm through a series of numerical experiments.
-Overall, our adaptive-rank algorithm offers a scalable and efficient solution for time-dependent advection-diffusion type PDEs with variable coefficients.
-It effectively balances computational cost and accuracy by efficiently constructing and truncating the Krylov basis and dynamically adjusting the solution rank.
-The ability to handle large meshes and time steps without sacrificing accuracy makes it a promising tool for a wide range of applications.
+Our algorithm is based on several key strategies: constructing dimension-wise basis from xKrylov subspaces, projecting the original matrix equation onto a reduced system to solve for a coefficient matrix, devising an effective Averaged-Coefficient Sylvester (ACS) preconditioner for the reduced system, and a computationally efficient procedure for the residual evaluation.
+This combined approach achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$, with $r$ being the rank of the solution during Krylov iteration process and $d$ the dimensionality, effectively mitigating the curse of dimensionality and making it comparable in efficiency to the constant-coefficient case [14].
+Such complexity scaling is numerically verified with our extensive tests, with the maximal rank $r$ found to be independent of the 1D grid size $N$ for practical grid resolutions and convergence tolerances.
 Future research directions include extending the algorithm to higher dimensions and nonlinear PDEs.
@@
 J.Q.
-was supported by Air Force Office of Scientific Research FA9550-22-1-0390, Department of Energy DE-SC0023164 and Air Force Office of Scientific Research (AFOSR) FA9550-24-1-0254 via the Multidisciplinary University Research Initiatives (MURI) Program.
+was also supported by Air Force Office of Scientific Research FA9550-22-1-0390, Department of Energy DE-SC0023164 and Air Force Office of Scientific Research (AFOSR) FA9550-24-1-0254 via the Multidisciplinary University Research Initiatives (MURI) Program.
 H.E.
@@
 Appendix A Discretization of the Diffusion Operator
-We provide a detailed derivation of the matrix formulation of the discretization of the diffusion operator for the two-dimensional advection-diffusion equation with separable diffusion coefficients.
-We consider the discretization of the diffusion term in the $x$-direction, with other directions treated similarly.
-Consider the diffusion term in the $x$-direction:
+We derive the matrix formulation of the discretization of the diffusion operator for the two-dimensional advection-diffusion equation with separable diffusion coefficients.
+We consider the discretization of the diffusion term in the $x$-direction, with other directions treated similarly:
 $\frac{\partial}{\partial x}\left(\phi^{x}(x,y)\frac{\partial u(x,y)}{\partial x}\right),$ (37)
@@
 $\phi^{x}(x,y)=\phi^{1}(x)\phi^{2}(y).$ (38)
-Our goal is to discretize this term using a 3-point stencil on a uniform grid with spacing $\Delta x$ in the $x$-direction.
+We discretize this term using a 3-point stencil on a uniform grid with spacing $\Delta x$ in the $x$-direction.
 Let $u_{i,j}\approx u(x_{i},y_{j})$, with $i=0,\cdots,N_{x}-1$ and $j=0,\cdots,N_{y}-1$.
-We approximate the spatial derivative using central differences. The discrete approximation at grid point $(i,j)$ is:
-$\left[\frac{\partial}{\partial x}\left(\phi^{x}\frac{\partial u}{\partial x}\right)\right]_{i,j}\approx\frac{1}{\Delta x}\left[\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i+\frac{1}{2},j}-\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i-\frac{1}{2},j}\right],$ (39)
-where the fluxes at the cell faces are approximated by:
-$\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i+\frac{1}{2},j}$ $\approx\phi^{x}_{i+\frac{1}{2},j}\cdot\frac{u_{i+1,j}-u_{i,j}}{\Delta x},$ (40)
-$\left(\phi^{x}\frac{\partial u}{\partial x}\right)_{i-\frac{1}{2},j}$ $\approx\phi^{x}_{i-\frac{1}{2},j}\cdot\frac{u_{i,j}-u_{i-1,j}}{\Delta x}.$ (41)
-Here $\phi^{x}_{i+\frac{1}{2},j}=\phi^{1}_{i+\frac{1}{2}}\phi^{2}_{j}=\phi^{1}(x_{i+\frac{1}{2}})\phi^{2}(y_{j}).$ Substituting back into Eq. (39), we obtain:
-$\left[\frac{\partial}{\partial x}\left(\phi^{x}\frac{\partial u}{\partial x}\right)\right]_{i,j}\approx\frac{1}{\Delta x^{2}}\left[\phi^{1}_{i+\frac{1}{2}}(u_{i+1,j}-u_{i,j})-\phi^{1}_{i-\frac{1}{2}}(u_{i,j}-u_{i-1,j})\right]\phi^{2}_{j}.$ (42)
-To express the discretization in matrix form, we define:
+We approximate the spatial derivative using central differences.
+The discrete approximation at grid point $(i,j)$ is:
+$\left[\frac{\partial}{\partial x}\left(\phi^{x}\frac{\partial u}{\partial x}\right)\right]_{i,j}\approx\frac{1}{\Delta x^{2}}\left[\phi^{1}_{i+\frac{1}{2}}(u_{i+1,j}-u_{i,j})-\phi^{1}_{i-\frac{1}{2}}(u_{i,j}-u_{i-1,j})\right]\phi^{2}_{j},$ (39)
+with $\phi^{1}_{i+\frac{1}{2}}\phi^{2}_{j}=\phi^{1}(x_{i+\frac{1}{2}})\phi^{2}(y_{j}).$ To express the discretization in matrix form, we define:
   • $\mathbf{F}$: matrix of solution values, $[\mathbf{F}]_{i,j}=f_{i,j}$.
@@
   • $\Phi_{2}$: diagonal matrix with elements $\left[\Phi_{2}\right]_{j,j}=\phi^{2}_{j}$.
-We also define the forward and backward difference matrices in the $x$-direction for the interior nodes $i=1,\cdots,N_{x}-2$, where $i=0$ and $i=N_{x}-1$ correspond to zero Dirichlet boundary conditions and rest at the boundary:
-$D_{+}$ $=\frac{1}{\Delta x}\begin{pmatrix}-1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\\ 0&\cdots&0&0&-1\end{pmatrix},$ (43)
-$D_{-}$ $=\frac{1}{\Delta x}\begin{pmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\end{pmatrix}.$ (44)
-Using these definitions, the discrete diffusion term in Eq. (42) can be represented in the matrix form as:
+We also define the forward and backward difference matrices in the $x$-direction for the interior nodes $i=1,\cdots,N_{x}-2$.
+The nodes $i=0$, $i=N_{x}-1$ rest at the boundary and enforce homogeneous Dirichlet boundary conditions (other boundary conditions can be readily considered):
+$D_{+}=\frac{1}{\Delta x}\begin{pmatrix}-1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\\ 0&\cdots&0&0&-1\end{pmatrix},D_{-}=\frac{1}{\Delta x}\begin{pmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\end{pmatrix}.$ (40)
+Using these definitions, the discrete diffusion term in Eq. (39) can be represented in matrix form as:
 $\left(\Phi^{1}_{+}D_{+}-\Phi^{1}_{-}D_{-}\right)\mathbf{F}{\Phi_{2}^{\top}}.$ (41)
-We note that the operator multiplying $\mathbf{F}$ from the left is analogous to $D^{\Phi}_{xx}$ in Equation (6). The discretization in the y-direction follows the same strategy.
+We note that the operator multiplying $\mathbf{F}$ from the left is analogous to $D^{\Phi}_{xx}$ in (9). The discretization in the $y$-direction follows the same strategy.
 Appendix B Discretization of the Advection Operator
-We provide a detailed derivation of the matrix formulation of the discretization of the advection operator for the two-dimensional advection-diffusion equation with variable, separable advection coefficients.
+We derive the matrix formulation of the discretization of the advection operator for the two-dimensional advection-diffusion equation with variable, separable advection coefficients.
 As before, we focus on the discretization of the advection term in the $x$-direction:
@@
 $\sigma^{x}(x,y)=\sigma^{1}(x)\sigma^{2}(y).$ (43)
-Our goal is to discretize this term on a uniform grid with spacing $\Delta x$ in the $x$-direction and $\Delta y$ in the $y$-direction.
-Let $u_{i,j}\approx u(x_{i},y_{j})$.
+We discretize this term on a uniform grid with spacing $\Delta x$ in the $x$-direction and $\Delta y$ in the $y$-direction.
 To preserve linearity of our discretization without introducing too much numerical dissipation, we employ a central difference scheme to approximate the spatial derivative.
-This discretization is not monotonicity-preserving unless the so-called cell Reynolds number condition is satisfied:
-$\text{Re}_{c}=\frac{\sigma_{\text{max}}\Delta x}{\phi_{\text{max}}}<2,$ (48)
-where $\sigma_{\text{max}}$ is the maximum magnitude of the advection velocity, and $\phi_{\text{max}}$ is the maximum diffusion coefficient.
+This discretization is monotonicity-preserving when the following condition is satisfied:
+$\Delta x<2\frac{\phi_{\text{max}}}{\sigma_{\text{max}}},$ (44)
+where $\sigma_{\text{max}}$ is the maximum magnitude of the advection velocity and $\phi_{\text{max}}$ is the maximum diffusion coefficient.
 This condition is practical in our formulation due to the very favorable scaling of our low-rank representation with mesh size and rank.
-The discrete approximation at grid point $(i,j)$ is:
-$\left[\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right)\right]_{i,j}\approx\frac{1}{\Delta x}\left[\left(\sigma^{x}u\right)_{i+\frac{1}{2},j}-\left(\sigma^{x}u\right)_{i-\frac{1}{2},j}\right],$ (49)
-where the fluxes at the cell faces are approximated by:
-$\left(\sigma^{x}u\right)_{i+\frac{1}{2},j}$ $\approx\sigma^{1}_{i+\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i+1,j}+u_{i,j}}{2},$ (50)
-$\left(\sigma^{x}u\right)_{i-\frac{1}{2},j}$ $\approx\sigma^{1}_{i-\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i,j}+u_{i-1,j}}{2}.$ (51)
-Here, $\sigma^{1}_{i+\frac{1}{2}}=\sigma^{1}(x_{i+\frac{1}{2}})$ is the advection coefficient evaluated at the midpoint between $x_{i}$ and $x_{i+1}$, and similarly for $\sigma^{1}_{i-\frac{1}{2}}$.
-Substituting back into the discrete approximation, we have:
-$\left[\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right)\right]_{i,j}$ $\approx\frac{1}{\Delta x}\left[\sigma^{1}_{i+\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i+1,j}+u_{i,j}}{2}-\sigma^{1}_{i-\frac{1}{2}}\sigma^{2}_{j}\cdot\frac{u_{i,j}+u_{i-1,j}}{2}\right]$ (52)
-$=\frac{1}{2\Delta x}\left[\sigma^{1}_{i+\frac{1}{2}}(u_{i+1,j}+u_{i,j})-\sigma^{1}_{i-\frac{1}{2}}(u_{i,j}+u_{i-1,j})\right]\sigma^{2}_{j}.$ (53)
+The discrete approximation at the grid point $(i,j)$ is:
+$\left[\frac{\partial}{\partial x}\left(\sigma^{x}(x,y)u(x,y)\right)\right]_{i,j}$ $\approx\frac{1}{2\Delta x}\left[\sigma^{1}_{i+\frac{1}{2}}(u_{i+1,j}+u_{i,j})-\sigma^{1}_{i-\frac{1}{2}}(u_{i,j}+u_{i-1,j})\right]\sigma^{2}_{j},$ (45)
+where $\sigma^{1}_{i+\frac{1}{2}}=\sigma^{1}(x_{i+\frac{1}{2}})$ is the advection coefficient evaluated at the midpoint between $x_{i}$ and $x_{i+1}$, and similarly for $\sigma^{1}_{i-\frac{1}{2}}$, and $\sigma^{2}_{j}$ is evaluated at $y_{j}$.
 To express the discretization in matrix form, we define:
@@
 We also define the difference matrices in the $x$-direction:
-$S_{+}$ $=\frac{1}{2}\begin{pmatrix}1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\\ 0&\cdots&0&0&1\end{pmatrix},$ (54)
-$S_{-}$ $=\frac{1}{2}\begin{pmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\end{pmatrix}.$ (55)
+$S_{+}=\frac{1}{2}\begin{pmatrix}1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\\ 0&\cdots&0&0&1\end{pmatrix},S_{-}=\frac{1}{2}\begin{pmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\end{pmatrix}.$ (46)
 Using these definitions, the discrete advection term in the $x$-direction can be represented as:
@@
 C.1 DIRK Scheme
-DIRK methods are characterized by the Butcher tableau, as shown in Table 1.
+DIRK methods are characterized by the Butcher tableau, Table 1.
 Table 1: Butcher tableau for an $s$-stage DIRK scheme.
@@
 We focus on stiffly accurate DIRK methods, where the coefficients satisfy $b_{i}=a_{si}$ for $i=1,\ldots,s$, ensuring that the final solution update corresponds to the solution at the last DIRK stage.
-For instance the Butcher tables for DIRK2 and DIRK3 are presented in Tables 3 and 3, respectively.
+The Butcher tables for DIRK2 and DIRK3 are presented in Tables 3 and 3, respectively.
+Table 2: DIRK2 Butcher table, $\gamma=1-\frac{\sqrt{2}}{2}$.
 $\gamma$ $\gamma$ 0
@@
 $1-\gamma$ $\gamma$
-Table 2: DIRK2 Butcher table with $\gamma=1-\frac{\sqrt{2}}{2}$.
+Table 3: DIRK3 Butcher table, $x=0.4358665215$.
 $x$ $x$ 0 0
@@
 $-\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $-\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
-Table 3: DIRK3 Butcher table with $x=0.4358665215$.
-Similarly to the backward Euler case, the matrix-based discretization of the advection-diffusion equation using the operator $\mathscr{L}$ defined in (8) results in a generalized Sylvester equation that must be solved at each DIRK stage.
+Similarly to the backward Euler case, the matrix-based discretization of the advection-diffusion equation using the operator $\mathscr{L}$ defined in (11) results in a GSE that must be solved at each DIRK stage.
 The corresponding DIRK scheme for the matrix differential equation (11) from $t^{(0)}$ to $t^{(1)}$ is written as:
@@
 $\begin{split}\mathbf{F}_{1}&=\mathbf{F}^{(s)}=\mathbf{F}_{0}+\Delta t\sum\limits_{k=1}^{s}b_{k}\mathbf{Y}_{k}.\end{split}$ (48c)
-More compactly, within each stage $k$ of the DIRK scheme, one must minimize the residual equation (32) given by:
-${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\quad k=1,2,\ldots,s.$
-As shown in Equation (33), the residual equations lead to a generalized Sylvester equation to be solved at each stage.
+Within each stage $k$ of the DIRK scheme, one must solve the residual equation given by:
+${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\quad k=1,2,\ldots,s.$ (49)
+Expanding the operator $\mathscr{L}$ in the residual equation above results in the following $k$-th stage GSE:
+$\mathbf{F}^{(k)}-a_{k,k}\Delta t\left(T_{1}\mathbf{F}^{(k)}\Phi_{2}{{}^{\top}}+\Phi_{1}\mathbf{F}^{(k)}T_{2}^{\top}+T_{3}\mathbf{F}^{(k)}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}^{(k)}T_{4}^{\top}\right)=\mathbf{B}^{(k)},\quad k=1,2,\ldots,s.$ (50)
+For the DIRK update, we adopt the strategy outlined in [14], with minor modifications to improve efficiency.
+In particular, we: (i) construct the basis using the time-step of the first stage and retain it for subsequent stages, (ii) evolve the projected system for the $S_{1}$ matrix across the DIRK stages, and (iii) truncate and evaluate the residual to determine whether to admit the solution or return to step (i) to further augment the basis.
+Key modifications vs.
+[14] include truncating the solution before the residual computation, and only checking the residual at the end of the DIRK step.
+The reduced matrix is solved using preconditioned GMRES, as before.
 C.2 Adaptive-Rank Strategy for High-Order Temporal Discretization via DIRK
@@
 […]
-$\mathbf{R}^{(k)}_{\tilde{\mathscr{L}}}=\mathbf{F}_{1}-a_{kk}\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell}.$
-Expanding this residual using $\tilde{\mathscr{L}}$, defined in (14), leads to the $k$th-stage Sylvester equation:
+$\mathbf{R}^{(k)}_{\tilde{\mathscr{L}}}=\mathbf{F}_{1}-a_{kk}\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell}.$ (51)
+Expanding this residual using $\tilde{\mathscr{L}}$, defined in (16), leads to the $k$th-stage SE:
 $\underbrace{\left[\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\alpha_{2}T_{1}\right)}_{A_{1}}+\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\gamma_{2}T_{3}\right)}_{A_{3}}\right]\mathbf{F}^{(k)}}_{P_{1}}+\mathbf{F}^{(k)}\underbrace{\left[\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\alpha_{1}T_{2}\right)^{\top}}_{A^{\top}_{2}}+\underbrace{\left(\frac{1}{4}I-a_{kk}\Delta t\,\gamma_{1}T_{4}\right)^{\top}}_{A^{\top}_{4}}\right]}_{P_{2}^{T}}=\mathbf{B}^{(k)}.$ (52)
 Note that the operators $A_{1:4}$ are fixed throughout the stages since the DIRK method uses constant diagonal Butcher tableau coefficients.
-We follow the strategy discussed in [10] to perform high-order integration.
-The key idea is to consider the basis fixed across DIRK stages, and only evolve the matrix of coefficients ${S}_{1}$ through the stages.
+We follow [14] to perform high-order integration.
+The key idea is to consider the basis fixed across DIRK stages, and only evolve the matrix of coefficients ${{\bf S}}_{1}$ through the stages.
 The proposed procedure is as follows:
@@
     1. (a)
-      Solve reduced Sylvester equation for $S^{(k)}$ using preconditioned GMRES,
+      Solve reduced SE for ${\bf S}^{(k)}$ using preconditioned GMRES,
       ${\bf S}^{(k)}-\Delta t\left(\tilde{T}_{1}{\bf S}^{(k)}\tilde{\Phi}^{\top}_{2}+\tilde{\Phi}_{1}{\bf S}^{(k)}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}^{(k)}\tilde{\Sigma}^{\top}_{2}+\tilde{\Sigma}_{1}{\bf S}^{(k)}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}^{(k)},$ (53)
@@
       where $\tilde{P}_{1}=U_{1}^{T}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{T}P_{1}V_{1}$, and is valid for all stages since the basis constructed is fixed throught the DIRK stages by assumption.
-Solve for $S^{(k)}$ and obtain the intermediate RK solutions as $U_{1}S^{(k)}V_{1}^{\top}$.
+Solve for ${\bf S}^{(k)}$ and obtain the intermediate RK solutions as $U_{1}S^{(k)}V_{1}^{\top}$.
 To further improve computational efficiency, from (53), we have,
-      $\tilde{Y}_{\ell}$ $=\frac{1}{a_{\ell\ell}\Delta t}\left({S^{(\ell)}}-{\tilde{B}^{(\ell)}}\right),$
-      leading to an efficient computation of $\tilde{B}^{(k)}$:
-      $\tilde{B}^{(k)}=\tilde{B}_{1}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({S^{(\ell)}}-{\tilde{B}^{(\ell)}}\right),$ (61)
-      with $\tilde{B}_{1}$ defined in (19). This avoids the need of evaluating the full size $B^{(k)}$ in (32).
+      $\tilde{Y}_{\ell}$ $=\frac{1}{a_{\ell\ell}\Delta t}\left({{\bf S}^{(\ell)}}-{\tilde{{\bf B}}^{(\ell)}}\right),$
+      leading to an efficient computation of $\tilde{{\bf B}}^{(k)}$:
+      $\tilde{{\bf B}}^{(k)}=\tilde{{\bf B}}_{1}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({{\bf S}^{(\ell)}}-{\tilde{{\bf B}}^{(\ell)}}\right),$ (55)
+      with $\tilde{{\bf B}}_{1}$ defined in (21). This avoids the need to evaluate the full size ${\bf B}^{(k)}$ in (51).
   3. Step 3.
-    Perform truncation of evolved solution, i.e. $\{U_{1},S_{1},V_{1}^{\top}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}^{\top}\})$, with $S_{1}:=S_{1}^{(s)}$.
+    Perform truncation of evolved solution, i.e. $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}^{\top}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}^{\top}\})$, with $S_{1}:=S_{1}^{(s)}$.
   4. Step 4.
     Efficiently evaluate the residual norm,
-    $\left\|{{\bf R}^{(s)}_{\mathscr{L}}}\right\|=\left\|{R}_{U}{\text{diag}\left(S_{1}-\tilde{B}^{(s)}_{1},-a_{ss}\Delta t\left(I_{{\mathcal{R}}_{\text{ranks}}}\otimes S_{1}\right)\right)}{R}_{V}^{\top}\right\|,$
-    where $R_{U}$ and $R_{V}$ are upper triangular matrices from a reduced QR decomposition as described in (23), which can be done once for all DIRK stages.
-  5. 1.
-    Compare the computed residual norm $\|{\bf R}^{(s)}_{\mathscr{L}}\|$ with the given error tolerance $\epsilon_{tol}$.
-If the tolerance is met, then admit the solution with ${\bf F}_{1}=U_{1}S_{1}V_{1}^{\top}$; otherwise, we go back to Step 1 to further augment Krylov subspaces.
+    $\left\|{{\bf R}^{(s)}_{\mathscr{L}}}\right\|=\left\|{R}_{U}{\text{diag}\left(-\tilde{B}^{(s)}_{1},S_{1,\epsilon},-a_{ss}\Delta t\left(I_{{\mathcal{R}}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)}{R}_{V}^{\top}\right\|,$
+    where $R_{U}$ and $R_{V}$ are upper triangular matrices from a reduced Q-less QR decomposition as described in (24), which can be done once for all DIRK stages.
+  5. Step 5.
+    Compare the computed residual norm $\|{\bf R}^{(s)}_{\mathscr{L}}\|$ with the given error tolerance $\epsilon_{\texttt{tol}}$.
+If the tolerance is met, then admit the solution with ${\bf F}_{1}=U_{1,\epsilon}S_{1,\epsilon}V_{1,\epsilon}^{\top}$; otherwise, we go back to Step 1 to further augment the xKrylov subspaces.
 C.3 DIRK Generalization to Multi-Rank Advection-Diffusion Coefficients
-We now extend the formulation to the multi-rank case, as described in (4). Using a DIRK method for temporal discretization, the $k$-th stage generalized Sylvester equation becomes:
+We now extend the formulation to the multi-rank case, as described in (7). Using a DIRK method for temporal discretization, the $k$-th stage GSE becomes:
 $\mathbf{F}^{(k)}-a_{kk}\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}^{(k)}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}^{(k)}{T_{2,j}}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}^{(k)}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}^{(k)}{T_{4,l}}^{\top}}_{\mathscr{L}(\mathbf{F}^{(k)})}\right)=\mathbf{B}^{(k)},\quad k=1,2,\ldots,s$ (56)
@@
 It also generalizes the backward Euler special case with a single stage and $a_{1,1}=1$, yielding $\mathbf{B}^{(1)}=\mathbf{F}_{0}$, consistently with (13).
-Using approximations in (25), we arrive at the $k$th stage approximated Sylvester equation:
+Using approximations in (62), we arrive at the $k$th stage approximated SE:
 […]
@@
 $P_{2}$ $=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{4,l}}=\sum_{j=1}^{\ell_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-a_{kk}\Delta t\alpha_{y,j}T_{2,j}\right)}+\sum_{l=1}^{k_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-a_{kk}\Delta t\gamma_{y,l}T_{4,l}\right)}.$
-In order to obtain a reduced equation for the $k$th stage matrix of coefficients $S^{(k)}$, we perform a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$ to obtain :
+In order to obtain a reduced equation for the $k$th stage matrix of coefficients ${\bf S}^{(k)}$, we perform a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$ to obtain :
 ${\bf S}^{(k)}-a_{kk}\Delta_{t}\left(\sum_{i=1}^{\ell_{x}}\tilde{T}_{1,i}{\bf S}^{(k)}\tilde{\Phi}_{i}^{2,x\top}+\sum_{j=1}^{\ell_{y}}\tilde{\Phi}_{j}^{1,y}{\bf S}^{(k)}\tilde{T}_{2,j}^{\top}+\sum_{k=1}^{k_{x}}\tilde{T}_{3,k}{\bf S}^{(k)}\tilde{\Sigma}_{k}^{2,x\top}+\sum_{l=1}^{k_{y}}\tilde{\Sigma}_{l}^{1,y}{\bf S}^{(k)}\tilde{T}_{4,l}^{\top}\right)=\tilde{{\bf B}}^{(k)}$ (58)
@@
 $\underbrace{\left[\sum_{i=1}^{\ell_{x}}{\tilde{A}_{1,i}}+\sum_{k=1}^{k_{x}}{\tilde{A}_{3,k}}\right]}_{\tilde{P}_{1}}\mathbf{S}^{(k)}+\mathbf{S}^{(k)}\underbrace{\left[\sum_{j=1}^{\ell_{y}}{\tilde{A}_{2,j}}+\sum_{l=1}^{k_{y}}{\tilde{A}_{4,l}}\right]}_{\tilde{P}_{2}^{\top}}=\tilde{{\bf B}}^{(k)}$ (59)
-Using the operators $\tilde{P}_{1}$ and $\tilde{P}_{2}$, we define the preconditioner mapping:
+Using the operators $\tilde{P}_{1}$ and $\tilde{P}_{2}$, we define the ACS-preconditioner mapping:
 ${\bf S}\mapsto\tilde{P}_{1}{\bf S}+{\bf S}\tilde{P}_{2}^{T},$ (60)
-Algorithm 4 solves the variable advection diffusion equation using DIRK methods.
+Algorithm 5 solves the time-dependent variable advection-diffusion equation using DIRK methods.
 Input: Initial condition matrices $U_{0}$, $V_{0}$, $S_{0}$;
@@
 Time step size $\Delta t$;
-Tolerances $\epsilon_{\kappa}$, $\epsilon$, ${\epsilon}_{\text{tol}}$;
-Maximum iterations max_iter;
+Tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, ${\epsilon}_{\text{tol}}$;
 Output: Updated bases $U_{1}$, $V_{1}$;
@@
 1 Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eq. (57);
-2
-3for $m=1$ to max_iter do
+2 Compute $l=2+\ell_{x}+\ell_{y}+k_{x}+k_{y}$;
+3 Set $U^{(i)}=U_{0}$ and $V^{(i)}=V_{0}$ for $i=1,\dots,l$;
+4 Set $U_{1}=U_{0}\quad V_{1}=V_{0}$;
+5 while not converged do
 6        // Step 1.
-5        Truncate to tolerance $\epsilon_{\kappa}$;
-6        $U_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right)\right)$;
-7        $V_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right)\right)$;
-8        Compute $\tilde{B}^{(1)}=(U_{1}^{\top}U_{0})S_{0}(V_{1}^{\top}V_{0})^{\top}$;
+7        Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
+8        $U_{1},\{U^{(1)}\cdots U^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}}\};\{U^{(1)}\cdots U^{(l)}\};U_{1};\epsilon_{\kappa}\right)$;
+9        $V_{1},\{V^{(1)}\cdots V^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}}\};\{V^{(1)}\cdots V^{(l)}\};V_{1};\epsilon_{\kappa}\right)$;
 10        // Step 2.
 11        for $k=1$ to $s$ do
-11               Compute $\tilde{B}^{(k)}=\tilde{B}^{(1)}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left(S^{(\ell)}-\tilde{B}^{(\ell)}\right)$;
-12               Solve the reduced equation (65) for $S^{(k)}$ using GMRES with ACS preconditioner (66) to tolerance $\epsilon_{\texttt{GMRES}}$;
-13               Compute and store $\frac{1}{a_{kk}}\left(S^{(k)}-\tilde{B}^{(k)}\right)$ and proceed to the next stage;
-14
-15       Truncate $\{U_{1},S_{1},V_{1}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S^{(s)},V_{1}\})$ to tolerance $\epsilon$;
+12               Compute $\tilde{{\bf B}}^{(k)}=\tilde{{\bf B}}^{(1)}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({\bf S}^{(\ell)}-\tilde{{\bf B}}^{(\ell)}\right)$;
+13               Solve the reduced equation (59) for ${\bf S}^{(k)}$ using GMRES with ACS preconditioner (60) to tolerance $\epsilon_{\texttt{GMRES}}$;
+14               Compute and store $\frac{1}{a_{kk}}\left({\bf S}^{(k)}-\tilde{{\bf B}}^{(k)}\right)$ and proceed to the next stage;
+15       Truncate $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S^{(s)},V_{1}\})$ to tolerance $\epsilon$;
 16        // Step 3.
-17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},T_{1,1:\ell_{x}}U_{1},\Phi_{1:\ell_{y}}^{1,y}U_{1},T_{3,1:k_{x}}U_{1},\Sigma_{1:k_{y}}^{1,y}U_{1}]\right)$ and
-18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},\Phi_{1:\ell_{x}}^{2,x}V_{1},T_{2,1:\ell_{y}}V_{1},\Sigma_{1:k_{x}}^{2,x}V_{1},T_{4,1:k_{y}}V_{1}]\right)$;
+17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
+18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
 19        // Step 4.
-20        Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(S_{1}-\tilde{B}^{(s)},-a_{ss}\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1}\right)\right)R_{V}^{\top}\right\|$;
+20        Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(-\tilde{B}^{(s)},S_{1,\epsilon},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)R_{V}^{\top}\right\|$;
 21
@@
 25       else
+26               Set $U_{1}\leftarrow U_{1,\epsilon}$; $U_{1}\leftarrow V_{1,\epsilon}$; $S_{1}=S_{1,\epsilon}$;
 27               Break;
 28               Exit loop;
-28
 29
 30
-Algorithm 4 s-Stages Adaptive-Rank DIRK Integrator for the Generalized Sylvester Equation.
-00footnotetext: The MATLAB implementation of the algorithm is publicly available on GitHub, Adaptive-rank-ACS, https://github.com/helkahza1/Adaptive-rank-ACS
+Algorithm 5 s-Stages Adaptive-Rank DIRK Integrator for the GSE.
+Appendix D Three-Dimensional Formulation
+We outline our approach to 3D next. In 3D, we replace the SVD algorithm with a tensor Tucker decomposition, specifically tailored for the three (and higher) dimension tensor-product grids.
+D.1 Tucker Tensor Equations
+We discretize the 3D time-dependent advection-diffusion equation on a three-dimensional tensor-product spatial grid with $N_{1}$, $N_{2}$, and $N_{3}$ grid points in the $x$, $y$, and $z$ directions, respectively.
+Here and in the following, we let ${\bf F}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$ denote a three-dimensional tensor.
+The mode-$n$ product of ${\bf F}$ with a matrix $A_{n}\in\mathbb{R}^{I\times N_{n}}$, for $n=1,2,3$, is defined as:
+${\bf F}\times_{1}A_{1}\in\mathbb{R}^{I\times N_{2}\times N_{3}},\,{\bf F}\times_{2}A_{2}\in\mathbb{R}^{N_{1}\times I\times N_{3}},\,{\bf F}\times_{3}A_{3}\in\mathbb{R}^{N_{1}\times N_{2}\times I}.$
+The resulting tensor in each case has its $n$-th dimension replaced by the number of rows of $A_{n}$. The entries of the mode-$n$ product are computed as:
+$({\bf F}\times_{n}A_{n})_{i_{1},\ldots,i_{n-1},j,i_{n+1},\ldots,i_{3}}=\sum_{i_{n}=1}^{N_{n}}{\bf F}_{i_{1},\ldots,i_{n},\ldots,i_{3}}\cdot(A_{n})_{j,i_{n}}.$
+For a more thorough review of tensor operations, we refer the reader to [22].
+The discretization of the 3D advection-diffusion equation using finite-differences in space combined with the backward Euler method in time, yields the following GSE:
+${\bf R}_{\mathscr{L}}=\mathbf{F}_{1}-\Delta t\mathscr{L}({\bf F}_{1})-\mathbf{F}_{0}=\mathbf{0},$ (61)
+where ${\bf F}_{0}$ denotes the initial tensor condition at time $t_{0}$, and ${\bf F}_{1}$ is the sought tensor solution at time $t_{1}$, and,
+$\mathscr{L}:\mathbf{F}\mapsto$ $\sum_{i=1}^{\ell_{x}}\mathbf{F}\times_{1}T_{1,i}\times_{2}\Phi_{i}^{2,x}\times_{3}\Phi_{i}^{3,x}+\sum_{i=1}^{\ell_{y}}\mathbf{F}\times_{1}\Phi_{i}^{1,y}\times_{2}T_{2,i}\times_{3}\Phi_{i}^{3,y}+\sum_{i=1}^{\ell_{z}}\mathbf{F}\times_{1}\Phi_{i}^{1,z}\times_{2}\Phi_{i}^{2,z}\times_{3}T_{3,i}$
+$+$ $\sum_{i=1}^{k_{x}}\mathbf{F}\times_{1}T_{4,i}\times_{2}\Sigma_{i}^{2,x}\times_{3}\Sigma_{i}^{3,x}+\sum_{i=1}^{k_{y}}\mathbf{F}\times_{1}\Sigma_{i}^{1,y}\times_{2}T_{5,i}\times_{3}\Sigma_{i}^{3,y}+\sum_{i=1}^{k_{z}}\mathbf{F}\times_{1}\Sigma_{i}^{1,z}\times_{2}\Sigma_{i}^{2,z}\times_{3}T_{6,i}.$
+The operators $T$ correspond to the composition of advection and diffusion difference operators with diagonal matrices, analogous to those defined in Section 2.
+Similarly, the operators $\Phi$ and $\Sigma$ are analogous to the diagonal matrices for diffusion and advection coefficients, respectively, used for pointwise multiplication, as defined in Section 2.
+To obtain an averaged (approximate) mapping for preconditioning purposes, we define the average coefficients:
+$\alpha_{i}^{2,x}$ $=\text{avg}(\Phi_{i}^{2,x}),\,\alpha_{i}^{3,x}$ $=\text{avg}(\Phi_{i}^{3,x}),\,\alpha_{i}^{1,y}$ $=\text{avg}(\Phi_{i}^{1,y}),\,\alpha_{i}^{3,y}$ $=\text{avg}(\Phi_{i}^{3,y}),\,\alpha_{i}^{1,z}$ $=\text{avg}(\Phi_{i}^{1,z}),\,\alpha_{i}^{2,z}$ $=\text{avg}(\Phi_{i}^{2,z}),$ (62)
+$\gamma_{i}^{2,x}$ $=\text{avg}(\Sigma_{i}^{2,x}),\,\gamma_{i}^{3,x}$ $=\text{avg}(\Sigma_{i}^{3,x}),\,\gamma_{i}^{1,y}$ $=\text{avg}(\Sigma_{i}^{1,y}),\,\gamma_{i}^{3,y}$ $=\text{avg}(\Sigma_{i}^{3,y}),\,\gamma_{i}^{1,z}$ $=\text{avg}(\Sigma_{i}^{1,z}),\,\gamma_{i}^{2,z}$ $=\text{avg}(\Sigma_{i}^{2,z}).$
+which yields the three-dimensional approximate operator $\tilde{\mathscr{L}}$ as:
+[…]
+$+\mathbf{F}\times_{2}\bigg{(}\sum_{j=1}^{\ell_{y}}\alpha_{j}^{1,y}\alpha_{j}^{3,y}T_{2,j}+\sum_{l=1}^{k_{y}}\gamma_{l}^{1,y}\gamma_{l}^{3,y}T_{5,l}\bigg{)}$
+$+\mathbf{F}\times_{3}\bigg{(}\sum_{k=1}^{\ell_{z}}\alpha_{k}^{1,z}\alpha_{k}^{2,z}T_{3,k}+\sum_{l=1}^{k_{z}}\gamma_{l}^{1,z}\gamma_{l}^{2,z}T_{6,l}\bigg{)}.$ (63)
+Using this approximate operator, and letting ${\mathcal{R}_{\text{ranks}}}=r_{x}+r_{y}+r_{z}+\ell_{x}+\ell_{y}+\ell_{z}$, we then define the averaged three-dimensional approximate SE of the form:
+${\bf F}_{1}\times_{1}P_{1}+{\bf F}_{1}\times_{2}P_{2}+{\bf F}_{1}\times_{3}P_{3}={\bf F}_{0},$ (64)
+with
+$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\alpha_{i}^{2,x}\alpha_{i}^{3,x}T_{1,i}\right)}_{{A_{1,i}}}+\sum_{k=1}^{k_{x}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\gamma_{k}^{2,x}\gamma_{k}^{3,x}T_{4,k}\right)}_{{A_{4,k}}}=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{4,k}},$
+$P_{2}$ $=\sum_{j=1}^{\ell_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\alpha_{j}^{1,y}\alpha_{j}^{3,y}T_{2,j}\right)}_{A_{2,i}}+\sum_{l=1}^{k_{y}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{l}^{1,y}\gamma_{l}^{3,y}T_{5,l}\right)}_{A_{5,i}}=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{5,l}},$
+$P_{3}$ $=\sum_{j=1}^{\ell_{z}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{3}}-\Delta t\alpha_{k}^{1,z}\alpha_{k}^{2,z}T_{3,k}\right)}_{A_{3,i}}+\sum_{l=1}^{k_{z}}\underbrace{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{l}^{1,z}\gamma_{l}^{2,z}T_{6,l}\right)}_{A_{6,i}}=\sum_{j=1}^{\ell_{z}}{A_{3,j}}+\sum_{l=1}^{k_{z}}{A_{6,l}}.$
+These one-dimensional matrix operators, analogous to those in (26), are used (i) to construct xKrylov subspaces for the Tucker basis factors and (ii) to precondition the solution of the three-dimensional GSE arising from the Galerkin condition on the residual in Eq.
+(61).
+D.2 GMRES-ACS Solution of the Reduced GSE in 3D
+We assume the existence of a low-rank Tucker factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}\times N_{3}}$ at time $t^{(1)}=t^{(0)}+\Delta t$, also in the low-rank Tucker format.
+This is represented as:
+[…]
+where $U_{0}\in\mathbb{R}^{N_{1}\times r^{x}_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r^{y}_{0}}$, $W_{0}\in\mathbb{R}^{N_{3}\times r^{z}_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r^{x}_{1}}$, $V_{1}\in\mathbb{R}^{N_{2}\times r^{y}_{1}}$, and $W_{1}\in\mathbb{R}^{N_{3}\times r^{z}_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r^{x}_{0}\times r^{y}_{0}\times r^{z}_{0}}$ and $S_{1}\in\mathbb{R}^{r^{x}_{1}\times r^{y}_{1}\times r^{z}_{1}}$ are core tensors at the corresponding times.
+The bases $U_{1}$, $V_{1}$, and $W_{1}$ are constructed from orthonormalization of dimension-wise xKrylov, as discussed in 3.2
+Using a Galerkin projection on the residual, $\mathbf{R}_{\mathscr{L}}\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}$, we obtain a reduced-size GSE for the core tensor ${\bf S}_{1}$ as follows:
+$\mathbf{S}_{1}-\Delta t\mathscr{L}(\mathbf{S}_{1})\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}=\mathbf{S}_{0}\times_{1}U_{1}^{\top}U_{0}\times_{2}V_{1}^{\top}V_{0}\times_{3}W_{1}^{\top}W_{0}.$ (66)
+This projected system has dimensions $r_{x}\times r_{y}\times r_{z}$, where $r_{x}$, $r_{y}$, and $r_{z}$ denote the dimensions of the truncated xKrylov subspaces in the $x$, $y$, and $z$ directions, respectively, and will be solved using the GMRES-ACS preconditioner.
+Eq. (66) is solved iteratively using GMRES, with $\mathscr{L}$-operator application cast as a series of tensor-matrix multiplications, along with the ACS preconditioning operator given by:
+$\mathcal{P}:\mathbf{S}\mapsto\mathbf{S}\times_{1}\tilde{P}_{1}+\mathbf{S}\times_{2}\tilde{P}_{2}+\mathbf{S}\times_{3}\tilde{P}_{3},$ (67)
+where $\tilde{P}_{i}=P_{i}\times_{1}U_{1}^{\top}\times_{2}V_{1}^{\top}\times_{3}W_{1}^{\top}$ denotes the projection of operators in Eq.
+(64), analogous to those introduced in Section 3.
+This operator is used to precondition each GMRES iteration of the inner system solve.
+The solution of the projected SE has a computational cost of $\mathcal{O}(r^{4})$ [9, 41].
+The tensor-times-matrix operations, executed using the ttm command in the Tensor Toolbox, incur a cost of $\mathcal{O}(r^{4})$.
+The total complexity in three dimensions is therefore $\mathcal{O}(Nr^{2}+r^{4})$.

### v3

@@
 Abstract
-We consider the adaptive-rank integration of 2D and 3D time-dependent advection-diffusion partial differential equations (PDEs) with variable coefficients.
-We employ a standard finite-difference method for spatial discretization coupled with diagonally implicit Runge-Kutta temporal schemes.
-The discrete equation is a generalized Sylvester equation (GSE), which we solve with an adaptive-rank algorithm structured around three key strategies: (i) constructing dimension-wise subspaces based on an extended Krylov strategy, (ii) developing an effective preconditioner for the reduced coefficient matrix, and (iii) efficiently computing the residual of the equation without explicitly reverting to the full-rank form.
-The low-rank decomposition is performed in 2D using SVD, and with high-order SVD (HOSVD) in 3D to represent the tensor in a compressed Tucker format.
-The computational complexity of the proposed approach is demonstrated numerically to be comparable to the constant-coefficient case [14], scaling as $\mathcal{O}(N{r^{2}}+{r^{d+1}})$ for $d$-dimensional problems (here, $d=2$ or $3$), with $N$ the resolution in one dimension and $r$ the maximal rank during the Krylov iteration (which we find to be largely independent of $N$).
-We present numerical examples that illustrate the computational efficacy and complexity of our algorithm.
+We consider the adaptive-rank integration of multi-dimensional time-dependent advection-diffusion partial differential equations (PDEs) with variable coefficients.
+We employ a standard finite-difference method for spatial discretization coupled with high-order diagonally implicit Runge-Kutta temporal schemes.
+The discrete equation is a generalized Sylvester equation (GSE), which we solve with a projection-based adaptive-rank algorithm structured around two key strategies: (i) constructing dimension-wise subspaces using a novel atypical extended Krylov strategy, and (ii) efficiently solving the basis coefficient matrix with a preconditioned GMRES solver.
+The low-rank decomposition is performed in 2D using SVD and with high-order SVD (HOSVD) in 3D to represent the tensor in a compressed Tucker format.
+For $d$-dimensional problems (here, $d=2$ or $3$), the computational complexity and memory storage of the approach are found numerically to scale as $\mathcal{O}(N{r^{2}})+\mathcal{O}({r^{d+1}})$ and $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$, respectively, with $N$ the one-dimensional resolution and $r$ the maximal rank during the Krylov iteration (which we find to be largely independent of $N$ on our numerical examples).
+We present numerical examples that illustrate the advertised properties of the algorithm.
 Keywords: Adaptive-rank, extended Krylov subspaces, implicit Runge-Kutta integrators, advection-diffusion equations, Sylvester preconditioning, matrix equations.
-1 Introduction:
-We propose an adaptive-rank numerical scheme for approximating solutions of time-dependent variable-coefficient advection-diffusion equations.
-The goal is to reduce the computational complexity of high-dimensional PDE solvers by exploiting the low-rank structure of the solution, building upon existing literature on low-rank PDE solvers [16, 13] and model order reduction (MOR) for solving high-dimensional linear and nonlinear systems [24, 16, 5, 40].
-Two primary strategies in the literature exploit evolving low-rank structures for time-dependent PDEs: the dynamical low-rank (DLR) approximation [21, 27, 8, 11, 26] and the step-and-truncate (SAT) method, which has both explicit [23, 10, 18, 17, 19] and implicit [35, 30, 14, 28, 29] variants.
-Our study employs SAT, discretizing implicitly in time with high-order accuracy, leading to matrix/tensor differential equations to be solved at every time step.
-For constant coefficients in 2D, the advection-diffusion PDE becomes a Sylvester equation (SE) of the form:
+1 Introduction
+We propose an adaptive-rank numerical scheme for approximating solutions of multi-dimensional time-dependent variable-coefficient advection-diffusion equations.
+The goal is to reduce the computational complexity of high-dimensional PDE solvers by dynamically exploiting the low-rank structure of the time-dependent solution when it exists, building upon existing literature on low-rank PDE solvers [16, 13] and model order reduction (MOR) for solving high-dimensional linear and nonlinear systems [25, 16, 5, 43].
+Two primary strategies in the literature exploit evolving low-rank structures for time-dependent PDEs: the dynamical low-rank (DLR) approximation [22, 28, 8, 11] and the step-and-truncate (SAT) method, which has both explicit [24, 10, 18, 17, 19] and implicit [38, 31, 14, 29, 30] variants.
+Recently, [13] provided an overview of these two approaches with a focus on time-dependent kinetic simulations.
+Our study employs the SAT approach, which is known for preserving the high-order spatial and temporal accuracy of traditional mesh-based solvers for time-dependent problems.
+Specifically, to capture the solution dynamics accurately, we apply a high-order diagonally implicit Runge-Kutta (DIRK) time discretization to the matrix differential equation arising from the semi-discretization of advection-diffusion equations on a tensor-product of 1D spatial grids.
+The backward Euler time discretization of a 2D constant-coefficient advection-diffusion equation yields a Sylvester equation (SE) of the form:
 $A\mathbf{X}_{1}+\mathbf{X}_{1}B^{\top}=\mathbf{X}_{0},$ (1)
-where $\mathbf{X}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ represents the solution at time $t_{1}$, $\mathbf{X}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$ is the initial condition at $t_{0}$, and $A$ and $B$ are difference operators arising from the PDE discretization.
-Here, $N_{1}$ and $N_{2}$ are the 2D grid dimensions.
-Krylov subspace methods have proven effective for solving SEs resulting from the discretization of PDEs [24, 39, 35, 31, 6].
-In these methods, the solution bases are chosen from a Krylov (or Krylov-like) subspace of $A$ and $B$ in (1), and the original system is Galerkin-projected onto this reduced subspace.
-To accelerate the convergence of iterative methods, extensions of Krylov subspaces have been developed [12] involving $A^{-1}$, $B^{-1}$, which have been successfully applied to matrix equations [39, 14].
-While substantial progress has been made for constant-coefficient problems, PDEs with variable coefficients remain challenging.
-Discretizing such systems leads to the so-called generalized Sylvester equation (GSE), of the form:
-$\sum_{i=1}^{l}A_{i}\mathbf{X}_{1}{B}^{\top}_{i}=\mathbf{X}_{0},$ (2)
-where $A_{i}$ and $B_{i}$ represent discretization operators for advection and/or diffusion terms.
-Several methods have been proposed to solve GSEs either with stationary iterative techniques  [5, 38] or Krylov methods  [5, 39, 32, 42].
-Stationary iterative methods leverage low-rank techniques to enhance efficiency, achieving computational complexity of $\mathcal{O}(N)$ when implemented effectively [38].
-However, their effectiveness critically depends on whether a dominating operator can be correctly identified to enable a contractive iterative process.
-For the Krylov approach, recent approaches leverage a Galerkin projection of the matrix equation on the dimension-wise Krylov bases when the right-hand side is low-rank [36, 40].
-Compared to a standard Kronecker Krylov iteration, dimension-wise Krylov subspaces are inherently richer and yield superior approximations [32, 42, 24].
-Several studies in the literature have considered using atypical extended Krylov subspaces to solve GSEs (see [5, 33, 20] and references therein).
-In [5], the authors present a series of techniques to tackle the Generalized Lyapunov equation, which is a special case of (2) of the form:
-$A\mathbf{X}+\mathbf{X}A^{T}+\sum_{j=1}^{m}M_{j}\mathbf{X}M_{j}^{T}=\mathbf{F}.$ (3)
-Among them, they discuss a low-rank projection-based scheme, in which extended Krylov subspaces are constructed using $A$ and standard Krylov subspaces are used for the remaining operators.
-In [20], a Krylov-based projection technique is developed for Generalized Lyapunov equations with $m=1$.
-The method builds its approximation subspaces by block augmentations generated from $A$ and $M_{1}$.
-In [33], a low-rank projection-based technique is employed for systems of disparate dimension sizes, $N_{2}\ll N_{1}$.
-The authors employ a nested sequence of block rational Krylov subspaces generated from all the operators in the $N_{1}$ direction, and solve the resulting small-sized equation with either matrix-based Conjugate Gradient or a Kronecker formulation (tractable since $N_{2}$ is small).
-In all of these studies, atypical approximation subspaces are formed by augmenting the basis with multiple applications of the operators in the multi‑term matrix equation, a strategy we adopt in this work.
-Another Krylov-based line of inquiry, particularly relevant to our study, employs a basis-update-and-Galerkin (BUG) DLR integrator as a low-rank GMRES preconditioner for the GSE arising from anisotropic diffusion time-dependent problems with variable coefficients [28].
-The BUG preconditioner evolves the basis and the coefficient matrix to provide guesses for the basis vectors and coefficient updates of the outer GMRES iteration, demonstrating significant potential in controlling GMRES iteration count and maximal Krylov rank.
-The approach is compatible with high-order temporal integrators.
-However, the reference offers no discussion of asymptotic computational complexity.
-In this study, we propose to solve the GSE for time-dependent advection-diffusion equations leveraging adaptive-rank approximations of solutions on tensor-product grids with $N$ grid points per dimension, and using extended Krylov subspaces to update the basis vectors as well as the rank (denoted as $r$).
-Our main goal is to minimize the computational cost of the adaptive rank implementation without sacrificing accuracy.
-Key innovations of our proposed algorithm include:
-  • Generalized Extended Krylov (xKrylov) Adaptive-Rank Integrators with Complexity Reduction for 2D and 3D Problems.
-We consider matrix solutions with factorizations of the form $X=USV^{\top}$, where basis vectors in $U$ and $V$ span the solution space and determine the rank of the solution.
-In the 2D setting, we employ a truncated singular value decomposition (SVD) to factorize the solution.
-In 3D, we use high-order singular value decomposition (HOSVD) to represent the 3D tensor in a compressed Tucker format.
-Our proposed approach constructs a set of xKrylov subspaces (for each dimension) that effectively capture the evolution of solution basis in an iterative manner.
-A Galerkin projection is then applied to project the full system to a reduced GSE.
-We propose a residual-checking mechanism that leverages the low-rank structure of solutions directly, thus avoiding forming the full residual matrix for norm evaluations.
-For the rest of the paper, we refer to the xKrylov iteration as the outer iteration.
-  • Effective Preconditioning Strategy: The effectiveness of these projection-based approaches depends on the availability of an effective solver for the reduced GSE (2) of size up to $r$, the maximum size of dimension-wise xKrylov subspaces.
-There are two main challenges to solve it: (i) an increased computational load due to the multi-term nature of the equation, leading to rapid increase of xKrylov subspace sizes, and (ii) slow convergence of iterative methods (e.g., GMRES) resulting from algebraic ill-conditioning when solved iteratively.
-These challenges highlight the need for an effective preconditioner.
-Inspired by [31], we propose an averaged-coefficient Sylvester (ACS) approximation operator to the GSE as a preconditioner to GMRES.
-The approximate operator is of the Sylvester form and compact-equivalent to the original GSE, ensuring excellent performance as a GMRES accelerator [1, 2].
-The direct inversion of the SE features a favorable computational complexity of $\mathcal{O}(r^{d+1})$, to be contrasted with the $\mathcal{O}(r^{3d})$ for $d$-dimensional problems when directly inverting the Kronecker-product form of the GSE.
-We refer to the preconditioned-GMRES reduced-GSE solve as the inner iteration.
-  • Efficient Computational Complexity: As proposed, our method achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$ for $d$-dimensional problems, offering a significant advantage over existing state-of-the-art adaptive-rank approaches.
-The $Nr^{2}$ term corresponds to the outer xKrylov iteration, and the $r^{d+1}$ term to the inner GMRES iteration, assuming an $r$-independent GMRES convergence rate (which we achieve with the ACS preconditioner).
-In principle, the rank $r$ may scale with $N$ [24], but we do not find this in practice.
-The favorable scaling properties of our approach will be verified in our numerical experiments.
+where $\mathbf{X}_{0},\mathbf{X}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ represent the solution snapshots at times $t_{0}$ and $t_{1}$, respectively, and $A$ and $B$ are coefficient matrices derived from the discretizations of advection and diffusion differential operators in two dimensions.
+Here $N_{1}$ and $N_{2}$ denote the number of grid points along each coordinate direction of the tensor product grid.
+Generalization to problems with variable advection and/or diffusion coefficients leads to the so-called generalized Sylvester equation (GSE) in matrix and Kronecker forms, respectively:
+$\sum_{i=1}^{l}A_{i}\mathbf{X}_{1}{B}^{\top}_{i}=\mathbf{X}_{0}\qquad\iff\qquad\underbrace{\left(\sum_{i=1}^{l}{B}_{i}\otimes A_{i}\right)}_{\mathcal{A}}\text{vec}(\mathbf{X}_{1})=\text{vec}(\mathbf{X_{0}}),$ (2)
+where $A_{i}$ and $B_{i}$ are coefficient matrices from discretization of the PDE operators, leveraging the separability of variable advection-diffusion coefficients.
+Higher-order DIRK schemes lead to SE and GSE in forms similar to (1)-(2), but with the right-hand side (RHS) term, $\mathbf{X}_{0}$, constructed from the previous DIRK stages.
+To exploit low-rank structure in matrices (for 2D) and tensors (for 3D and beyond), such RHS term can be efficiently compressed using SVD rounding in 2D, Tucker decomposition in 3D, and hierarchical Tucker or tensor train decomposition for problems beyond 3D [32, 27].
+Ultimately, a successful development of efficient and robust solvers for the GSE (2) is of critical importance to enable adaptive-rank integrators that accurately capture the transient states of time-dependent problems.
+Two broad families of algorithms are commonly employed to tackle GSEs (and SEs as a special case): projection-type methods [5, 20, 45, 36, 39, 25, 43, 41] and matrix-oriented classical Krylov methods [45, 35, 5, 34, 33, 26].
+Projection-type schemes first construct low-dimensional subspaces that capture the dominant features of the solution.
+The GSE is then projected dimension-by-dimension via a Galerkin condition, yielding a much smaller GSE for the expansion coefficients that compose the bases together.
+Matrix-oriented Krylov iterations, in contrast, construct an approximation subspace by applying powers of the operator $\mathcal{A}$ in (2) to the vectorized RHS, but exploit the Kronecker structure so that the action of $\mathcal{A}$ is performed at the matrix level, and the full Kroneckerized operator is never formed explicitly.
+Inner products and other Krylov routines are likewise executed at the matrix-level, with low-rank truncation applied throughout to control memory.
+Although an appropriately chosen projection space can make a projection-type method formally equivalent to a matrix-oriented Krylov scheme, the projection-based framework offers additional freedom to design nonstandard subspaces that may accelerate convergence [45].
+On the downside, projection methods are less amenable to precondition: their efficiency hinges on the efficacy of the chosen subspaces and on the accuracy with which the reduced problem is solved.
+Classical Krylov methods, in contrast, benefit from well-established preconditioning strategies but offer no flexibility to alter the search space, as the basis is typically generated by an Arnoldi or Lanczos-type process applied directly to $\mathcal{A}$.
+In summary, each approach has complementary strengths and weaknesses, and the choice between them should be guided by the structure of the particular GSE under investigation.
+Below we survey representative work in both directions.
+In the context of solving SEs, projection-type methods based on Krylov subspace techniques have proven highly effective, particularly by exploiting the low-rank structure of PDE solutions [25, 42, 38, 6, 14].
+These methods typically construct dimension-wise solution bases from Krylov or Krylov-like subspaces and apply a Galerkin projection onto the reduced subspace spanned by the outer product of dimension-wise Krylov subspaces [39, 25, 43].
+To accelerate convergence, extended Krylov subspace methods (first proposed in [12] for the approximation of matrix functions) that incorporate $A$ and $A^{-1}$, as well as $B$ and $B^{-1}$, in each direction, respectively, were successfully applied to matrix equations in [42, 14] within a projection-based scheme termed Krylov-Plus-Inverted-Krylov (KPIK).
+Applying projection-type methods to GSEs is much more challenging.
+The major difficulty lies in preventing the rapid growth of the solution approximation spaces due to the (possibly large) multi-term structure of the equation, which otherwise results in large reduced systems (of size $r^{d}$, where $r$ is the dimension of the per-direction Krylov subspace and $d$ is the number of dimensions).
+Such large reduced system demands efficient, tailored solution strategies.
+Generalized Lyapunov equations of the form ($A\mathbf{X}_{1}+\mathbf{X}_{1}A+\sum_{i=1}^{l}M_{i}X_{1}M_{i}=\mathbf{X}_{0}$), a special case of the GSE matrix equation (2), have received significant attention  [5, 20] exemplifying such challenges.
+In [5], projection spaces are built from extended Krylov subspaces from the operator $A$ and standard Krylov subspaces from the remaining operators ($M_{1},\cdots M_{l}$), but do not address how to solve the reduced system efficiently.
+In [20], for the case $l=1$, projection spaces are built from a rational Krylov from $A$ and inverted Krylov from $A+M_{1}$.
+The authors then use the Sherman-Woodbury formula to express the projected equation as a series of Lyapunov solves, relying on the availability of a low-rank decomposition of the Galerkin projected generalized operators in the system.
+The resulting complexity for the reduced system is $\mathcal{O}(r^{d+1})$ under the assumption of fast decaying singular values of the operators, but deteriorates to $\mathcal{O}(r^{3d})$ otherwise.
+Low-rank matrix-oriented Krylov methods such as Conjugate Gradient (CG) or GMRES [5, 26, 33, 46] provide an alternative to projection-type solvers.
+In [5], several low-rank matrix-oriented Krylov methods, including CG and BiCGSTAB, are surveyed, and a comparative numerical study against projection-based methods is presented.
+The study shows that, across a few benchmark problems, projection methods (for the specific approximations spaces utilized) underperforms preconditioned BiCGSTAB Krylov method in terms of both convergence rate and memory usage.
+In contrast, Reference  [45] shows that projection-based methods can be computationally superior to the classical conjugate gradient method when an appropriate solution subspace is chosen, and that classical Krylov methods suffer from performance degradation due to loss of orthogonality from the low-rank truncation of the subspaces.
+Reference [34] confirms the loss of orthogonality issues in low-rank Krylov subspace methods, and proposes techniques to mitigate it.
+They further find that preconditioning is key to avoid substantial subspace growth in matrix-oriented low-rank Krylov methods.
+Preconditioning is indeed key to low-rank Krylov solvers for GSEs.
+Reference [34] proposes an interesting preconditioner for GSEs stemming from variable advection–diffusion equations, which will be of particular relevance to the present study.
+By averaging the PDE coefficients—a strategy we term here the average-coefficient-Sylvester (ACS) approximation—they obtain a constant-coefficient preconditioner of an SE form.
+The GSE system is then solved using matrix-oriented GMRES (but without low-rank truncation), where the operator application exploits the Kronecker product structure of the GSE, and the preconditioner is applied using KPIK [42].
+In another related work on anisotropic diffusion problems [29], a basis-update-and-Galerkin (BUG) DLR integrator is employed as a preconditioner for a low-rank matrix-oriented GMRES solver for the associated GSE.
+The BUG preconditioner evolves the basis and the coefficient matrix to provide effective initial guesses for both the basis vectors and coefficient updates in the GMRES iteration, demonstrating significant potential in reducing the GMRES iteration count and the maximal Krylov rank.
+Therefore, the performance of (low-rank) matrix-oriented Krylov methods critically depends on the availability of a cheap and suitable preconditioner and an efficient truncation mechanism to limit the iteration count and control subspace growth.
+In this study, we build upon the advances above for SE and GSE solvers within the projection-based framework and develop an adaptive-rank implicit integrator for 2D and 3D time-dependent advection-diffusion equations with variable coefficients.
+The main innovation of the algorithm lies in the specific treatment of both the basis construction and coefficient-finding steps within a standard projection-based Krylov framework, in which we combine solution elements proposed elsewhere to produce a novel algorithm.
+Notably, from the projection-based framework, we borrow the idea of enriched atypical Krylov subspaces to ensure an efficient representation of the solution bases.
+From classical matrix-oriented Krylov methods, we adopt the idea of truncating the subspaces along the iterations to keep it as compact as possible, and the idea of preconditioning the GSE with a simpler equation that is spectrally close.
+We assemble these ingredients in the following way: For the basis-building step (which we will refer to as the outer iteration), we introduce a new Krylov subspace construction strategy that combines (i) inverted Krylov subspaces for each differentiation coefficient matrix, (ii) standard Krylov subspaces for diagonal matrices arising from separability of advection-diffusion coefficients, and (iii) extended Krylov subspaces for averaged-coefficient matrices formed from the ACS approximation.
+We control excessive growth of the rank of the so-constructed Krylov subspaces bases by a targeted truncation procedure using SVD-based QR and modified Gram-Schmidt orthogonalization.
+In the coefficient-finding step (which we refer to as the inner iteration), we follow a standard Galerkin projection and solve the associated system with GMRES, preconditioned by the SE resulting from the Galerkin projection of the ACS approximation.
+This preconditioner is expected to perform optimally because it is compact equivalent to the original system [1, 2] and therefore spectrally close.
+The key distinction of this approach from others in the literature is applying the ACS SE preconditioner to the reduced-system of the projection-based framework with $\mathcal{O}(r^{d+1})$, rather than to the full-scale system (as was done for instance in the matrix-oriented GMRES framework proposed in [35]).
+These algorithmic innovations enable the method to achieve linear scaling with the one-dimensional resolution $N$, both for computational complexity [as $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{d+1})$] and for memory storage [as $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$], and allow us to perform 3D simulations of advection-diffusion problems with equivalent resolutions of $\mathcal{O}\left(10^{12}\right)$ degrees of freedom on a standard laptop, orders of magnitude faster than an optimally performing multigrid method scaling as $N^{d}\log(N)$ [21].
+Additionally, as in [14], the adaptive low-rank integrator is amenable for use with DIRK methods for high-order temporal accuracy (up to third-order is considered in this study).
+All these advertised properties will be demonstrated numerically in the test examples.
 The rest of this paper is organized as follows.
 In Section 2, we consider the case of rank-one advection and diffusion coefficients and derive the GSE that results from a finite-difference discretization in space and backward Euler in time.
-In Section 3, we introduce an adaptive-rank solution strategy for 2D problems, including the coefficient-averaged approximation operator, construction of xKrylov subspaces, preconditioning technique in solving the reduced system, and low-rank computation of the residual norm.
+In Section 3, we introduce an adaptive-rank solution strategy for 2D problems, including the coefficient-averaged approximation operator, the construction of extended Krylov subspaces (which we abbreviate as xKrylov in the sequel), the preconditioning technique in solving the reduced system, and the low-rank computation of the residual norm.
 The extension of the technique to high-order temporal discretizations with multi-rank coefficients is provided in Appendix C, and for 3D problems in Appendix D.
-In Section 4, numerical results are presented to demonstrate the efficacy of the proposed algorithms.
-Finally, we conclude and provide outlook for future research in Section 5.
+In Section 4, numerical results are presented to demonstrate the efficacy of the algorithm.
+Finally, we conclude and provide an outlook for future research in Section 5.
 2 Problem Description and Generalized Sylvester Equation
-For simplicity, we focus this section on 2D, with the extension to 3D elaborated in Appendix D.
+For simplicity, we focus the discussion in this section on 2D, with the extension to 3D elaborated in Appendix D.
 We establish the following notation conventions.
@@
 We assume that the diffusion and advection coefficients are separable with finite rank. Specifically, the diagonal diffusion coefficients, $\phi^{x}(x,y)$ and $\phi^{y}(x,y)$, are given by:
-$\phi^{x}(x,y)=\sum_{i=1}^{\ell_{x}}\phi^{1,x}_{i}(x)\phi^{2,x}_{i}(y),\quad\phi^{y}(x,y)=\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\phi^{2,y}_{i}(y).$ (5)
+$\phi^{x}(x,y)=\sum_{i=1}^{\ell_{x}}\phi^{1,x}_{i}(x)\phi^{2,x}_{i}(y),\qquad\phi^{y}(x,y)=\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\phi^{2,y}_{i}(y).$ (4)
 Similarly, for the advection coefficients we have:
-$\sigma^{x}(x,y)=\sum_{i=1}^{k_{x}}\sigma^{1,x}_{i}(x)\sigma^{2,x}_{i}(y),\quad\sigma^{y}(x,y)=\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\sigma^{2,y}_{i}(y).$ (6)
+$\sigma^{x}(x,y)=\sum_{i=1}^{k_{x}}\sigma^{1,x}_{i}(x)\sigma^{2,x}_{i}(y),\qquad\sigma^{y}(x,y)=\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\sigma^{2,y}_{i}(y).$ (5)
 Substituting these coefficients in their separable forms into Equation (3), we obtain:
@@
 To simplify the presentation, we assume for now rank-one advection and diffusion coefficients:
-$\phi^{x}=\phi^{1,x}(x)\phi^{2,x}(y),\quad\phi^{y}=\phi^{1,y}(x)\phi^{2,y}(y),$
-$\sigma^{x}=\sigma^{1,x}(x)\sigma^{2,x}(y),\quad\sigma^{y}=\sigma^{1,y}(x)\sigma^{2,y}(y).$
+$\phi^{x}=\phi^{1,x}(x)\phi^{2,x}(y),\qquad\phi^{y}=\phi^{1,y}(x)\phi^{2,y}(y),$
+$\sigma^{x}=\sigma^{1,x}(x)\sigma^{2,x}(y),\qquad\sigma^{y}=\sigma^{1,y}(x)\sigma^{2,y}(y).$
 This will be generalized to multiple ranks below. Equation (6) then simplifies to:
@@
 We consider the point-wise multiplication of the advection and diffusion coefficients in the discrete by setting:
-$\Phi_{1}$ $=\text{diag}(\phi^{1,y}(x_{i})),\Sigma_{1}=\text{diag}(\sigma^{1,y}(x_{i})),\quad i=0,\cdots,N_{1}-1,$
-$\Phi_{2}$ $=\text{diag}(\phi^{2,x}(y_{j})),\Sigma_{2}=\text{diag}(\sigma^{2,x}(y_{j})),\quad j=0,\cdots,N_{2}-1.$
+$\Phi_{1}$ $=\text{diag}(\phi^{1,y}(x_{i})),\Sigma_{1}=\text{diag}(\sigma^{1,y}(x_{i})),\qquad i=0,\cdots,N_{1}-1,$
+$\Phi_{2}$ $=\text{diag}(\phi^{2,x}(y_{j})),\Sigma_{2}=\text{diag}(\sigma^{2,x}(y_{j})),\qquad j=0,\cdots,N_{2}-1.$
 We approximate the spatial derivative with a second-order finite-difference scheme using a three-point stencil, and obtain the following matrix differential equation for $\mathbf{F}$, for which its $(i,j)$ entry $F_{i,j}$ corresponds to the approximation of $f(x_{i},y_{j})$:
@@
 For notational simplicity, let:
-$T_{1}=D_{xx}^{\Phi},\quad T_{2}=D_{yy}^{\Phi},\quad T_{3}=-D_{x}^{\Sigma},\quad T_{4}=-D_{y}^{\Sigma}.$ (10)
+$T_{1}=D_{xx}^{\Phi},\qquad T_{2}=D_{yy}^{\Phi},\qquad T_{3}=-D_{x}^{\Sigma},\qquad T_{4}=-D_{y}^{\Sigma}.$ (9)
 Equation (8) can be written compactly as:
@@
 3 Krylov-Based Adaptive-Rank Solver for Generalized Sylvester Equation
-We propose an adaptive-rank algorithm for the GSE (13).
+We describe next the projection-based framework for solving the GSE (12) [5, 20, 45, 36].
 We assume a low-rank factorization of the initial condition, $\mathbf{F}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$, which evolves to an updated solution $\mathbf{F}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ at time $t^{(1)}=t^{(0)}+\Delta t$ also in the low-rank format, i.e.:
-[…]
+$\mathbf{F}_{0}=U_{0}S_{0}V_{0}^{\top}\xrightarrow{\text{evolve in time}}\mathbf{F}_{1}=U_{1}S_{1}V_{1}^{\top},$ (14)
 where $U_{0}\in\mathbb{R}^{N_{1}\times r_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r_{1}}$, and $V_{1}\in\mathbb{R}^{N_{2}\times r_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$ and $S_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$ are diagonal matrices with singular values at the corresponding times.
 Projection methods involve constructing bases $U_{1}$ and $V_{1}$ and then projecting the original equation onto them to determine the coefficients of $S_{1}$.
-A key aspect of the proposed adaptive-rank approach is that we never form the full solutions explicitly.
-Instead, we seek bases in a dimension-by-dimension manner and then use them to perform a Galerkin projection.
-Key questions for solving the GSE that we address in this study include:
-  • How do we choose suitable subspaces effectively to build $U_{1}$ and $V_{1}$ from $U_{0}$ and $V_{0}$?
-  • How can we solve the reduced equation for $S_{1}$ efficiently?
-  • How can we check the residual of the GSE without forming the matrices explicitly?
-Our main innovation leverages an approximate operator $\tilde{\mathscr{L}}$ (which is of the Sylvester form) to the original operator ${\mathscr{L}}$ in (12).
-This approximation not only allows us to construct dimension-wise Krylov bases $U_{1}$ and $V_{1}$ effectively and efficiently, but also enables effective preconditioning for solving the projected reduced system for $S_{1}$, in the spirit of [31].
-In the following, we cover the construction of the approximated operator $\tilde{\mathscr{L}}$ in Section 3.1; we discuss the dimension-wise basis construction in Section 3.2; we present the reduced equation for the matrix of coefficients $S_{1}$ and introduce our ACS preconditioner in Section 3.3, and outline the residual computation in Section 3.4.
+Efficiently solving the GSE requires choosing suitable solution subspaces effectively to build $U_{1}$ and $V_{1}$ from $U_{0}$ and $V_{0}$, and solving the reduced equation for $S_{1}$ efficiently.
+To accomplish both, in the spirit of the work in [35] we leverage an approximate operator $\tilde{\mathscr{L}}$ (which is of the Sylvester form) to the original operator $\mathscr{L}$} in (11).
+In the following, we describe the construction of the approximated operator $\tilde{\mathscr{L}}$ in Section 3.1; we discuss the dimension-wise basis construction in Section 3.2; we present the reduced equation for the matrix of coefficients $S_{1}$ and introduce our particular implementation of the ACS preconditioner in Section 3.3, and outline the residual computation in Section 3.4.
 Finally, Section 3.5 extends the algorithm to the case of multi-rank coefficients.
 3.1 Averaged-Coefficient Approximate Operator $\tilde{\mathscr{L}}$
-Our approach incorporates ideas from [31] and [14] to construct an approximate operator $\tilde{\mathscr{L}}$ and efficiently solve the time-dependent inhomogeneous advection-diffusion equation using a projection-based adaptive-rank integration framework.
-The approximated operator $\tilde{\mathscr{L}}$ serves two purposes: (i) constructing the solution bases $U_{1}$ and $V_{1}$ as xKrylov subspaces, and (ii) preconditioning the solution of the coefficient matrix $S_{1}$.
+We incorporate ideas from [35] and [14] to construct an approximate operator $\tilde{\mathscr{L}}$ and efficiently solve the time-dependent inhomogeneous advection-diffusion equation using a projection-based adaptive-rank integration framework.
+The approximated operator $\tilde{\mathscr{L}}$ serves two purposes: (i) contribute to constructing the solution bases $U_{1}$ and $V_{1}$ from xKrylov subspaces, and (ii) preconditioning the solution of the coefficient matrix $S_{1}$.
 We seek an approximation to $\mathscr{L}$} that captures its spectral content and is easier to invert.
@@
 This approximation yields the following mapping:
-~${\tilde{\mathscr{L}}}:\mathbf{F}\mapsto\left(\alpha_{2}T_{1}+\gamma_{2}T_{3}\right)\mathbf{F}+\mathbf{F}\left(\alpha_{1}T_{2}+\gamma_{1}T_{4}\right)^{\top}.$ (16)
+${\tilde{\mathscr{L}}}:\mathbf{F}\mapsto\left(\alpha_{2}T_{1}+\gamma_{2}T_{3}\right)\mathbf{F}+\mathbf{F}\left(\alpha_{1}T_{2}+\gamma_{1}T_{4}\right)^{\top}.$ (15)
 A backward Euler temporal discretization using (15) leads to the SE:
@@
 with:
-$A_{1}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{2}T_{1}\right)\quad A_{3}=\left(\frac{1}{4}I-\Delta t\,\gamma_{2}T_{3}\right),\quad P_{1}=A_{1}+A_{3},$ (18)
-$A_{2}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{1}T_{2}\right)\quad A_{4}=\left(\frac{1}{4}I-\Delta t\,\gamma_{1}T_{4}\right),\quad P_{2}=A_{2}+A_{4}.$ (19)
-The solution to (17) is an approximation to that of (13), but it can be efficiently computed using low-rank approximations [14, 39, 37, 36].
-It can therefore be exploited as a viable preconditioner for a GMRES iterative solution of the GSE (13).
+$A_{1}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{2}T_{1}\right)\qquad A_{3}=\left(\frac{1}{4}I-\Delta t\,\gamma_{2}T_{3}\right),\qquad P_{1}=A_{1}+A_{3},$ (17)
+$A_{2}$ $=\left(\frac{1}{4}I-\Delta t\,\alpha_{1}T_{2}\right)\qquad A_{4}=\left(\frac{1}{4}I-\Delta t\,\gamma_{1}T_{4}\right),\qquad P_{2}=A_{2}+A_{4}.$ (18)
+Equation (16) is an SE approximation to the GSE (12), which can be efficiently solved using low-rank approximations [14, 42, 40, 39].
+It can be exploited as a preconditioner for a GMRES iterative solution of the GSE (12).
 This was done in [35] with $\mathcal{O}(N^{3})$ computational complexity.
 In this work, we extend this approach by performing a Galerkin projection of both the original operator $\mathscr{L}$ and its approximation $\tilde{\mathscr{L}}$ on the low-rank bases before solving the GSE with GMRES.
-This reduces the problem to a lower-dimensional subspace, where preconditioning can be applied more effectively on a smaller $d$-dimensional problem with $\mathcal{O}(r^{d+1})$ complexity, resulting in an overall computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$.
+This reduces the problem to a lower-dimensional subspace, where preconditioning can be applied more effectively on a smaller $d$-dimensional problem with $\mathcal{O}(r^{d+1})$ complexity.
 3.2 Basis Construction
-We begin by defining the xKrylov subspaces used for our basis construction at the $m$-th outer iteration as:
-$\kappa_{m}(M_{1},\cdots,M_{l},U)=\text{{\color[rgb]{0,0,0}Range}}\Big{\{}U,\underbrace{M_{1}U,\cdots,M_{l}U}_{\text{1st augmentation}},\ldots,\underbrace{M_{1}^{m+1}U,\cdots,M_{l}^{m+1}U}_{\text{mth augmentation}}\Big{\}}.$ (20)
-where $M_{1},\cdots,M_{l}$ are operators of interest and their inverses arising in our GSE, and $l$ is the number of operators used for each basis augmentation step.
-The xKrylov basis construction involves several operators from the original GSE (13) and its approximation (17).
-Following [14], we construct an xKrylov subspace in the $x$-direction for the approximated SE (17) using $P_{1}$ and its inverse.
-We also include the inverses of the individual operators $A_{1}$ and $A_{3}$, as they may contain information not captured by the inverse of $P_{1}$.
-Additionally, we consider the action of the diagonal operators $\Phi_{1}$, $\Sigma_{1}$, $\Phi_{2}$, and $\Sigma_{2}$, which are averaged to obtain $\tilde{\mathscr{L}}$.
-We have found that including all these matrices is effective in capturing the range of the solution.
-Using these operators, we construct a generalized Krylov subspace in the $x$-direction as follows:
-$\kappa_{m}\left(M_{1},\cdots,M_{6},U_{0}\right),\quad\text{where}$ $M_{1}=P_{1},$ $M_{2}=P_{1}^{-1},$ $M_{3}=A_{1}^{-1},$ $M_{4}=A_{3}^{-1},$ $M_{5}=\Phi_{1},$ $M_{6}=\Sigma_{1}.$
-Depending on the structure of the GSE (2), $l$ may be large, and the generated subspace vectors at each augmentation may exhibit linear dependencies.
+The construction of the the xKrylov subspaces per dimension at the $m$-th outer iteration is performed as:
+$\kappa_{m}(M_{1},\cdots,M_{l},U)=\text{{\color[rgb]{0,0,0}Range}}\Big\{U,\underbrace{M_{1}U,\cdots,M_{l}U}_{\text{1st augmentation}},\ldots,\underbrace{M_{1}^{m+1}U,\cdots,M_{l}^{m+1}U}_{\text{mth augmentation}}\Big\},$ (19)
+where $M_{1},\cdots,M_{l}$ are operators of interest and their inverses arising in the GSE, and $l$ is the number of operators used for each basis augmentation step.
+For the $M_{l}$ operators, we include $P_{1}$ and its inverse from the SE approximation to the GSE (16), the individual operators $A_{1}$ and $A_{3}$ and their inverses, as they may contain information not captured by the inverse of $P_{1}$, and the diagonal operators $\Phi_{1}$, $\Sigma_{1}$, $\Phi_{2}$, and $\Sigma_{2}$, which are averaged to obtain $\tilde{\mathscr{L}}$.
+Using these operators, the generalized Krylov subspace (for instance in the $x$-direction) considered is as follows:
+$\kappa_{m}\left(M_{1},\cdots,M_{6},U_{0}\right),\qquad\text{where}$ $M_{1}=P_{1},$ $M_{2}=P_{1}^{-1},$ $M_{3}=A_{1}^{-1},$ $M_{4}=A_{3}^{-1},$ $M_{5}=\Phi_{1},$ $M_{6}=\Sigma_{1}.$
+We have found that including all these matrices is important to capture the range of the solution.
+Depending on the structure of the GSE (2), $l$ in (19) may be large, and the generated subspace vectors at each augmentation may exhibit linear dependencies.
 To limit subspace inflation, we employ an SVD-based truncation strategy (Algorithm 1), such that only basis vectors corresponding to significant singular values are retained.
-The construction of the dimension-wise xKrylov subspaces (20), following [39], is as follows.
+With this in mind, the construction of the dimension-wise xKrylov subspaces (19), following [42], is as follows.
 At each iteration $m$, Algorithm 2 augments the previously constructed basis $\mathcal{V}_{(m-1)}$ by applying powers of operators $M_{1}$ through $M_{l}$ to their corresponding basis blocks $V^{(1)},\dots,V^{(l)}$.
@@
 The truncation step in line 2 of Algorithm 2 removes any linear dependencies within each block during the power iterations; line 2 applies a global truncation across all newly generated blocks, eliminating redundancies among them.
-As a result, the basis of the subspace is kept optimal in size.
+As a result, the subspace is kept optimal in size.
 Input: Krylov Subspace $\kappa_{m}$; Tolerance $\epsilon_{\kappa}$.
-Output: Truncated bases ${Q}$.
+Output: Truncated bases Q{Q}.
 1 Compute the reduced QR decomposition: $\{Q,R\}=\texttt{QR}\left(\kappa_{m}\right)$;
-2 Perform reduced SVD decomposition: $Z_{1}\tilde{R}Z_{2}=\texttt{SVD}(R)$, where $\tilde{R}=\text{diag}(\sigma_{j})$ Identify the last index $r$ such that $\sigma_{r}>\epsilon_{\kappa}$;
+2 Perform reduced SVD decomposition: $Z_{1}\tilde{R}Z_{2}=\texttt{SVD}(R)$, where $\tilde{R}=\text{diag}(\sigma_{j})$ ;
+Identify the last index $r$ such that $\sigma_{r}>\epsilon_{\kappa}$ ;
+// Optional: Set Maximum Allowable Rank
 3 Update the basis vectors: $Q\leftarrow QZ_{1}(:,1:r)$;
@@
 Algorithm 2 Construction of xKrylov Subspace $\kappa_{m}^{\texttt{trunc}}$
-We comment next on the performance of the xKrylov iteration.
+We comment next on the expected performance of the xKrylov iteration.
 In principle, the number of xKrylov iterations may depend on the mesh size, $N$.
-For instance, [24, 4] theoretically show for the 2D and 3D Poisson equations that the dimension of the extended Krylov subspace (denoted by $r^{(m)}$ in the basis construction above) scales as $\mathcal{O}(\kappa^{1/4})\sim\mathcal{O}(N^{1/2})$, with $\kappa\sim N^{2}$ the condition number.
-While this provides an upper bound of the convergence rate (without considering solution smoothness), our numerical results suggest that it may be pessimistic in practice.
-In particular, for moderately tight tolerances, we find that the required Krylov subspace dimensions remain independent of grid resolution for a wide range of grid sizes $N$.
+For instance, [25, 4] theoretically show for the 2D and 3D Poisson equations that the dimension of the extended Krylov subspace (denoted by $r^{(m)}$ in the basis construction above) scales as $\mathcal{O}(\kappa^{1/4})\sim\mathcal{O}(N^{1/2})$ in 2D, and $(\kappa^{2/3})\sim\mathcal{O}(N^{4/3})$ in 3D, with $\kappa\sim N^{2}$ the condition number.
+While this provides an upper bound of the convergence rate (without considering solution smoothness), our numerical results suggest that it may be pessimistic in practice, with Krylov subspace dimensions remaining largely independent of grid resolution for moderate convergence tolerances and mesh refinements.
 3.3 Solving for the Matrix of Coefficients, $S_{1}$: ACS-Preconditioned GMRES
@@
 The reduced GSE for ${\bf S}_{1}$ arises from projecting the residual via the Galerkin condition $U_{1}^{\top}\mathbf{R}_{\mathscr{L}}V_{1}=0$, yielding:
-${\bf S}_{1}-\Delta t\left(\tilde{T}_{1}{\bf S}_{1}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}{\bf S}_{1}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}_{1}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}{\bf S}_{1}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}_{1},\qquad\text{with}\qquad\tilde{{\bf B}}_{1}=(U_{1}^{\top}U_{0})S_{0}(V_{1}^{\top}V_{0})^{\top}.$ (21)
+${\bf S}_{1}-\Delta t\left(\tilde{T}_{1}{\bf S}_{1}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}{\bf S}_{1}\tilde{T}_{2}^{\top}+\tilde{T}_{3}{\bf S}_{1}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}{\bf S}_{1}\tilde{T}_{4}^{\top}\right)=\tilde{{\bf B}}_{1},\hskip 18.49988pt\text{with}\hskip 18.49988pt\tilde{{\bf B}}_{1}=(U_{1}^{\top}U_{0})S_{0}(V_{1}^{\top}V_{0})^{\top}.$ (20)
 This matrix-based equation is equivalent to the classical Kronecker formulation:
@@
 The projected operators are defined as follows:
-$\tilde{T}_{1}$ $=U_{1}^{\top}T_{1}U_{1},\quad\tilde{T}_{3}=U_{1}^{\top}T_{3}U_{1},\quad\tilde{\Phi}_{1}=U_{1}^{\top}\Phi_{1}U_{1},\quad\tilde{\Sigma}_{1}=U_{1}^{\top}\Sigma_{1}U_{1},$
-$\tilde{T}_{2}$ $=V_{1}^{\top}T_{2}V_{1},\quad\tilde{T}_{4}=V_{1}^{\top}T_{4}V_{1},\quad\tilde{\Phi}_{2}=V_{1}^{\top}\Phi_{2}V_{1},\quad\tilde{\Sigma}_{2}=V_{1}^{\top}\Sigma_{2}V_{1}.$
-Solving the Kronecker formulation $\mathcal{A}\text{vec}({{\bf S}_{1}})=\text{vec}({\tilde{{\bf B}}_{1}})$ in (22) with a direct solver scales as $\mathcal{O}(r^{6})$ in 2D and $\mathcal{O}(r^{9})$ in 3D, where $r=\text{max}(r_{x},r_{y})$, which may be too expensive even for moderate $r$-values.
+$\tilde{T}_{1}$ $=U_{1}^{\top}T_{1}U_{1},\qquad\tilde{T}_{3}=U_{1}^{\top}T_{3}U_{1},\qquad\tilde{\Phi}_{1}=U_{1}^{\top}\Phi_{1}U_{1},\qquad\tilde{\Sigma}_{1}=U_{1}^{\top}\Sigma_{1}U_{1},$
+$\tilde{T}_{2}$ $=V_{1}^{\top}T_{2}V_{1},\qquad\tilde{T}_{4}=V_{1}^{\top}T_{4}V_{1},\qquad\tilde{\Phi}_{2}=V_{1}^{\top}\Phi_{2}V_{1},\qquad\tilde{\Sigma}_{2}=V_{1}^{\top}\Sigma_{2}V_{1}.$
+Solving the Kronecker formulation $\mathcal{A}\text{vec}({{\bf S}_{1}})=\text{vec}({\tilde{{\bf B}}_{1}})$ in (21) with a direct solver scales as $\mathcal{O}(r^{3d})$, with $r$ the maximum rank of all dimensions and $d$ the dimensionality, which may be too expensive even for moderate $r$-values.
 Instead, we solve this equation iteratively using preconditioned GMRES.
-Instead of adapting GMRES to the matrix equation, we wrap GMRES around the Kronecker form, but exploiting the matrix-equation form (21) for both the $\mathcal{A}$-times-vector operation and the preconditioning step as indicated below, with overall complexity $\mathcal{O}(r^{3})$ in 2D (and $\mathcal{O}(r^{4})$ in 3D).
+We wrap standard GMRES around the Kronecker form of the GSE, but exploiting the matrix-equation form (20) for both the $\mathcal{A}$-times-vector operation and the preconditioning step as indicated below.
+The resulting overall complexity is $\mathcal{O}(r^{d+1})$.
 For the GMRES operator-vector product, given a vector $s_{k}=\text{vec}(\mathbf{S}_{k})$, the product of the matrix $\mathcal{A}$ times a vector $s_{k}$ is performed as follows:
@@
   2. 2.
-    Compute $\hat{\mathbf{Z}}_{k}\leftarrow\mathbf{S}_{k}-\Delta t\left(\tilde{T}_{1}\mathbf{S}_{k}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}\mathbf{S}_{k}\tilde{T}_{2}^{\top}+\tilde{T}_{3}\mathbf{S}_{k}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}\mathbf{S}_{k}\tilde{T}_{4}^{\top}\right)$, with complexity $\mathcal{O}(r^{3})$.
+    Compute $\hat{\mathbf{Z}}_{k}\leftarrow\mathbf{S}_{k}-\Delta t\left(\tilde{T}_{1}\mathbf{S}_{k}\tilde{\Phi}_{2}+\tilde{\Phi}_{1}\mathbf{S}_{k}\tilde{T}_{2}^{\top}+\tilde{T}_{3}\mathbf{S}_{k}\tilde{\Sigma}_{2}+\tilde{\Sigma}_{1}\mathbf{S}_{k}\tilde{T}_{4}^{\top}\right)$, with complexity $\mathcal{O}(r^{d+1})$.
   3. 3.
@@
 For efficiency, GMRES requires preconditioning so that the eigenvalues of the preconditioned matrix are sufficiently clustered to deliver $r$-independent convergence rates  [7].
-Here, we employ the ACS preconditioner, with $\mathcal{P}=I_{r_{y}}\otimes\tilde{P}_{1}+\tilde{P}_{2}\otimes I_{r_{x}}$ arising from the Galerkin projection of (17).
-The ACS preconditioner is expected to deliver optimal rank-independent, superlinear convergence (both observed in our numerical experiments), owing to its connection to the theory of compact-equivalent operators for diffusion-like equations, [^1] The averaged-coefficient advection-diffusion equation is compact-equivalent to the variable-coefficient one.
-which are known to be effective preconditioners for Krylov methods [1, 2].
-We consider left preconditioning as:
+Here, we employ the ACS preconditioner, with $\mathcal{P}=I_{r_{y}}\otimes\tilde{P}_{1}+\tilde{P}_{2}\otimes I_{r_{x}}$ arising from the Galerkin projection of (16), with $\tilde{P}_{1}=U_{1}^{\top}P_{1}U_{1}$ and $\tilde{P}_{2}=V_{1}^{\top}P_{2}V_{1}$.
+The ACS preconditioner delivers optimal, rank-independent, superlinear convergence (all observed in our numerical experiments), owing to its connection to the theory of compact-equivalent operators for diffusion-like equations.
+[^1] The averaged-coefficient advection-diffusion equation is compact-equivalent to the variable-coefficient one, which are known to be effective preconditioners for Krylov methods [1, 2].
+In particular, we consider left preconditioning as:
 $\mathcal{P}^{-1}\mathcal{A}\text{vec}({\bf S}_{1})=\mathcal{P}^{-1}\text{vec}(\tilde{{\bf B}}_{1}).$ (22)
@@
   2. 2.
-    Solve $\mathcal{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\mathcal{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k}$ with complexity $\mathcal{O}(r^{d+1})$, for example with the Bartels-Stewart algorithm [15].
+    Solve $\mathcal{P}_{1}\mathbf{Z}_{k}+\mathbf{Z}_{k}\mathcal{P}_{2}^{\top}=\hat{\mathbf{Z}}_{k}$ with complexity $\mathcal{O}(r^{d+1})$, e.g. with the Bartels-Stewart algorithm [15].
   3. 3.
@@
 As a result, the $S$-matrix solve scales as $O(r^{d+1})$ overall, a great reduction in complexity compared to the $O(r^{3d})$ scaling of a naïve implementation of GMRES with the Kronecker form.
-3.4 Truncation, Efficient Residual Computation
+3.4 Truncation and Residual Computation
 Once the solution ${\bf F}_{1}=U_{1}S_{1}V_{1}^{\top}$ is obtained, we apply truncation to control rank growth and reduce the residual computational load.
@@
 The SVD truncation gives ${\bf F}_{1}=U_{1,\epsilon}S_{1,\epsilon}V_{1,\epsilon}^{\top}$, with $U_{1,\epsilon}\in\mathbb{R}^{N_{1}\times r_{1}}$, $V_{1,\epsilon}\in\mathbb{R}^{N_{2}\times r_{1}}$, and $S_{1,\epsilon}\in\mathbb{R}^{r_{1}\times r_{1}}$, with $r_{1}\leq\min(r_{x},r_{y})$ (Algorithm 3).
-Following [14], we implement a cost-effective way to compute the residual using the identity:
+Following [14], we compute the residual using the identity:
 $\|\mathbf{R_{{\mathscr{L}}}}\|$ $=\|\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}\|$
 $=\underbrace{\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}\\ \end{bmatrix}R_{V}^{\top}\right\|}_{\in\mathbb{R}^{(r_{x}+5r_{1})\times(r_{y}+5r_{1})}}$
-$={\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{S_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,(I_{\mathcal{R}_{\text{ranks}}}\otimes{S_{1,\epsilon}})\end{bmatrix}R_{V}^{\top}\right\|}\leq\epsilon_{\texttt{tol}},$
+$={\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{S_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,(I_{\mathcal{R}_{\text{ranks}}}\otimes{S_{1,\epsilon}})\end{bmatrix}R_{V}^{\top}\right\|}\leq\epsilon_{\texttt{tol}},$ (23)
 where $\mathcal{R}_{\text{ranks}}={\ell_{x}+k_{x}+\ell_{y}+k_{y}}$ is the total rank number of the advection and diffusion coefficients, and $R_{U}$ and $R_{V}$ are the upper triangular matrices arising from the following Q-less QR decomposition:
-{_$\{\_\ ,R_{U}\}$ $=\texttt{QR}\left(\begin{bmatrix}{U}_{1}&{U}_{1,\epsilon}&{T_{1}}{U}_{1,\epsilon}&\Phi_{1}{U}_{1,\epsilon}&{T_{3}}{U}_{1,\epsilon}&\Sigma_{1}{U}_{1,\epsilon}\end{bmatrix}\right),$ (24)
-{_$\{\_\ ,R_{V}\}$ $=\texttt{QR}\left(\begin{bmatrix}{V}_{1}&{V}_{1,\epsilon}&\Phi_{2}{V}_{1,\epsilon}&{T_{2}}{V}_{1,\epsilon}&\Sigma_{2}{V}_{1,\epsilon}&{T_{4}}{V}_{1,\epsilon}\end{bmatrix}\right).$
+[…]
+[…]
 The above residual result follows from the property that $U_{1}(U_{1})^{\top}{\bf F}_{0}V_{1}(V_{1})^{\top}={\bf F}_{0}$, with ${\bf F}_{0}=U_{0}S_{0}V_{0}^{\top}$ (where $U_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ and $V_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ have orthonormal columns), since $U_{1}$ and $V_{1}$ contain $U_{0}$ and $V_{0}$, respectively.
+We note that, as discussed in [33] and analyzed in detail later in this study, forming, storing, and evaluating the norm of the residual tensor in (23) can be a very expensive operation when the number of GSE terms, $\mathcal{R}_{\text{ranks}}$, becomes large.
+This is so because the complexity of a naïve implementation of the residual evaluation scales as $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d+1})$ and the memory storage scales as $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d})$ (although the latter has been reduced in our implementation to $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{2})$ by careful algorithmic choreography).
+While this is not a pressing issue in our numerical experiments (which feature small-to-moderate values of $\mathcal{R}_{\text{ranks}}$), it could become a limiting factor in applications where $\mathcal{R}_{\text{ranks}}$ becomes large.
+In such instances, it may be necessary to use alternate strategies as proxy measures of convergence such as the difference between consecutive iterates [33], which is commensurate in cost with the rest of our algorithm.
 Input: Bases $U$, $V$; matrix of coefficients $\bf S$}; Tolerance $\epsilon$.
@@
 $\mathbf{F}_{1}-\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}_{1}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}_{1}T_{2,j}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}_{1}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}_{1}T_{4,l}^{\top}}_{\mathscr{L}(\mathbf{F}_{1})}\right)={\bf F}_{0}.$ (25)
-[…]
+As before, we find $\mathscr$$\tilde${L}} by averaging the diagonal terms $(\Phi_{i}^{2,x},\Phi_{j}^{1,y},\Sigma_{k}^{2,x},\Sigma_{l}^{1,y})$ from (25) to arrive at the approximated SE:
 ${P_{1}}\mathbf{F}_{1}+\mathbf{F}_{1}{P_{2}^{\top}}=\mathbf{F}_{0},$ (26)
@@
 This adjustment ensures that errors from the inner solve do not pollute the accuracy of the outer solve.
-The SVD truncation threshold $\epsilon$ is chosen to be one to three orders of magnitude smaller than the tolerance $\epsilon_{\text{tol}}$.
+The SVD truncation threshold $\epsilon$ is chosen to be one to three orders of magnitude smaller than the tolerance $\epsilon_{\texttt{tol}}$.
 Since solution truncation occurs before residual evaluations, the truncation error in the solution should remain below the residual tolerance to prevent residual pollution by over-truncation.
@@
 Truncated singular values $S_{1}$
+1ex
 1Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eqs. (27), (28);
@@
 3 Set $U^{(i)}=U_{0}$ and $V^{(i)}=V_{0}$ for $i=1,\dots,l$;
-4 Set $U_{1}=U_{0}\quad V_{1}=V_{0}$;
+4 Set $U_{1}=U_{0}\qquad V_{1}=V_{0}$;
 5 while not converged do
@@
 16   // Step K4.
-17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
-18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
+17    Compute $\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
+18    $\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
 19
@@
 Algorithm 4 Backward Euler Adaptive-Rank Integrator for the GSE.
-3.6 Computational Complexity
-We analyze next the computational complexity of Algorithm 4 for backward Euler.
+3.6 Analysis of Computational and Memory Complexity
+We analyze next the computational complexity and memory storage of Algorithm 4 for backward Euler.
 For conciseness, we consider the same spatial resolution per dimension, i.e., $N=N_{1}=N_{2}$.
 Starting from an initial condition $U_{0}S_{0}V_{0}^{\top}$, where $U_{0}\in\mathbb{R}^{N\times r_{0}}$, $V_{0}\in\mathbb{R}^{N\times r_{0}}$, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$, the $m$-th xKrylov iteration generates $U_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, $V_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, and $S_{1}\in\mathbb{R}^{r_{\kappa}\times r_{\kappa}}$, where $r_{\kappa}=(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0}+r^{(m-1)}$.
-Here, $r^{(m-1)}$ denotes the basis size from the previous xKrylov iteration (initialized with $r^{(0)}=r_{0}$), and $r^{\ast}_{0}\leq r_{0}$ is the possibly truncated initial rank as produced by line (2) of Algorithm 2.
+Here, $r^{(m-1)}$ denotes the basis size from the previous xKrylov iteration (with $r^{(0)}=r_{0}$), and $r_{0}^{\ast}\leq r_{0}$ is the (possibly truncated) initial rank produced at line (2) of Algorithm 2.
+While not used in this work, one may impose a maximum allowable rank $r_{\max}$ in the orthogonalization step of Algorithm 1 to prevent the subspace from exceeding available computational resources.
 Using SVD-truncated QR, this dimension is then reduced to smaller ranks, $r_{x}$ and $r_{y}$.
-For simplicity, let $\tilde{r}=\max(r_{x},r_{y})$.
+Let $\tilde{r}=\max(r_{x},r_{y})$.
 The matrix $S_{1}$ from the $m$-th inner iteration solve, initially of size $r_{x}\times r_{y}$, is truncated to rank $r_{1}\leq\min(r_{x},r_{y})$ using Algorithm 3.
@@
 Additionally, applying $P_{1}U_{0}$, $\Phi_{1,3}U_{0}$, $\Sigma_{1,3}U_{0}$, $P_{2}V_{0}$, $\Phi_{2,4}V_{0}$, and $\Sigma_{2,4}V_{0}$, can be done with $\mathcal{O}(Nr^{\ast}_{0})$ computational complexity for sparse diagonal and tridiagonal matrices, leading to a total complexity of $\mathcal{O}(N(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0})$ to augment the xKrylov subspaces.
-The use of Algorithm 2 to compute the orthonormal truncated bases $U_{1}$ and $V_{1}$ incurs a complexity of $\mathcal{O}\left(N\,\max\left[\left(2+\mathcal{R}_{\text{ranks}}\right)\,r_{0}^{\ast}\,r^{(m-1)},\;\left(\left(2+\mathcal{R}_{\text{ranks}}\right)\,r_{0}^{\ast}\right)^{2}\right]\right)\sim\mathcal{O}(Nr^{2})$.
+The use of Algorithm 2 to compute the orthonormal truncated bases $U_{1}$ and $V_{1}$ incurs a complexity of $\mathcal{O}\left(N\,\max\left(r_{\kappa}\,r^{(m-1)},\;r_{\kappa}^{2}\right)\right)\sim\mathcal{O}(Nr^{2})$.
+The memory complexity to store the truncated bases is $\mathcal{O}\left(N\,\left(r^{(m-1)}+r_{\kappa}\right)\right)\sim\mathcal{O}(Nr)$.
 Generally, we expect $r$ to be much smaller and independent of $N$ when the solution is sufficiently smooth and exhibits a low-rank structure.
@@
 The application of the reduced GSE operator using matrix-matrix multiplication and summation in the Krylov method (GMRES) features a complexity of $\mathcal{O}\left(\tilde{r}^{d+1}\right)$.
-Applying the ACS preconditioner requires solving an SE of size $r_{x}\times r_{y}$, which can be achieved using the Bartels-Stewart algorithm [15] with a complexity of $\mathcal{O}(\tilde{r}^{d+1})$.
-Hence, the $S$-step has $\mathcal{O}(r^{d+1})$ complexity.
+Applying the ACS preconditioner requires solving an SE of size $r_{x}\times r_{y}$, which can be achieved using the Bartels-Stewart algorithm [15] with a complexity of $\mathcal{O}(\tilde{r}^{d+1})\sim\mathcal{O}(r^{d+1})$.
+Hence, the computation of the $S$-step has $\mathcal{O}(r^{d+1})$   complexity.
+The memory complexity to store the S matrix/tensor is $\mathcal{O}\left(\tilde{r}^{d}\right)\sim\mathcal{O}(r^{d})$.
 Step K3
@@
 Step K4
-The QR decomposition to obtain $R_{U}$ and $R_{V}$ in lines (11)-(4) of Algorithm 4 features a complexity of $\mathcal{O}\left(N\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{2}\right)$.
-The residual computation in line (4) is of $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d+1}\right)$ due to the matrix multiplications, with an additional $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d}\right)$ for the Frobenius norm.
-It follows that, as advertised, the overall complexity of the algorithm scales as $Nr^{2}$ and as $r^{d+1}$, with $d$ the dimensionality, which will be numerically verified in Section 4.
+The QR decomposition to obtain $R_{U}$ and $R_{V}$ in lines (11)–(4) of Algorithm 4 has a complexity of $\mathcal{O}\left(N\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{2}\right)\sim\mathcal{O}(Nr^{2})$.
+The residual computation in line (4) requires $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d+1}\right)\sim\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d+1}r^{d+1})$ due to matrix multiplications, with an additional $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d}\right)\sim\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d}r^{d})$ for the Frobenius norm.
+The memory storage requirements for the residual tensor may be quite prohibitive if stored fully [it would scale as $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d}r^{d})$].
+Instead, we store only the $R_{U}$ and $R_{V}$ factor matrices and the $S$ matrix, and build the residual for the norm calculation on the fly still with computational complexity $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d+1}r^{d+1})$.
+Storing $R_{U}$ and $R_{V}$ requires $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{2}\right)\sim\mathcal{O}(\mathcal{R}_{\text{ranks}}^{2}r^{2})$, and the residual tensor blocks in Eq.
+23 require $\mathcal{O}(r^{d})$, resulting in a total storage of $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{2}r^{2})+\mathcal{O}(r^{d})$.
+Nevertheless, it is clear from this analysis that the residual evaluation may become prohibitively expensive when $\mathcal{R}_{\text{ranks}}\gg 1$.
+In such situations, the difference between successive solution iterates may be used as a alternate measure of convergence [33], commensurate in cost to the rest of the algorithm.
+It follows that the overall complexity of the algorithm scales as $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{d+1})$, and the memory complexity scales to leading order as $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$, with $d$ the dimensionality.
+Both claims will be numerically verified in Section 4.
 4 Numerical Experiments
-We demonstrate the overall efficacy of the proposed algorithm with several 2D and 3D numerical examples.
-We assume homogeneous Dirichlet boundary conditions for simplicity; other boundary conditions (.e.g, periodic, Neumann, etc.) can be easily incorporated in the formulation through the operators $T_{1:4}$ (10).
+We demonstrate the overall efficacy of the algorithm with several 2D and 3D numerical examples.
+We assume homogeneous Dirichlet boundary conditions for simplicity; other boundary conditions (e.g., periodic, Neumann, etc.) can be easily incorporated in the formulation through the operators $T_{1:4}$ (9).
 All simulations are performed using MATLAB.
 For efficient computational handling in 3D, we employ tools from the Tensor Toolbox [3] and the Htucker package [27].
+High-order temporal integration using DIRK is used throughout the numerical section.
+Details on the the DIRK scheme implementation in our adaptive-rank algorithm are provided in Appendix C.
 Throughout this section, we denote the diffusion Courant number and the advection Courant-Friedrichs-Lewy (CFL) number by $\lambda_{D}$ and $\lambda_{A}$, respectively, defined as:
-$\lambda_{D}=\frac{\Delta t\phi_{\text{max}}}{\Delta x^{2}},\qquad\lambda_{A}=\frac{\Delta t\sigma_{\text{max}}}{\Delta x},$ (34)
+$\lambda_{D}=\frac{\Delta t\phi_{\text{max}}}{\Delta x^{2}},\hskip 18.49988pt\lambda_{A}=\frac{\Delta t\sigma_{\text{max}}}{\Delta x},$ (34)
 where $\phi_{\text{max}}$ and $\sigma_{\text{max}}$ are the maximum of the diffusion and advection coefficients, respectively.
-These numbers are measures of the condition number of the associated GSEs.
+The numbers $\lambda_{D}$ and $\lambda_{A}$ are appropriate measures of the maximum eigenvalue of the associated GSE, and by extension of its condition number since the minimum eigenvalue is of $\mathcal{O}(1)$.
 For all simulations presented in this section, we did not need to set a maximum number of outer xKrylov iterations because the algorithm consistently converged within a small number of iterations.
@@
 $\sigma^{z}(x,y,z)$ $=\sigma^{1}_{3}(x)\sigma^{2}_{3}(y)\sigma^{3}_{3}(z),$ $\sigma^{1}_{2}(x)$ $=4x,$ $\sigma^{2}_{2}(y)$ $=2y,$ $\sigma^{3}_{3}(z)$ $=1-z^{2}.$
-and a rank-3 variable diffusion coefficient:
+Note that this flow is divergence-free and therefore a pure vorticity flow, which results in interesting flow dynamics such as differential rotation.
+For diffusion, we specify a rank-3 variable coefficient as follows:
 $\phi^{x}(x,y,z)=\phi^{y}(x,y,z)=\sum_{i=1}^{3}\phi^{1}_{i}(x)\phi^{2}_{i}(y)\phi^{3}_{i}(z),$
@@
 ${\phi}^{1}_{3}(x)$ $=\exp\left(-\left(x-0.6\sin(2\pi x)\right)^{2}\right),$ ${\phi}^{2}_{3}(y)$ $=\exp\left(-\left(y-0.6\sin(2\pi y)\right)^{2}\right)$ ${\phi}^{3}_{3}(z)$ $=\exp\left(-\left(y-0.6\sin(2\pi z)\right)^{2}\right).$
-The initial condition is a rank-2 Gaussian:
+These diffusion coefficients are relatively balanced and therefore mostly isotropic. We will investigate anisotropic diffusion in future work. The initial condition is a rank-2 Gaussian:
 $f(x,y,z,t=0)$ $=0.5\exp\left(-400\left((x-0.3)^{2}+(y-0.35)^{2}+(z-0.2)^{2}\right)\right)$
@@
 We use a spatial grid with $N=300$ in all dimensions, and the time-step $\Delta t$ is set to $\lambda_{D}\Delta x^{2}$, with $\lambda_{D}$ ranging from $140$ to $1125$.
-We test our adaptive-rank algorithm using three different integrators: backward Euler (BE), DIRK2, and DIRK3, as detailed in Tables 3 and 3 in the Appendix.
+We test the adaptive-rank algorithm using three different integrators: backward Euler (BE), DIRK2, and DIRK3, as detailed in Tables 4 and 4 in Appendix C.
 The residual tolerance is $\epsilon_{\texttt{tol}}=10^{-4}$ for BE, DIRK2, and DIRK3, the basis truncation threshold is $\epsilon_{\kappa}=10^{-6}$, the SVD truncation threshold is $\epsilon=10^{-6}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-6}$.
-We investigate numerically the following: (1) high-order temporal convergence, (2) effectiveness of the xKrylov iterations in capturing the range of solution, (3) performance of the proposed ACS preconditioner, and (4) verification of the overall computational complexity.
+We investigate numerically the following: (1) high-order temporal convergence, (2) effectiveness of the xKrylov iterations in capturing the range of solution, (3) performance of the ACS preconditioner to solve the reduced GSE, and (4) verification of the overall computational complexity.
 High-Order Temporal Convergence: Figure 1(a) illustrates the $L_{1}$ error norm vs time stepping size (characterized by $\lambda_{D}$) by showcasing the performance of the adaptive-rank integrators for BE, DIRK2, and DIRK3.
-The expected asymptotic temporal error order of convergence for our proposed adaptive-rank integrators, depicted by lines with markers, is observed.
+The expected asymptotic order of convergence of the temporal error for the adaptive-rank integrators, depicted by lines with markers, is observed.
 Additionally, the algorithm adeptly captures the rank evolution of the solution in the $x$, $y$, and $z$ directions, as illustrated in Figure 1(b).
@@
 We perform a single time-step integration with a fixed step size $\Delta t=10^{-3}$ with grid refinements in $N$.
-To isolate the effect of $\lambda_{D}$ (a measure of the GSE condition number) on the convergence behavior of xKrylov, we mitigate other potential influencing factors, such as excessive truncation of the Krylov subspace, insufficient convergence of the GMRES-ACS solver, and truncation errors in the constructed solution, by setting related thresholds to $\epsilon=\epsilon_{\texttt{GMRES}}=\epsilon_{\kappa}=10^{-10}$.
-Figure 2 illustrates the xKrylov subspace dimension required for convergence as a function of the diffusion Courant number $\lambda_{D}$ for a fixed $\Delta t$ and varying spatial resolutions $N$.
+The convergence bound for the extended Krylov method depends on the condition number of the advection-diffusion operator, $\kappa$.
+Assuming that the diffusion operator dominates the condition number, and that the variable coefficients perturb its eigenvalues only by factors of order unity, $\lambda_{D}$ provides a good estimate of the condition number (i.e., $\kappa=\mathcal{O}(\lambda_{D})$ in the case of the diffusion equation), and is therefore a good figure of merit to assess the convergence of the xKrylov subspaces.
+As reported in [25], the convergence factor for the extended Krylov subspace method applied to an equation with Kronecker product structure is $\kappa^{1/4}$ in 2D and $\kappa^{2/3}$ in 3D.
+To isolate the effect of $\lambda_{D}$ on the convergence of xKrylov, we mitigate other potential influencing factors, such as excessive truncation of the Krylov subspace, insufficient convergence of the GMRES-ACS solver, and truncation errors in the constructed solution, by setting related thresholds to $\epsilon=\epsilon_{\texttt{GMRES}}=\epsilon_{\kappa}=10^{-10}$.
+Figure 2 illustrates the xKrylov subspace dimension required for convergence as a function of the diffusion Courant number $\lambda_{D}$ for a fixed $\Delta t$ and varying the spatial 1D resolution, $N$.
 Our results suggest that, for a fixed tolerance, the convergence of the Krylov subspaces remains largely insensitive to grid refinement across a broad range of $\lambda_{D}$.
-However, for extremely fine mesh resolutions, beyond typical practical computational needs, the convergence rate of xKrylov scales as $\lambda_{D}^{1/4}$, consistently with the $\kappa^{1/4}$ scaling (with $\kappa$ the condition number) previously reported for the Poisson equation [24, 4].
+For extremely fine 1D mesh resolutions, $N\sim 10^{4}$, beyond typical practical computational needs, the convergence rate of xKrylov scales as $\lambda_{D}^{2/3}$, consistently with the theoretically expected $\kappa^{2/3}$ scaling in 3D [25, 4] .
 Figure 2: Example 4.1. Convergence study using the xKrylov subspace method with DIRK3 adaptive-rank integrator for several outer basis tolerances, $\epsilon_{tol}$.
-Performance of ACS-Preconditioner: We compare the performance of the GMRES solver with and without the ACS preconditioner for the inner solve.
+Performance of ACS-Preconditioner: Ill-conditioning for the projected system stems from both grid resolution (the projected system becomes more ill-conditioned as we refine the mesh) and rank growth (the projected system becomes larger in size).
+For effectiveness, the preconditioner must be able to deal with both sources of ill-conditioning.
+We consider first the scaling of the ACS-preconditioned GMRES inner solve with respect to the 1D resolution $N$.
+For this, we compare performance with and without the ACS preconditioner at the first time step ($\Delta t=0.01$) and the third outer-basis augmentation (i.e., keeping the subspace size constant, and therefore the rank), varying the mesh size.
 Figure 3 illustrates the GMRES residual history over iterations for ACS-preconditioned and unpreconditioned cases.
-Results at the first time step ($\Delta t=0.01$) and the third outer-basis augmentation are shown, comparing different mesh sizes.
 The results confirm that ACS-preconditioned GMRES delivers mesh-independent and super-linear convergence rates, which is not the case for unpreconditioned GMRES.
@@
 Simulations are conducted for a single time step with $\Delta t=0.01$ at the third outer-basis augmentation, using grid resolutions $N_{x}=N_{y}=N_{z}=200$ in (a), $1000$ in (b), and $2500$ in (c).
-Figure 4 (a) illustrates the average condition number of the Galerkin-projected diffusion and advection operators as the dimension of the proposed xKrylov subspaces increases across 12 outer iterations.
-For these results, the grid size is fixed at $N=1000$ in 3D, and a single time step of $\Delta t=0.01$ is used.
+We consider next the performance of the GMRES inner solve with respect to the solution rank $r$.
+For this test, the grid size is fixed at $N=1000$ in 3D, and a single time step $\Delta t=0.01$ is used.
+Figure 4 (a) illustrates the average condition number of the Galerkin-projected diffusion and advection operators as the dimension of the xKrylov subspaces increases across 12 outer iterations, confirming that the projected system gets more ill-conditioned as the rank increases.
 Figures 4 (b) - (d) show the GMRES residual histories for the ACS-preconditioned and unpreconditioned cases, corresponding to the second, fourth, and twelfth outer xKrylov iterations.
 These correspond to xKrylov subspace dimensions of $r=37$, $r=73$, and $r=217$, respectively.
-As the dimension of the xKrylov subspace grows, unpreconditioned GMRES convergence for the projected problem becomes progressively slower, as expected.
-In contrast, the ACS-preconditioner effectively maintains rapid convergence rates independent of subspace size $r$.
+As $r$ grows, unpreconditioned GMRES converges progressively slower, as expected.
+In contrast, the ACS-preconditioner maintains rapid convergence rates independent of subspace size $r$.
+To assess the relative cost of the preconditioner, we measure the ratio of the wall-clock timings for a single GMRES iteration with and without preconditioning for the different solution ranks considered in Fig.
+4, yielding $1.64$, $2.09$, and $1.85$, respectively.
+A preconditioned GMRES iteration is only about twice as expensive as an unpreconditioned one, but results in a dramatic improvement on GMRES’ convergence rate.
 Figure 4: Example 4.1.
@@
 Simulations are conducted for a single time step with $\Delta t=0.01$ and a grid of $N_{x}=N_{y}=N_{z}=1000$.
-Verification of the Computational Complexity scaling: We verify the $\mathcal{O}(N)$ computational complexity claim next.
+Verification of the Computational Complexity and Memory Scalings: We verify the $\mathcal{O}(N)$ computational complexity and memory storage claims next.
 We vary $N$ from $100$ to $30000$ and the time step is fixed at $\Delta t=10^{-3}$.
-Such range of mesh sizes is chosen to prevent the $O(r^{4})$ contributions of the ACS-GMRES solver from polluting the mesh-resolution scaling.
-Figure 5(a) displays $N$ vs.
-computational time for $10$ time steps of integration.
-We observe that the complexity remains linear with respect to $N$.
-The finite y-intercept corresponds to the $\mathcal{O}(r^{4})$ cost of the GMRES-ACS solve in 3D.
+Figure 5(a) displays a profiling study of the algorithm, depicting the computational time of various components vs.
+$N$ obtained for $10$ time steps.
+We observe that only the xKrylov component scales with mesh refinement $N$, and does so linearly.
+Other components do not scale with $N$, consistently with the rank of the solution being bounded.
+Furthermore, the inner solve remains the dominant computational cost of the algorithm until, for grids of size $N\approx 20{,}000$, the xKrylov routine begins to dominate.
+This, in turn, highlights the importance of an efficient reduced-system solution strategy.
 Figure 5(b) displays the computational time required for the preconditioned GMRES-ACS solver vs.
-rank for a single time step, using BE with $N=10,000$ and $\epsilon_{\texttt{GMRES}}=10^{-10}$.
-For this test, we allowed the basis to grow without truncation, recording the simulation time required for the GMRES solve at each basis augmentation iteration.
+rank for a single time step, using BE with $N=10^{4}$ and $\epsilon_{\texttt{GMRES}}=10^{-10}$.
+For this test, the xKrylov basis was allowed to continue to grow boundlessly, albeit with truncation applied at each augmentation step.
+The simulation time required for the GMRES solve was recorded at each basis augmentation.
 The expected $\mathcal{O}(r^{4})$ scaling in 3D is found.
-These studies verify that the complexity scaling of our approach is $\mathcal{O}(Nr^{2}+r^{4})$ in 3D.
+These studies verify that the complexity scaling of the approach is $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{4})$ in 3D.
+Figure 5(c) depicts the results of a memory scaling study demonstrating that the memory requirements also scale as $\mathcal{O}(N)$.
+The memory usage reported corresponds to the total size of all variables allocated within the solver function, as measured by MATLAB’s whos command at the end of the function call.
+A similar trend is expected (and observed in our simulations) in terms of memory storage: the inner tensor dominates the storage for coarser grids, while storing the xKrylov subspace dominates for finer grids.
 Figure 5: Example 4.1.
-(a) Demonstration of ${\cal O}(N)$ computational complexity of DIRK3 adaptive-rank integrator in 3D.
+(a) Profiling study depicting the computational cost of various components of the algorithm, demonstrating overall ${\cal O}(N)$ computational complexity of DIRK3 adaptive-rank integrator in 3D.
 (b) Demonstration of ${\cal O}(r^{4})$ computational complexity of the GMRES-ACS solver in 3D.
+For this test only, we allow the rank to grow indefinitely by setting the outer residual tolerance, $\epsilon_{\texttt{tol}}$, to machine precision.
+(c) Demonstration of ${\cal O}(N)$ memory storage complexity of DIRK3 adaptive-rank integrator in 3D.
 The execution times were recorded using MATLAB’s timeit function.
+The memory usage was recorded using MATLAB’s whos command at the end of each time step, and the values reported correspond to the average memory usage across all time steps in megabytes.
+To provide additional perspective on our results, we compare the performance of the adaptive-rank integrator for this example with a hypothetical optimal full-rank multigrid solver, which at best would scale as $\mathcal{O}(N^{d}\log N)$ [21].
+Assuming that the low-rank and multigrid solvers are at parity for $N=100$ (Appendix E), Table 1 reports the theoretical wall-clock-time speedup of the adaptive-rank integrator vs.
+the full-rank multigrid solver.
+The table suggests vast efficiency gains from our approach vs.
+optimal multigrid for this problem, scaling as 1D with grid refinement and effectively mitigating the curse of dimensionality.
+Table 1: Theoretical speedup of the adaptive-rank solver versus a hypothetical optimal multigrid solver for various 1D grid resolutions NN in a 3D domain.
+1D grid size NN Wall-clock-time speedup in 3D
+$1000$ $1446.1$
+$4000$ $99229.3$
+$10000$ $1418193.2$
 Example 4.2: 3D Swirling Deformation Flow
 We consider next a 3D swirling deformation flow:
-$f_{t}+\left[(1-x^{2})\cdot 2y\cdot-2z\cdot f\right]_{x}+\left[2x\cdot(1-y^{2})\cdot 2z\cdot f\right]_{y}+\left[-2x\cdot 2y\cdot 2(1-z^{2})\cdot f\right]_{z}=\nu\Delta f,\quad x,y,z\in[-1,1].$ (35)
-We set the diffusion coefficient $\nu=1/500$. The initial condition is a bi-Gaussian function given by:
+$f_{t}+\left[-4(1-x^{2})yzf\right]_{x}+\left[-4x(1-y^{2})zf\right]_{y}+\left[8xy(1-z^{2})f\right]_{z}=\nu\Delta f,\qquad x,y,z\in[-1,1].$ (35)
+We set the diffusion coefficient $\nu=1/500$. The flow is solenoidal, and therefore a pure-vorticity flow. The initial condition is a bi-Gaussian function given by:
 $f(x,y,0)$ $=0.5\exp\left(-\frac{1}{2\sigma^{2}}\left[(x-0.3)^{2}+(y-0.35)^{2}+(z-0.3)^{2}\right]\right)$
@@
 Figure 6 shows rank evolution.
-The solution rank rises sharply in the $x$ and $y$ directions [Figures 6 (a)-(b)].
-The solution rank in the $z$ direction evolves more gently [Figure 6 (c)] and is roughly one-third of that in the $x$ and $y$ directions.
+The solution rank rises sharply in the $x$- and $y$-directions [Figures 6 (a)-(b)].
+The solution rank in the $z$-direction evolves more gently [Figure 6 (c)] and is roughly one-third of that in the $x$ and $y$ directions.
 Figures 6 (d)-(e) showcase how the solution deforms according to the prescribed advection coefficients and diffuses slowly over time.
-Figure 7 shows snapshots of the solution in the $x\text{--}z$ plane at $y=-0.5$.
+Figure 7 shows snapshots of the solution in the $x\text{--}z$ plane at y=0.5y=-0.5.
 As the flow-driven deformation increases the complexity of the solution structure, the rank grows to capture these features; when the diffusion process starts to dominate, the rank decreases as the solution smooths.
-This example shows that the low-rank assumption is not intrinsic to our proposed algorithm.
+This example shows that the low-rank assumption is not intrinsic to the algorithm.
 Even when the rank becomes moderately large, the algorithm remains capable of capturing the underlying dynamics effectively.
 Naturally, the benefit of a low-rank approximation decreases as the solution approaches a full-rank structure.
+From a computational perspective, the per-time-step cost breakdown is as follows: xKrylov construction and orthogonalization (10%), ACS-GMRES inner solve (47.5%), and residual evaluation (42.5%).
+The relatively small contribution of the cost of xKrylov construction is partly a consequence of the truncation mechanism, together with the effectiveness of the chosen projection spaces, which collectively keep the approximation spaces compact.
+On average, in the $x$-$y$ directions, the truncation reduces the subspace dimension by a factor of about $4.5$, from roughly $450$ to $100$.
+In the $z$ direction, the reduction amounts to a factor of approximately $3$, from about $150$ to $50$.
+On the other hand, the relatively large contribution of the residual evaluation to the simulation time originates in the poor scaling of the computational complexity with the number of GSE terms (eight in this example, including the RHS) and suggests that a more economical stopping criterion (as discussed in Section 3.4) should be used.
 Figure 6: Example 4.2.
@@
 (d)–(e) Solution profiles at $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$.
+The red dots in (a)–(c) mark these time instants and correspond to the profiles shown in (d)–(e).
 The final simulation time is $T_{f}=10$.
@@
 3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid.
-(a)–(f) snapshots of the solution in the $x\text{--}z$ plane at $y=-0.5$, corresponding to times $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$, respectively.
+(a)–(f) snapshots of the solution in the $x\text{--}z$ plane at y=0.5y=-0.5, corresponding to times $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$, respectively.
 Example 4.3: A 2D Steady-State Balanced-Advection-Diffusion Problem
 The last example prescribes diffusion and advection coefficients that exactly balance each other, resulting in a non-trivial steady-state solution.
+This test demonstrates the ability of the algorithm to take large timesteps stably to solve directly for the steady state with expected spatial asymptotic accuracy.
 We seek an equilibrium solution satisfying the steady-state condition:
@@
 which satisfies homogeneous Dirichlet boundary conditions. The diffusion coefficients are defined as:
-$\phi^{x}(x,y)=\phi^{y}(x,y)$ $=\phi^{1}(x)\phi^{2}(y),\qquad\phi^{1}(x)=x^{2}(1-x)^{2},\qquad\phi^{2}(y)=y^{2}(1-y)^{2}.$
+$\phi^{x}(x,y)=\phi^{y}(x,y)$ $=\phi^{1}(x)\phi^{2}(y),\hskip 18.49988pt\phi^{1}(x)=x^{2}(1-x)^{2},\hskip 18.49988pt\phi^{2}(y)=y^{2}(1-y)^{2}.$
 The variable advection coefficients are $\sigma^{x}(x,y)=\sigma_{1}^{1}(x)\sigma_{1}^{2}(y)$, $\sigma^{y}(x,y)=\sigma_{2}^{1}(x)\sigma_{2}^{2}(y),$ where:
-$\sigma_{1}^{1}(x)$ $=2x(1-3x+2x^{2}),\qquad\sigma_{1}^{2}(y)=y^{2}(1-y)^{2},\qquad\sigma_{2}^{1}(x)=x^{2}(1-x)^{2},\qquad\sigma_{2}^{2}(y)=2y(1-3y+2y^{2}).$
+$\sigma_{1}^{1}(x)$ $=2x(1-3x+2x^{2}),\hskip 18.49988pt\sigma_{1}^{2}(y)=y^{2}(1-y)^{2},\hskip 18.49988pt\sigma_{2}^{1}(x)=x^{2}(1-x)^{2},\hskip 18.49988pt\sigma_{2}^{2}(y)=2y(1-3y+2y^{2}).$
 We select the following rank-1 initial condition: $f(x,y,0)=\left|\sin(2\pi x)\sin(2\pi y)\right|$.
 The outer basis residual tolerance is $\epsilon_{\texttt{tol}}=10^{-3}$, the basis truncation threshold is $\epsilon_{\kappa}=10^{-8}$, the SVD truncation threshold is $\epsilon=10^{-8}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-8}$.
-Figure 8 shows the error of the numerical steady-state solution, reached using ten time steps, relative to the prescribed analytical steady state as a function of mesh size.
-The advection CFL number, $\lambda_{A}$, for a fixed time step $\Delta t=1000$ ranges from $1,190$ to $36,076$ as the mesh is refined.
-The results are obtained by rescaling the numerical solution so that its normalization constant is the same as the analytical one.
-They confirm second-order spatial convergence and demonstrate the ability of our algorithm to take large timesteps to reach the correct steady state with expected asymptotic accuracy.
+Figure 8 shows the error of the numerical steady-state solution relative to the prescribed analytical steady state as a function of mesh size.
+The solution is obtained using ten time steps of size $\Delta t=1000$.
+The advection CFL number $\lambda_{A}$ ranges from $1,190$ to $36,076$ as the mesh is refined.
+The error is obtained by comparing the numerical and analytical solutions, with the numerical solution rescaled so that its normalization constant is the same as the analytical one.
+The results confirm second-order spatial convergence and the ability of the algorithm to solve directly for the correct steady state.
 Figure 8: Example 4.3.
@@
 5 Conclusion
-We have proposed a computationally efficient adaptive-rank implicit algorithm for solving time-dependent advection-diffusion partial differential equations with variable coefficients.
-Our algorithm is based on several key strategies: constructing dimension-wise basis from xKrylov subspaces, projecting the original matrix equation onto a reduced system to solve for a coefficient matrix, devising an effective Averaged-Coefficient Sylvester (ACS) preconditioner for the reduced system, and a computationally efficient procedure for the residual evaluation.
-This combined approach achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$, with $r$ being the rank of the solution during Krylov iteration process and $d$ the dimensionality, effectively mitigating the curse of dimensionality and making it comparable in efficiency to the constant-coefficient case [14].
-Such complexity scaling is numerically verified with our extensive tests, with the maximal rank $r$ found to be independent of the 1D grid size $N$ for practical grid resolutions and convergence tolerances.
+We propose a computationally efficient adaptive-rank implicit algorithm for solving time-dependent advection-diffusion partial differential equations with variable coefficients.
+We formulate the corresponding implicit system as a GSE, for which we follow a projection type approach, namely: constructing dimension-wise basis from xKrylov subspaces, projecting the original matrix equation onto a reduced system, and solving for the coefficient matrix.
+Our approach innovates in the way we construct the Krylov subspaces, and in the solution of the reduced system, which leverages GMRES effectively preconditioned with an effective Averaged-Coefficient Sylvester (ACS) approximation.
+The approach achieves a 1D-like computational complexity (of $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{d+1})$, with $r$ being the rank of the solution during Krylov iteration process and $d$ the dimensionality) and memory storage [of $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$], effectively mitigating the curse of dimensionality when the solution rank is bounded and making it comparable in efficiency to the constant-coefficient case [14].
+Such complexity scaling is numerically verified with extensive tests, with the maximal rank $r$ found to be independent of the 1D grid size $N$ for practical grid resolutions and convergence tolerances.
 Future research directions include extending the algorithm to higher dimensions and nonlinear PDEs.
@@
 with $\phi^{1}_{i+\frac{1}{2}}\phi^{2}_{j}=\phi^{1}(x_{i+\frac{1}{2}})\phi^{2}(y_{j}).$ To express the discretization in matrix form, we define:
-  • $\mathbf{F}$: matrix of solution values, $[\mathbf{F}]_{i,j}=f_{i,j}$.
+  • $\mathbf{F}$: matrix of solution values, $\mathbf{F}]_{i,j}=f_{i,j}$.
   • $\Phi^{1}_{+}$: diagonal matrix with elements $\left[\Phi^{1}_{+}\right]_{i,i}=\phi^{1}_{i+\frac{1}{2}}$.
@@
 The nodes $i=0$, $i=N_{x}-1$ rest at the boundary and enforce homogeneous Dirichlet boundary conditions (other boundary conditions can be readily considered):
-$D_{+}=\frac{1}{\Delta x}\begin{pmatrix}-1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\\ 0&\cdots&0&0&-1\end{pmatrix},D_{-}=\frac{1}{\Delta x}\begin{pmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\end{pmatrix}.$ (40)
+$D_{+}=\frac{1}{\Delta x}\begin{pmatrix}-1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\\ 0&\cdots&0&0&-1\end{pmatrix},D_{-}=\frac{1}{\Delta x}\begin{pmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\end{pmatrix}$.
+(40)
 Using these definitions, the discrete diffusion term in Eq. (39) can be represented in matrix form as:
@@
 We discretize this term on a uniform grid with spacing $\Delta x$ in the $x$-direction and $\Delta y$ in the $y$-direction.
-To preserve linearity of our discretization without introducing too much numerical dissipation, we employ a central difference scheme to approximate the spatial derivative.
+To preserve linearity of the discretization without introducing too much numerical dissipation, we employ a central difference scheme to approximate the spatial derivative.
 This discretization is monotonicity-preserving when the following condition is satisfied:
@@
 where $\sigma_{\text{max}}$ is the maximum magnitude of the advection velocity and $\phi_{\text{max}}$ is the maximum diffusion coefficient.
-This condition is practical in our formulation due to the very favorable scaling of our low-rank representation with mesh size and rank.
+This condition is practical in our formulation due to the very favorable scaling of the low-rank representation with mesh size and rank.
 The discrete approximation at the grid point $(i,j)$ is:
@@
 To express the discretization in matrix form, we define:
-  • $\mathbf{F}$: matrix of solution values, $[\mathbf{F}]_{i,j}=f_{i,j}$.
+  • $\mathbf{F}$: matrix of solution values, $\mathbf{F}]_{i,j}=f_{i,j}$.
   • $\Sigma^{1}_{+}$ and $\Sigma^{1}_{-}$: diagonal matrices with elements $\left[\Sigma^{1}_{+}\right]_{i,i}=\sigma^{1}_{i+\frac{1}{2}}$ and $\left[\Sigma^{1}_{-}\right]_{i,i}=\sigma^{1}_{i-\frac{1}{2}}$.
@@
 We also define the difference matrices in the $x$-direction:
-$S_{+}=\frac{1}{2}\begin{pmatrix}1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\\ 0&\cdots&0&0&1\end{pmatrix},S_{-}=\frac{1}{2}\begin{pmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\end{pmatrix}.$ (46)
+$S_{+}=\frac{1}{2}\begin{pmatrix}1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\\ 0&\cdots&0&0&1\end{pmatrix},S_{-}=\frac{1}{2}\begin{pmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 0&1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&1\end{pmatrix}$.
+(46)
 Using these definitions, the discrete advection term in the $x$-direction can be represented as:
@@
 $c_{1}$ $a_{11}$ 0 $\ldots$ 0
-$c_{2}$ $a_{21}$ $a_{22}$ldots 0
+$c_{2}$ $a_{21}$ $a_{22}$ $\ldots$ 0
 $\vdots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$
-$c_{s}$ $a_{s1}$ $a_{s2}$ldots $a_{ss}$
-$b_{1}$ $b_{2}$ldots $b_{s}$
+$c_{s}$ $a_{s1}$ $a_{s2}$ $\ldots$ $a_{ss}$
+$b_{1}$ $b_{2}$ $\ldots$ $b_{s}$
 We focus on stiffly accurate DIRK methods, where the coefficients satisfy $b_{i}=a_{si}$ for $i=1,\ldots,s$, ensuring that the final solution update corresponds to the solution at the last DIRK stage.
-The Butcher tables for DIRK2 and DIRK3 are presented in Tables 3 and 3, respectively.
+The Butcher tables for DIRK2 and DIRK3 are presented in Tables 4 and 4, respectively.
 Table 3: DIRK2 Butcher table, $\gamma=1-\frac{\sqrt{2}}{2}$.
 $\gamma$ $\gamma$ 0
-$1$ $1-\gamma$ $\gamma$
+11 $1-\gamma$ $\gamma$
 $1-\gamma$ $\gamma$
@@
 $\frac{1+x}{2}$ $\frac{1-x}{2}$ $x$ 0
-$1$ $-\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $-\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
-$-\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $-\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
+11 $\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
+$\frac{3x^{2}}{2}+4x-\frac{1}{4}$ $\frac{3x^{2}}{2}-5x+\frac{5}{4}$ $x$
 Similarly to the backward Euler case, the matrix-based discretization of the advection-diffusion equation using the operator $\mathscr{L}$ defined in (10) results in a GSE that must be solved at each DIRK stage.
 The corresponding DIRK scheme for the matrix differential equation (10) from $t^{(0)}$ to $t^{(1)}$ is written as:
-$\begin{split}\mathbf{F}^{(k)}&=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k}a_{k\ell}\mathbf{Y}_{\ell},\qquad k=1,2,\ldots,s,\end{split}$ (48a)
-$\begin{split}\mathbf{Y}_{k}&=\mathscr{L}(\mathbf{F}^{(k)};t^{(k)}),\qquad t^{(k)}=t^{0}+c_{k}\Delta t,\qquad k=1,2,\ldots,s,\end{split}$ (48b)
+$\begin{split}\mathbf{F}^{(k)}&=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k}a_{k\ell}\mathbf{Y}_{\ell},\hskip 18.49988ptk=1,2,\ldots,s,\end{split}$ (48a)
+$\begin{split}\mathbf{Y}_{k}&=\mathscr{L}(\mathbf{F}^{(k)};t^{(k)}),\hskip 18.49988ptt^{(k)}=t^{0}+c_{k}\Delta t,\hskip 18.49988ptk=1,2,\ldots,s,\end{split}$ (48b)
 $\begin{split}\mathbf{F}_{1}&=\mathbf{F}^{(s)}=\mathbf{F}_{0}+\Delta t\sum\limits_{k=1}^{s}b_{k}\mathbf{Y}_{k}.\end{split}$ (48c)
 Within each stage $k$ of the DIRK scheme, one must solve the residual equation given by:
-${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\quad k=1,2,\ldots,s.$ (49)
+${\bf R}^{(k)}_{\mathscr{L}}=\mathbf{F}^{(k)}-a_{kk}\Delta t\mathscr{L}(\mathbf{F}^{(k)})-\mathbf{B}^{(k)}=\mathbf{0},\qquad\text{with}\qquad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell},\qquad k=1,2,\ldots,s.$ (49)
 Expanding the operator $\mathscr{L}$ in the residual equation above results in the following $k$-th stage GSE:
-$\mathbf{F}^{(k)}-a_{k,k}\Delta t\left(T_{1}\mathbf{F}^{(k)}\Phi_{2}{{}^{\top}}+\Phi_{1}\mathbf{F}^{(k)}T_{2}^{\top}+T_{3}\mathbf{F}^{(k)}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}^{(k)}T_{4}^{\top}\right)=\mathbf{B}^{(k)},\quad k=1,2,\ldots,s.$ (50)
+$\mathbf{F}^{(k)}-a_{k,k}\Delta t\left(T_{1}\mathbf{F}^{(k)}\Phi_{2}{{}^{\top}}+\Phi_{1}\mathbf{F}^{(k)}T_{2}^{\top}+T_{3}\mathbf{F}^{(k)}\Sigma_{2}^{\top}+\Sigma_{1}\mathbf{F}^{(k)}T_{4}^{\top}\right)=\mathbf{B}^{(k)},\qquad k=1,2,\ldots,s.$ (50)
 For the DIRK update, we adopt the strategy outlined in [14], with minor modifications to improve efficiency.
@@
 C.2 Adaptive-Rank Strategy for High-Order Temporal Discretization via DIRK
-[…]
-[…]
-$\mathbf{R}^{(k)}_{\tilde{\mathscr{L}}}=\mathbf{F}_{1}-a_{kk}\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}-\mathbf{B}^{(k)}=\mathbf{0},\quad\text{with}\quad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell}.$ (51)
+The high-order DIRK method features a stage-by-stage generalized-Sylvester-equation solve.
+The strategy follows closely that of the backward Euler adaptive-rank treatment, i.e., we construct an approximate $\mathscr$$\tilde${L}} used for both the basis construction of $U_{1}$ and $V_{1}$ and the preconditioning of the reduced system to solve $S_{1}$.
+At each stage $k$ of the DIRK scheme, the corresponding DIRK formulation for the matrix differential equation from $t^{(0)}$ to $t^{(1)}$ requires solving an SE for the approximated operator $\mathscr$$\tilde${L}} by solving the $k$th-stage residual:
+$\mathbf{R}^{(k)}_{\tilde{\mathscr{L}}}=\mathbf{F}_{1}-a_{kk}\Delta t\,\tilde{\mathscr{L}}(\mathbf{F}_{1})-\mathbf{F}_{0}-\mathbf{B}^{(k)}=\mathbf{0},\qquad\text{with}\qquad\mathbf{B}^{(k)}=\mathbf{F}_{0}+\Delta t\sum\limits_{\ell=1}^{k-1}a_{k\ell}\mathbf{Y}_{\ell}.$ (51)
 Expanding this residual using $\tilde{\mathscr{L}}$, defined in (15), leads to the $k$th-stage SE:
@@
 The key idea is to consider the basis fixed across DIRK stages, and only evolve the matrix of coefficients ${{\bf S}}_{1}$ through the stages.
-The proposed procedure is as follows:
+The procedure is as follows:
   1. Step 1.
@@
 To further improve computational efficiency, from (53), we have,
-      $\tilde{Y}_{\ell}$ $=\frac{1}{a_{\ell\ell}\Delta t}\left({{\bf S}^{(\ell)}}-{\tilde{{\bf B}}^{(\ell)}}\right),$
+[…]
       leading to an efficient computation of $\tilde{{\bf B}}^{(k)}$:
-      $\tilde{{\bf B}}^{(k)}=\tilde{{\bf B}}_{1}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({{\bf S}^{(\ell)}}-{\tilde{{\bf B}}^{(\ell)}}\right),$ (55)
+[…]
       with $\tilde{{\bf B}}_{1}$ defined in (20). This avoids the need to evaluate the full size ${\bf B}^{(k)}$ in (51).
@@
 We now extend the formulation to the multi-rank case, as described in (6). Using a DIRK method for temporal discretization, the $k$-th stage GSE becomes:
-$\mathbf{F}^{(k)}-a_{kk}\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}^{(k)}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}^{(k)}{T_{2,j}}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}^{(k)}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}^{(k)}{T_{4,l}}^{\top}}_{\mathscr{L}(\mathbf{F}^{(k)})}\right)=\mathbf{B}^{(k)},\quad k=1,2,\ldots,s$ (56)
+$\mathbf{F}^{(k)}-a_{kk}\Delta t\left(\underbrace{\sum_{i=1}^{\ell_{x}}T_{1,i}\mathbf{F}^{(k)}{\Phi_{i}^{2,x}}^{\top}+\sum_{j=1}^{\ell_{y}}\Phi_{j}^{1,y}\mathbf{F}^{(k)}{T_{2,j}}^{\top}+\sum_{k=1}^{k_{x}}T_{3,k}\mathbf{F}^{(k)}{\Sigma_{k}^{2,x}}^{\top}+\sum_{l=1}^{k_{y}}\Sigma_{l}^{1,y}\mathbf{F}^{(k)}{T_{4,l}}^{\top}}_{\mathscr{L}(\mathbf{F}^{(k)})}\right)=\mathbf{B}^{(k)},\qquad k=1,2,\ldots,s$ (56)
 where $\mathbf{B}^{(k)}$ is as defined in (51).
@@
 Time step size $\Delta t$;
-Tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, ${\epsilon}_{\text{tol}}$;
+Tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, ${\epsilon}_{\texttt{tol}}$;
 Output: Updated bases $U_{1}$, $V_{1}$;
 Truncated singular values $S_{1}$;
-1 Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eq. (57);
+1
+21exCompute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eq. (57);
 3 Compute $l=2+\ell_{x}+\ell_{y}+k_{x}+k_{y}$;
 4 Set $U^{(i)}=U_{0}$ and $V^{(i)}=V_{0}$ for $i=1,\dots,l$;
-4 Set $U_{1}=U_{0}\quad V_{1}=V_{0}$;
+5 Set $U_{1}=U_{0}\qquad V_{1}=V_{0}$;
 6 while not converged do
@@
 12    for $k=1$ to $s$ do
-12               Compute $\tilde{{\bf B}}^{(k)}=\tilde{{\bf B}}^{(1)}+\Delta t\sum_{\ell=1}^{k-1}\frac{a_{k\ell}}{a_{\ell\ell}}\left({\bf S}^{(\ell)}-\tilde{{\bf B}}^{(\ell)}\right)$;
+[…]
 14        Solve the reduced equation (59) for ${\bf S}^{(k)}$ using GMRES with ACS preconditioner (60) to tolerance $\epsilon_{\texttt{GMRES}}$;
@@
 17    // Step 3.
-17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
-18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
+18    Compute $\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
+19    $\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
 20    // Step 4.
 21    Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(-\tilde{B}^{(s)},S_{1,\epsilon},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)R_{V}^{\top}\right\|$;
-21
+22
 23   if $\|\mathbf{R}\|/\|\mathbf{F}_{0}\|\geq\epsilon_{\texttt{tol}}$ then
 24        Return to Step 1 to augment bases further;
-24
+25
 26   else
@@
 29        Exit loop;
-29
 30
+31
 Algorithm 5 s-Stages Adaptive-Rank DIRK Integrator for the GSE.
 Appendix D Three-Dimensional Formulation
-We outline our approach to 3D next. In 3D, we replace the SVD algorithm with a tensor Tucker decomposition, specifically tailored for the three (and higher) dimension tensor-product grids.
+We outline the approach to 3D next. In 3D, we replace the SVD algorithm with a tensor Tucker decomposition, specifically tailored for the three (and higher) dimension tensor-product grids.
 D.1 Tucker Tensor Equations
@@
 This is represented as:
-[…]
+$\mathbf{F}_{0}=S_{0}\times_{1}U_{0}\times_{2}V_{0}\times_{3}W_{0}\xrightarrow{\text{evolve in time}}\mathbf{F}_{1}=S_{1}\times_{1}U_{1}\times_{2}V_{1}\times_{3}W_{1},$ (65)
 where $U_{0}\in\mathbb{R}^{N_{1}\times r^{x}_{0}}$, $V_{0}\in\mathbb{R}^{N_{2}\times r^{y}_{0}}$, $W_{0}\in\mathbb{R}^{N_{3}\times r^{z}_{0}}$, $U_{1}\in\mathbb{R}^{N_{1}\times r^{x}_{1}}$, $V_{1}\in\mathbb{R}^{N_{2}\times r^{y}_{1}}$, and $W_{1}\in\mathbb{R}^{N_{3}\times r^{z}_{1}}$ are orthonormal bases for their respective dimensions, and $S_{0}\in\mathbb{R}^{r^{x}_{0}\times r^{y}_{0}\times r^{z}_{0}}$ and $S_{1}\in\mathbb{R}^{r^{x}_{1}\times r^{y}_{1}\times r^{z}_{1}}$ are core tensors at the corresponding times.
@@
 The total complexity in three dimensions is therefore $\mathcal{O}(Nr^{2}+r^{4})$.
+Appendix E Technical Details of the Hypothetical Multigrid Comparison
+We provide here a brief explanation of the assumptions involved in the performance comparison between a hypothetical optimally scaling full-rank multigrid method and the adaptive-rank approach.
+Optimal multigrid methods exhibit a computational scaling of $\mathcal{O}(N^{d}\log N)$ [21].
+Let us denote the wall-clock time (WCT) of a multigrid method as:
+$WCT_{\mathrm{MG}}=C_{\mathrm{MG}}N^{d}\log N.$
+From Fig. 1-a, the wall-clock time of the adaptive-rank method may be written as:
+$WCT_{\mathrm{LR}}=C_{\mathrm{LR}}(0.0023\,N+55.2734).$
+The speedup of the adaptive-rank method vs. multigrid is given by the ratio:
+$\frac{WCT_{\mathrm{MG}}}{WCT_{\mathrm{LR}}}=C_{\mathrm{WCT}}\frac{N^{3}\log N}{0.0023\,N+55.2734},$ (68)
+where $C_{\mathrm{WCT}}=1.2\times 10^{-5}$ is found assuming that both methods achieve comparable wall-clock times for $N=100$. This formula is used to populate Table 1.